### v3

# Still More Shades of Null: An Evaluation Suite for Responsible Missing Value Imputation

Abstract.
Data missingness is a practical challenge of sustained interest to the scientific community. In this paper, we present Shades-of-Null, an evaluation suite for responsible missing value imputation. Our work is novel in two ways (i) we model realistic and socially-salient missingness scenarios that go beyond Rubin’s classic Missing Completely at Random (MCAR), Missing At Random (MAR) and Missing Not At Random (MNAR) settings, to include multi-mechanism missingness (when different missingness patterns co-exist in the data) and missingness shift (when the missingness mechanism changes between training and test) (ii) we evaluate imputers holistically, based on imputation quality, as well as on the predictive performance, fairness and stability of the models that are trained and tested on the data post-imputation.
We use Shades-of-Null to conduct a large-scale empirical study involving 23,940 experimental pipelines, and find that while there is no single best-performing imputation approach for all missingness types, interesting trade-offs arise between predictive performance, fairness and stability, based on the combination of missingness scenario, imputer choice, and the architecture of the predictive model. We make Shades-of-Null publicly available, to enable researchers to rigorously evaluate missing value imputation methods on a wide range of metrics in plausible and socially meaningful scenarios.
1. Introduction
As AI becomes more widely deployed into society, data — most importantly, openly accessible high quality AI-ready data — becomes a precious shared commodity. Among the factors affecting data quality is data missingness, a prevailing practical challenge of sustained interest to the data management, statistics and data science communities, and to the scientific community writ large (Little and Rubin, 1987; Rubin, 1976; Pigott, 2010; Mitra et al., 2023; Zhou et al., 2024; Shadbahr et al., 2023; Graham, 2009; Le Morvan et al., 2021; Zhang and Long, 2021a; Howell, 2007; Ilyas and Chu, 2019).
Specifically in data management, debates have been raging since the very inception of the field regarding whether — and how — to handle missing values, see  Date (1984) for a classic example. One difficulty, at the operational level, is that, while missing values are typically denoted with the token null, that may not always be the case. There could be hidden missing values in the data, like when ‘AL’ is selected as the state on a job application because it’s the first item in the drop-down list. Another difficulty, at the semantic level, is that a null can take on many meanings: it may signify that the value is unknown, inapplicable, intentionally withheld, etc.
In this paper, we do not concern ourselves with this semantic debate, and we also do not consider hidden missing values (Qahtan et al., 2018). Rather, we focus our attention on a very specific case. We are given a dataset $X$ (a single relation), in which the values of some features are missing, denoted by the marker null. The meaning of null is that the feature does take on a value in the real world (i.e., it is applicable) but that value is unobserved in $X$. Further, we intend to use $X$ in a machine learning (ML) setting: to train a model or to make predictions about the tuples in $X$ using a model that has already been trained and deployed. This model cannot deal with null values directly, and so these values must be imputed (filled in) as part of data pre-processing.
As our starting point, we will use Rubin’s missingness framework (Rubin, 1976) that, nearly 50 years since it was proposed, still remains the most popular approach to modeling missing data. Consider a dataset $X$ of $n$ samples, each with $p$ features, and an indicator $R$ such that $R_{i,j}=1$ when the value of the $j$’s feature of $X_{i}$ is missing: $X_{i,j}\text{ is null}$, and $R_{i,j}=0$ when that the value is observed: $X_{i,j}\text{is not null}$. Rubin identified three data missingness scenarios:
Missing Completely at Random (MCAR)
Consider a dataset of job applicants that includes their most recent salary and the number of years of relevant work experience. The MCAR assumption holds, for example, when a job applicant’s most recent salary is missing from their record for purely administrative reasons that are uncorrelated with either their (potentially missing) salary or with their (observed) experience. More formally, the probability that a feature value is missing is independent of $X$: $P(R|X)=P(R)$.
Missing at Random (MAR)
Suppose that job applicants with fewer years of relevant work experience are more likely to withhold their most recent salary (in hopes of being offered a higher salary by the prospective employer), and that this can be explained by observed covariates (i.e., by the number of years of work experience). The MAR assumption holds in this example, where the probability that some feature value is missing depends only on the observed covariates $X_{\text{obs}}$, and not on the missing feature values: $P(R|X)=P(R|X_{\text{obs}})$.
Missing Not at Random (MNAR)
Suppose that a job applicant’s most recent salary does not correlate with their number of years of experience but rather is based on some combination of their geographic location and the results of a skills test, neither of which are captured in the data. Further, suppose that individuals whose most recent salary was low may be likely to withhold their salary information (again, in hopes of being offered a higher salary by the prospective employer). The MNAR assumption holds in this example, because missingness is correlated with the missing value itself, and it cannot be explained by the observed covariates (i.e., by the number of years of experience): $P(R|X)\not=P(R|X_{\text{obs}})$.
Missing value imputation (MVI)
Rubin’s framework has formed the basis of a large body of technical work on missing value imputation, which has been covered in several excellent and comprehensive surveys (Abdelaal et al., 2023; Hasan et al., 2021; Le Quy et al., 2022; Mena et al., 2022; Pintz et al., 2022; Emmanuel et al., 2021; Peherstorfer et al., 2018; Lin and Tsai, 2020; Pigott, 2010; Zhou et al., 2024; Alabadla et al., 2022; García Laencina et al., 2010; Joel and Doorsamy, 2022; Ilyas and Chu, 2019). Broadly speaking, MVI approaches fall into several categories: (1) statistical, such as imputing missing observations with the median or mode (Schafer and Graham, 2002); (2) learning-based impute-then-classify, which use the observed data to iteratively learn and impute missing values, using k-nearest neighbors (Batista and Monard, 2003), clustering (Gajawada and Toshniwal, 2012), decision trees (Twala, 2009), or ensembles (Stekhoven and Bühlmann, 2012); and (3) joint data cleaning and model training (Karlaš et al., 2020; Krishnan et al., 2017, 2016), based on Rubin’s seminal multiple imputation framework (Rubin, 1987).
Beyond Rubin’s framework: Mixing scenarios and dealing with missingness shift.
Rubin’s framework, while clean and amenable to analysis, does not fully capture real-world missingness. First, it has been argued that the assumptions behind MCAR rarely hold in practice, and that most real-world data exists on a continuum between MAR and MNAR, depending on how that data was collected (Graham, 2009). Second, while it is useful to distinguish between missingness scenarios based on their statistical properties to draw valid inferences, we are likely to encounter multi-mechanism missingness in practice, where several scenarios may co-exist in the same dataset, affecting subsets of its features, tuples, or both (Zhou et al., 2024). For example,  Mitra et al. (2023) introduce the notion of a data missingness life cycle and posit that data pre-processing steps such as assimilating data from different sources or from different heterogeneous populations (each with their own missingness patterns) can result in ‘structured missingness’ that cannot be represented using Rubin’s framework. Third, it has been observed that, in the context of data-centric AI, the assumptions that hold about missingness in the training data may no longer hold post-deployment, a phenomenon termed missingness shift (by analogy with data distribution shift) (Zhou et al., 2023).
Evaluating MVI techniques.
Despite an abundance of MVI techniques using different model architectures being proposed each year, making meaningful progress on learning from incomplete data ultimately comes down to how realistically we model missingness: our proposed methods are only as good as the evaluation scenarios we rigorously test them on. If our evaluation scenarios do not realistically capture the kinds of missingness that we encounter in practice, then our data cleaning methods are unlikely to be effective. Further, newly proposed MVI techniques are often only evaluated on the correctness of the imputation (imputation quality) and the correctness of the predictive model (accuracy or F1). However, in recent years, the scientific community is starting to uncover trade-offs between MVI efficacy and other normatively important dimensions of model performance, including algorithmic fairness (Guha et al., 2024; Feng, 2023), and stability and robustness (Zhou et al., 2024), beyond accuracy.
Missingness as a form of pre-existing bias.
Consider again the job applicant screening example, and suppose that the applicants’ gender and age are among the features. Female job applicants who suspect that they have been receiving lower salaries than their male colleagues, despite having comparable qualifications and job experience, may withhold salary information more frequently then men, hoping to narrow the gender wage gap in their next position. As a result there will be more missing salary values for women, and the probability that the value is missing depends on the observed covariate, gender. This situation is consistent with MAR missingness, and it points to pre-existing bias — a situation where data reflects a history of societal discrimination (Friedman and Nissenbaum, 1996). For our final example, suppose that disability status is among the features. Job applicants who live with a disability are more likely to omit a value for this feature. Assuming that this feature does not correlate with the other features, this situation is consistent with MNAR, with missingness itself serving as a proxy for disadvantage.
Summary of Contributions
We implemented an experimental benchmark called Shades-of-Null to rigorously and comprehensively evaluate state-of-the-art MVI techniques on a variety of realistic missingness scenarios (including single- and multiple-mechanism missingness and missingness shift), on a suite of evaluation metrics (including fairness and stability), in the context of data pre-processing in a machine learning pipeline.
Our work is (1) novel: to the best of our knowledge, the settings of multi-mechanism missingness and missingness shifts have not been empirically studied before; (2) comprehensive: we evaluate a suite of 8 MVI techniques on 7 benchmark datasets using 6 model types, running a total of 23,940 pipelines, and is the first study of such scale in the missing data domain, to the best of our knowledge; (3) normatively grounded: we focus on decision-making contexts such as lending, hiring, and healthcare, where missingness is socially salient. Mitigating social harm such as algorithmic discrimination is a leading concern in these domains (Barocas and Selbst, 2016), and we evaluate the impact of MVI approaches on downstream model fairness and stability (which have been understudied in the context of missing data), in addition to classically studied imputation quality and model correctness metrics.
While developing the Shades-of-Null evaluation suite, we found and fixed several bugs in existing MVI implementations, including data leakage and omitted hyperparameter tuning. See Appendix A.2 for details.
We make Shades-of-Null publicly available [^1] https://github.com/FalaahArifKhan/data-cleaning-stability and hope to enable researchers to comprehensively and rigorously evaluate new MVI methods on a wide range of evaluation metrics, under plausible and socially meaningful missingness scenarios.
2. Related Work
Missing value imputation techniques
Learning-based approaches have become increasingly popular, and include k-nearest neighbors, decision trees, support vector machines, clustering, and ensembles (Batista and Monard, 2003; Lin and Tsai, 2020; Hasan et al., 2021).  Zhou et al. (2024) and Liu et al. (2023) review deep learning-based approaches (variational auto-encoders and generative adversarial networks) and representation learning (graph neural networks and diffusion-based methods). Multiple imputation (Schafer and Graham, 2002; Zhou et al., 2024) and expectation maximization (Pigott, 2010; Wang et al., 2024) are also influential, but too computationally expensive to be popular in practice (Alabadla et al., 2022).
Beyond impute-then-classify, the data management community has proposed holistic methods like CPClean (Karlaš et al., 2020) and ActiveClean (Krishnan et al., 2016), that jointly perform data cleaning and model training, deriving from the multiple imputation framework (Rubin, 1987). These methods detect and repair a variety of errors including outliers, mislabels, duplicates, and missing values, and hence are less directly optimized to model missingness, instead focusing on improving data quality holistically. BoostClean (Krishnan et al., 2017) aims to reduce the human effort in error repair by learning efficiently from a few gold standard annotations (from a human oracle).
Evaluating MVI techniques
Empirical studies have been overwhelmingly focused on medical datasets, and only evaluate missingness under MCAR (Lin and Tsai, 2020; Alabadla et al., 2022; Farhangfar et al., 2008). Further, most proposed methods only perform direct evaluation, using metrics such as MAE, MSE, RMSE, and AUC (Hasan et al., 2021; Joel and Doorsamy, 2022). Very few studies use both direct and indirect evaluation, and even studies that include indirect evaluation only evaluate predictive accuracy (Lin and Tsai, 2020). This leaves a significant research gap, specially for desiderata such as stability and self-consistency, which have recently been shown to be highly sensitive to data-centric interventions (Black and Fredrikson, 2021; Black et al., 2022; Cooper et al., 2023), and unfairness, which can be worsened by improperly chosen data cleaning techniques (Guha et al., 2024). Additionally, the performance of MVI techniques under multi-mechanism missingness (Zhou et al., 2024) and missingness shifts (Zhou et al., 2023) remains unexplored, despite these conditions being more likely to occur in practice due to distribution shifts in production deployments (Graham, 2009).
Notably, and worryingly, overwhelming evidence in the literature indicates that there is no single “best-performing” MVI approach on accuracy  (Hasan et al., 2021; Lin and Tsai, 2020; Shadbahr et al., 2023; García Laencina et al., 2010; Farhangfar et al., 2008), and that model performance (narrowly measured based on ‘correctness’ thus far) depends on dataset characteristics such as size and correlation between variables (Alabadla et al., 2022) and missingness rates in the train and test sets (Lin and Tsai, 2020; Shadbahr et al., 2023).
Fairness and missingness
There has been some recent interest in studying the social harm that can come from poorly chosen MVI techniques (Caton et al., 2022; Zhang and Long, 2022, 2021a, 2021b; Jeong et al., 2022; Wang and Singh, 2021; Martínez-Plumed et al., 2019; Feng et al., 2024). Most empirical studies (Wang and Singh, 2021; Zhang and Long, 2021b, 2022; Caton et al., 2022) have worked with the COMPAS (Larson et al., 2016) and Adult  [^2] https://archive.ics.uci.edu/dataset/2/adult datasets, the latter of which has been “retired” from community use due to issues with provenance (Ding et al., 2021). Further, these experimental studies employ randomly-generated missingness: usually by randomly sampling or using a fixed set of columns, and randomly picking rows in which to replace values with null. We critique this approach, since detecting and mitigating unfairness requires broader socio-technical thinking, such as having higher rates of missingness for minority groups and in features that are highly correlated with sensitive attributes (called proxy variables in the fairness literature) (Caton et al., 2022).
A notable exception is  Martínez-Plumed et al. (2019), who map social mechanisms such as prejudicial access and self-reporting bias to different categories of missingness such as missing-by-design and item non-response. Further, they analyse the correlations between all features to study the effect of different kinds of missingness. We use a similar methdology to simulate realistic missingnesses in this work. Importantly, however, we find conceptual limitations in their framing of algorithmic fairness. Martínez-Plumed et al. (2019) write:“The surprising result was to find that, […] the examples with missing values seem to be fairer than the rest.” Asserting that some rows of data are more or less fair does not make sense, in that samples (e.g., job applicants) do not demonstrate the property of being fair. It is the model (e.g., used to make hiring decisions) that asserts the property of being fair or not, by the inclusion or exclusion of certain samples amongst those given a positive outcome. We re-interpret the findings of  Martínez-Plumed et al. (2019) to say that excluding samples with missingness can increase model unfairness, which is further evidence against the use of deletion as a strategy for dealing with missing data.
Zhang and Long (2021b) evaluate MVI methods on imputation fairness, defined as the imputation accuracy difference on privileged and disadvantaged demographic groups. They find that (1) higher missingness disparity leads to higher imputation unfairness; (2) higher rates of missingness lead to higher imputation unfairness; (3) more imbalance in the data (or larger disparity in the sample sizes across groups) leads to higher imputation unfairness; (4) the same imputation method has different prediction fairness under different missingness mechanisms. Their evaluation is limited to randomly generated null values in COMPAS. We extend their analysis of imputation fairness to additional datasets, missingness scenarios and imputation fairness definitions (such as KL divergence difference).
Zhang and Long (2022) introduce the notion of imputation fairness risk and provide bounds for “correctly specified” imputation methods. While there is very little theoretical work in this space, and this is a commendable contribution, we call into question the broader implications of this result: it is well-known that imputation quality metrics do not fully capture downstream model performance (Zhang and Long, 2021a). Put differently, a classifier built on poor imputation quality can still give good downstream performance (Shadbahr et al., 2023). With this in mind, we ask the question: Does minimizing imputation unfairness minimize model unfairness? Our empirical findings indicate it does not, as is discussed in Section 4.4.
Missingness and stability. We are not aware of any work investigating the effect of missing value imputation on model stability.
3. Benchmark Overview
3.1. Methodology for Simulating Missingness
We start with datasets in which there are no null values, and then simulate missingness. We make this choice because we are interested in comparing MVI performance under single-mechanism versus multi-mechanism missingness, and under missingness shifts, and, to the best of our knowledge, there are no datasets with naturally-occurring documented missingness of this form.
Our methodology for simulating missingness is based on evaluation scenarios, defined by the missingness mechanism during training and testing, shown in Table 1: (1) single-mechanism missingness, injected similarly into train and test sets (S1 - S3); (2) single-mechanism missingness, injected differently into train and test sets (missingness shift) (S4 - S9); and (3) missingness is mixed, to include all three missingness mechanisms, and is injected similarly into train and test sets (S10).
Table 1. Table 1: Evaluation Scenarios
Train     Test
Scenario MCAR MAR MNAR MCAR MAR MNAR
S1
S2
S3
S4
S5
S6
S7
S8
S9
S10
For each dataset in our study (see Section 3.5), we designed socially-salient missingness scenarios corresponding to the three missingness mechanisms (MCAR, MAR, MNAR). Following (Joel and Doorsamy, 2022; Martínez-Plumed et al., 2019), we selected columns in which to inject null values based on the correlation coefficients between all covariates and the target variable, as well as feature importance computed using scikit-learn’s built-in functionality. We used sensitive attributes in the design of MAR missingness to simulate pre-existing bias, as described in Section 1, and injected higher rates of missingness for disadvantaged groups wherever possible (in some cases there were too few samples from disadvantaged groups), aiming to reflect realistic instances of disparate non-response caused by disparate access, distrust, or procedural injustice (Akpinar et al., 2024). Furthermore, we selected columns that may plausibly contain missingness. For example, blood pressure or cholesterol in diabetes is expected to always be observed, but family history or physical activity might be omitted or withheld.
Table 3 presents missingness scenarios for the diabetes dataset, based on the correlation coefficients and feature importances in Figure 2. Details for other datasets are available in Appendix B.2.
3.2. Missing Value Imputation (MVI) Techniques
As discussed in Section 2, there is a large number of competitive MVI techniques in the literature. For our study, we picked 8 techniques, from 5 broad categories in the taxonomy of  Emmanuel et al. (2021), namely: (1) deletion; (2) statistical techniques: median-mode and median-dummy; (3) machine learning-based techniques: miss-forest (Stekhoven and Bühlmann, 2011) and clustering (Gajawada and Toshniwal, 2012); (4) deep learning-based techniques: datawig (Biessmann et al., 2019) and auto-ml (Jäger et al., 2021); and (5) multiple imputation: boostclean (Krishnan et al., 2017). Additional details about these methods are deferred to Appendix A.1.
3.3. Evaluation Metrics
Following (Lin and Tsai, 2020; Hasan et al., 2021), we evaluate MVI techniques in two ways: directly using imputation quality metrics and indirectly based on downstream model performance.
3.3.1. Imputation Quality
Shadbahr et al. (Shadbahr et al., 2023) report that distributional metrics capture downstream model performance better than classically-used discrepancy metrics. To confirm or refute this claim, we use a mix of both. To assess agreement with true values, we compute Root Mean Square Error (RMSE) for numerical features and F1 score for categorical features. To assess distributional alignment, we compute KL-divergence (i.e., the Shannon entropy) between the true and the predicted feature distributions, for both numerical and categorical features, measured for the imputed columns only as well as for the full dataset. For categorical features, we obtain the probability distributions using the value_counts method with normalization from pandas. For numerical features, we use Gaussian kernel density estimation from scipy, with 1000 samples. Finally, to assess imputation fairness (Zhang and Long, 2021b, 2022), we compute F1 score difference, RMSE difference, and KL divergence difference between privileged (priv) and disadvantaged (dis) groups.
3.3.2. Model Performance
To assess the impact of MVI techniques on model correctness, we report the F1 score because it is a more reliable metric than accuracy for imbalanced data.
For evaluating model stability, we report average Label Stability (Darling and Stracuzzi, 2018; Khan et al., 2023) over the full test set (closely related to the self-consistency metric from  Cooper et al. (2023)), computed per-sample for binary classification as Label Stability = $\frac{|B_{+}-B_{-}|}{B}$, where $B_{+}$ is the number of times the sample is classified into the positive class and $B_{-}$ is the number of times the sample is classified into the negative class, and $B=B_{+}+B_{-}$ models are trained by bootstrapping over the train set. We set $B=50$ in all our experiments.
Lastly, we report model fairness based on group-specific error rates, namely True Positive Rate Difference (TPRD), True Negative Rate Difference (TNRD), Selection Rate Difference (SRD), and Disparate Impact (DI). (Note that DI computes the ratio of selection rates, but we refer to it as DI as is standard in the literature (Feldman et al., 2015).)
$TPRD=\frac{TP_{dis}}{TP_{dis}+FN_{dis}}-\frac{TP_{priv}}{TP_{priv}+FN_{priv}}$
$TNRD=\frac{TN_{dis}}{TN_{dis}+FP_{dis}}-\frac{TN_{priv}}{TN_{priv}+FP_{priv}}$
$SRD=\frac{TP_{dis}+FP_{dis}}{N_{dis}}-\frac{TP_{priv}+FP_{priv}}{N_{priv}}$
$DI=\frac{TP_{dis}+FP_{dis}}{N_{dis}}\div\frac{TP_{priv}+FP_{priv}}{N_{priv}}$
3.4. Shades-of-Null Architecture
The architecture of Shades-of-Null is shown in Figure 1. Our goal when developing this evaluation suite was to support the comprehensive evaluation of different MVI techniques (Section 3.2), under different missingness scenarios (Section 3.1), according to different aspects of model performance (Section 3.3). Shades-of-Null is novel in several ways. First, it is flexible and supports a variety of missingness scenarios, namely, single-mechanism, multi-mechanism and missingness shift. Second, it supports holistic evaluation of both imputation quality (including imputation fairness) and model performance — including accuracy, stability and fairness.
The central component of the architecture in Figure 1 is the benchmark controller that executes user-specified missingness scenarios by applying error injectors to input datasets. It then imputes missing values according to user-specified MVI technique(s) and pre-processes the imputed datasets using standard scaling for numerical features and one-hot encoding for categorical features. The controller then executes ML model training, including hyperparameter tuning. Finally, the evaluation module assesses imputation quality and model performance. For comprehensive model profiling, the evaluation module uses Virny (Herasymuk et al., 2024), a Python library that computes accuracy, stability and fairness metrics for a set of initialized models, for multiple sensitive attributes and their intersections. We store experimental results in MongoDB.
Figure 1. Shades-of-Null architecture
Table 2. Table 2: Dataset Information
Name Domain Size Sensitive Attributes
diabetes healthcare 905 sex
german finance 1,000 sex, age
folk-income finance 15,000 sex, race
law-school education 20,798 sex, race
bank marketing 40,004 age
heart healthcare 70,000 sex
folk-employment hiring 302,640 sex, race
Shades-of-Null incorporates two optimizations to streamline the execution of multiple experiments. Firstly, it decouples missing value imputation and model training. This enables the preservation of imputation results, including imputed datasets stored in a file system, facilitating their reuse during subsequent model training and evaluation stages. Secondly, the experimental pipeline within Shades-of-Null supports evaluation on multiple test sets simultaneously (e.g., with different rates or types of missingness), to speed up evaluation. For example, the running time of executing a pipeline with one training set and one test set is close to that of a pipeline with one train set and multiple test sets.
3.5. Datasets and Tasks
As noted in Section 1, in this work we are interested in missingness that is socially salient. With this in mind, we selected $7$ datasets from a variety of social decision-making contexts including lending, hiring, marketing, admissions, and healthcare, summarized in Table 2. For each dataset, there is a binary classification task, where being assigned a positive label corresponds to the attainment of a desirable social good (such as education, employment, or access to healthcare). We selected these datasets to have good coverage of (i) social domains and (ii) dataset sizes and other characteristics.
Table 3. Table 3: Missingness scenarios for an error rate of 30% for diabetes. SoundSleep is a numerical column; Family__\__Diabetes, PhysicallyActive and RegularMedicine are categorical columns.
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
MCAR SoundSleep, Family__\__Diabetes, PhysicallyActive, RegularMedicine N/A 0.3 0.3
MAR Family__\__Diabetes, RegularMedicine Sex 0.2 (female) 0.1 (male)
PhysicallyActive, SoundSleep Age 0.2 ($\geq 40$) 0.1 ($<40$)
MNAR Family__\__Diabetes Family__\__Diabetes 0.25 (yes) 0.05 (no)
RegularMedicine RegularMedicine 0.2 (yes) 0.1 (no)
PhysicallyActive PhysicallyActive 0.25 (none, $<\frac{1}{2}$ hour) 0.05 ($>\frac{1}{2}$ hour, $>1$ hour)
SoundSleep SoundSleep 0.2 ($<5$) 0.1 ($\geq 5$)
(a) Correlation with label
(b) Feature importance
Figure 2. EDA for designing missingness scenarios in diabetes.
diabetes [^3] https://www.kaggle.com/datasets/tigganeha4/diabetes-dataset-2019 (Tigga and Garg, 2020) was collected in India through a questionnaire including 18 questions related to health, lifestyle, and family background. A total of 952 participants are characterized by 17 attributes (13 categorical, 4 numerical) and a binary target variable that represents whether a person is diabetic. Here, sex is the sensitive attribute, with “female” as the disadvantaged group.
german [^4] https://archive.ics.uci.edu/dataset/144/statlog+german+credit+data (Hofmann, 1994) is a popular fairness dataset that contains records of creditworthiness assessments, classifying individuals as high or low credit risks. It contains information on 1,000 individuals characterized by 21 attributes (14 categorical, 7 numerical), including credit history, occupation and housing information. Here, sex and age are the sensitive attributes, with “female” and “age25absent25$\leq$ 25≤ 25” as the disadvantaged groups.
Folktables (Ding et al., 2021) is another popular fairness dataset derived from US Census data from all 50 states between 2014-2018. The dataset has several associated tasks, of which we selected two: (i) ACSIncome (folk-income) is a binary classification task to predict whether an individual’s annual income is above \$50,000, from 10 features (8 categorical, 2 numerical) including educational attainment, work hours per week, marital status, and occupation. We use data from Georgia from 2018, subsampled to 15k rows. (ii) ACSEmployment (folk-employment) is a binary classification task to predict whether an individual is employed, from 16 features (15 categorical, 1 numerical) including educational attainment, employment status of parent, military status, and nativity. We use data from California from 2018, with 302,640 rows (we do not subsample). In both tasks, sex and race are the sensitive attributes, with “female” and “non-White” as the disadvantaged groups.
law-school [^5] https://www.kaggle.com/datasets/danofer/law-school-admissions-bar-passage  (Wightman, 1998) was gathered through a survey conducted by the Law School Admission Council (LSAC) across 163 law schools in the US in 1991, and contains admissions records of 20,798 applicants, characterized by 11 attributes (5 categorical, 6 numerical), including LSAT scores and college GPAs. The task is to predict whether a candidate would pass the bar exam. Here, sex and race are the sensitive attributes, with “female” and “non-White” as the disadvantaged groups.
bank [^6] https://archive.ics.uci.edu/dataset/222/bank+marketing  (Moro et al., 2014) contains data from direct marketing campaigns by a Portuguese bank, between 2008 and 2013. It contains information on 40,004 potential customers, with 16 attributes (10 categorical, 6 numerical), including occupation, marital status, education, and a binary target that indicates whether the individual subscribed for a term deposit. Here, age is the sensitive attribute, with $<25$ and $>60$ as the disadvantaged group.
heart [^7] https://www.kaggle.com/datasets/sulianova/cardiovascular-disease-dataset consists of patient measurements with respect to cardiovascular diseases, with information of 70,000 individuals characterized by 11 attributes (6 categorical, 5 numerical), including age, height, weight, blood pressure, and a binary target indicating whether the patient has heart disease. Here, sex is the sensitive attribute, with “female” as the disadvantaged group.
3.6. Model Types
We evaluate predictive performance of 6 ML models: (i) decision tree (dt__\__clf) with a tuned maximum tree depth, minimum samples at a leaf node, number of features used to decide the best split, and criteria to measure the quality of a split; (ii) logistic regression (lr__\__clf) with tuned regularization penalty, regularization strength, and optimization algorithm; (iii) gradient boosted trees (lgbm__\__clf) with tuned number of boosted trees, maximum tree depth, maximum tree leaves, and minimum number of samples in a leaf; (iv) random forest (rf__\__clf) with a tuned number of trees, maximum tree depth, minimum samples required to split a node, and minimum samples at a leaf node (v) neural network, historically called the multi-layer perceptron (mlp__\__clf) with two hidden layers, each with 100 neurons, and a tuned activation function, optimization algorithm, and learning rate; (vi) a deep table-learning method called GANDALF (Joseph and Raj, 2022) (gandalf__\__clf) with a tuned learning rate, number of layers in the feature abstraction layer, dropout rate for the feature abstraction layer, and initial percentage of features to be selected in each Gated Feature Learning Unit (GFLU) stage.
4. Single and Multi-Mechanism Missingness
To simulate single-mechanism missingness (S1-S3 in Table 1) we inject 30% of each training and test sets with nulls, according to the missingness scenarios described in Section 3.1. For multi-mechanism or mixed missingness, when MCAR, MAR and MNAR co-exist (S10 in Table 1), we inject 10% of nulls for each of the three mechanism into both training and test sets, for a total of 30% nulls.
To evaluate model correctness, we report results for F1, see Appendix C.1 for accuracy results. For fairness, we use binary group definitions throughout. For datasets with two sensitive attributes, we define the doubly-disadvantaged group as disadvantaged (dis) and everyone else as privileged (priv). For example: on the law-school, folk-income and folk-employment datasets, non-White women are the dis group, and White women, non-White men and White men collectively form the priv group. We report results for TPRD, and defer results for other fairness metrics to Appendix C.1. For stability, we used a bootstrap of 50 estimators, each seeing a random 80% of the training set (Efron and Tibshirani, 1994). Higher values of F1 and label stability are better, and values of TPRD close to zero are better.
Different models are the best-performing on different datasets. In Figures 34 and 5, we report on the best-performing models (according to F1) for each dataset, and compare performance against a model trained on the clean training set using the same model architecture (performance shown as a blue line in the plots).
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 3. F1 of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Datasets are in increasing order by size. Blue line shows performance of the model trained on clean data.
4.1. Correctness of the Predictive Model
Figure 3 shows the F1 of models trained with different MVI techniques. We find interesting trends in MVI performance based on characteristics of the dataset and missingness type. All techniques are competitive for all missingness mechanisms, including mixed missingness, on heart and law-school. boostclean, which uses multiple imputation (MI), is otherwise only competitive on small datasets (diabetes and german), and only under MCAR and mixed missingness on german. None of the MVI techniques are able to match the F1 of the model trained on clean data on folk-employment, and this effect is strongest under MNAR (notably, stronger than under mixed missingness). boostclean shows particularly poor performance on folk-income, with a 0.08 decrease in F1 compared to other methods, for all missingness types. We discuss this unexpected performance of MI further in Section 6.
auto-ml, datawig and miss-forest are generally the best performing MVI techniques, but simpler statistical techniques (e.g., median-mode under MAR) are also competitive. This underscores the need to evaluate novel DL-based and ML-based approaches more holistically (e.g., on a variety of missingness scenarios) to ensure that they justify the additional training overhead and complexity they introduce, compared to simple statistical methods.
In line with conventional wisdom (Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), we find that deletion worsens predictive performance. This effect is strongest for small datasets like diabetes, with F1 decreasing as much as 0.1 under MNAR, compared to the model trained on clean data.
Note that the F1 score on the bank dataset is extremely low (0.32). This is due to severe class imbalance in the dataset (the base rate is 0.117, as reported in Table 8 in Appendix B). Further, from Figure 18 in Appendix C.1, we can see that the accuracy of the lgbm__\__clf is 0.89, which is marginally better than the accuracy reported using Logistic Regression (0.885) in (Le Quy et al., 2022) in Table 15.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 4. True Positive Rate Difference (unfairness) of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets ($x$-axis), and missingness mechanisms (subplots). Values close to 0 are desirable. Datasets are in increasing order by size. Blue line shows TPRD of the model trained on clean data.
4.2. Fairness of the Predictive Model
Figure 4 shows the effect of MVI on fairness, according to TPRD.  Wang and Singh (2021) posit that models will exhibit more unfairness under MAR and MNAR compared to MCAR, but we only find weak empirical evidence towards this, even for deletion. A nuance here is that we designed missingness scenarios, described in Section 3.1, to be realistic — including MAR scenarios where people from disadvantaged groups withhold information that could hurt their chances of getting the desired outcome. Hence, dropping these rows can in fact improve fairness under MAR and MNAR, as observed on bank.
In contrast to Wang and Singh (2021), we find that the effect of MVI on fairness is strongly correlated with fairness of the model trained on clean data, corroborating the findings of  Guha et al. (2024). Fairness depends on two things: dataset characteristics and model type. All MVI techniques except for boostclean have the same model type as the model trained on clean data (because they are impute-then-classify approaches) and generally preserve fairness of that model, under all missingness mechanisms. Notably, this is agnostic to whether the TPRD of the model trained on clean data is low (close to 0.01 on heart and folk-employment, and 0 on german) or high (close to -0.1 on folk-income and 0.2 on bank).
On the other hand, boostclean, is a joint data cleaning and model training approach and thereby constitutes it own model type, and shows fairness trends that deviate from the model trained on clean data. boostclean significantly improves fairness on
folk-income (TPRD close to 0, compared to -0.1 for the clean model) and bank (TPRD close to 0.1, compared to 0.2 for the clean model), but marginally worsens fairness on law-school (TPRD -0.1 compared to -0.08 for the clean model) and heart (TPRD -0.02 compared to 0.01 for the clean model), for all missingness types.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
(d) Mixed missingness (MCAR & MAR & MNAR)
Figure 5. Label Stability of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 1 are desirable. Datasets are ordered in increasing order by size. The blue line indicates performance of the model trained on clean data.
4.3. Stability of the Predictive Model
Figure 5 shows label stability of models trained using different MVI techniques. In line with conventional wisdom (Domingos, 2000), stability depends most strongly on dataset characteristics (mainly size) and model type. Impute-then-classify approaches like miss-forest, auto-ml and datawig, which were the best-performing according to F1, also match the stability of the model trained on clean data under all missingness types and dataset sizes. clustering, which as was poor-performing according to F1 (correctness) is also poor-performing according to stability on small datasets (diabetes and german, with 905 and 1k samples, respectively). A notable and surprising exception is under mixed missingness on german, where clustering is competitive with other impute-then-classify approaches, despite not being competitive on MCAR, MAR or MNAR individually.
Deletion worsens stability compared to the model trained on clean data for all missingness types on diabetes, but, notably, only under MNAR on german. boostclean, which constitutes its own model type, shows a deviation from the stability of the model trained on clean data on all datasets except bank and all missingness types: worsening stability compared to the clean model on folk-income, law-school and folk-employment (except under MCAR), but, surprisingly, improving it on heart, even under mixed missingness.
(a) F1 (imputation) vs F1 (model)
(b) RMSE (imputation) vs F1 (model)
(c) KL divergence (imputation) vs F1 (model)
Figure 6. Imputation quality vs. model performance: imputation correctness (F1, RMSE and KL divergence) may not be indicative of model correctness (F1).
4.4. Imputation Quality and Fairness
In Figure 6, we report the imputation quality of different MVI techniques according to F1 score (for categorical columns), RMSE (for numerical columns), and KL divergence (for both numerical and categorical columns, computed over the columns with nulls only) and compare it with the F1 of the downstream model.
Imputation Quality.
We find that MVI techniques with vastly different imputation quality can give rise to models with comparable F1 scores, indicating that imputation correctness is not sufficiently predictive of model correctness (Zhang and Long, 2021a). For example, in Figure 6(a) on the diabetes dataset, median-dummy has an F1 score (imputation) close to 0; clustering, datawig, median-mode and deletion have an F1 score (imputation) between 0.5-0.6, and auto-ml and miss-forest have an F1 score (imputation) close to 1, but the models trained with all of these techniques have an F1 close to 1. This trend is consistent across all datasets and missingness types, and for numerical columns as well (see Figure 6(b)).
We observe similar trends for KL divergence (see Figure 6(c)), indicating that neither discrepancy nor distributional imputation quality metrics are sufficiently indicative of downstream model performance, refuting Shadbahr et al. (2023)’s claim.
(a) F1 score difference (imputation) vs TPRD (model)
(b) RMSE difference (imputation) vs TPRD (model)
(c) KL divergence difference (imputation) vs TPRD (model)
Figure 7. Imputation fairness vs model fairness: imputation fairness (F1 difference, RMSE difference and KL divergence difference) may not be indicative of model fairness (TPRD).
Imputation fairness.
In Figure 7, we report the imputation fairness of different MVI techniques, according to F1 score difference (for categorical columns), RMSE difference (for numerical columns), and KL divergence difference (for both numerical and categorical columns, computed over the columns with nulls only) and compare it with the fairness of the downstream model, according to TPRD. We find that, while model fairness is generally agnostic to missingness type (as discussed in Section 4.2), imputation fairness is highly sensitive to missingness mechanism. For example, in Figure 7(c), miss-forest has good imputation fairness on german under MAR (KL difference of -0.4) but significant imputation unfairness under MCAR (KL difference of 2.25), MNAR (KL difference of 1.1) and mixed missingness (KL difference of 1.4).
Further, imputation fairness is insufficiently predictive of model fairness. For example, on german under mixed missingness, median-dummy has KL difference close to 0, datawig and miss-forest have KL difference between 1 and 1.5, and median-mode and clustering have KL difference between -1 and -1.5, but the models trained using all five of these techniques have a TPRD close to -0.02. Conversely, on diabetes under MAR, auto-ml, miss-forest, median-dummy and clustering all have near perfect imputation fairness (KL difference close to 0), but different model fairness (TPRD between 0.04 and 0.12). We see similar trends for other imputation fairness metrics such as F1 score difference (Figure 7(a)) and RMSE difference (Figure 7(b)).
5. Missingness Shift
Next, we evaluate the correctness, fairness, and stability of MVI techniques under missingness shift. We simulate missingness shift in two ways: (i) by varying the missingness mechanism between training and test (S4-9 in Table 1); and (ii) by varying the missingness rates between training and test. First, we hold the fraction of nulls in the test set constant (at 30%) and vary the fraction of nulls in the training set (10%, 30% and 50%). Then, we hold the fraction of nulls in the training set constant (at 30%) and vary the fraction of nulls in the test set ($10\%$, $20\%$, $30\%$, $40\%$ and $50\%$). Note that we have fewer settings for training missingness rates because varying the test set is less computationally demanding (as discussed in Section 3.4). We discuss results on diabetes, and defer results on other datasets, with fixed and variable training and test missingness rates, to Appendix D.
5.1. Correctness of the Predictive Model
Figure 8. F1 of the Random Forest model on diabetes, as a function of training set missingness rate. Dashed line shows performance of the model trained on clean data.
Training set missingness. Figure 8 shows the F1 of the Random Forest model on diabetes as a function of training missingness rate. Of all MVI techniques, deletion is most strongly affected by missingness: F1 degrades with increasing missingness rate, and this effect is strongest under MNAR. This includes when MNAR is encountered both during training (the bottom row in Figure 8 shows the steepest decline in F1 compared to other rows—training missingness) and during testing (the right-most column in Figure 8 has the lowest F1 compared to other columns —test missingness).
All other MVI techniques, including boostclean, are generally robust to higher training missingness rates, and only show a 5% decrease in F1 (compared to 10% decrease with deletion), even at rates as high as 50%. A notable exception is clustering, which, surprisingly and somewhat counter-intuitively, has higher F1 at higher training missingness rates, and is actually better under MNAR than under MAR and MCAR, for all missingness rates and scenarios.
Test set missingness.
Figure 45 in Appendix D.2 shows the F1 of the Random Forest model on diabetes as a function of test missingness rate. We find that F1 generally decreases with an increase in test missingness, corroborating the findings of  Shadbahr et al. (2023). A notable exception is miss-forest, which is robust to both forms of missingness shift (i.e., changing missingness mechanisms and missingness rates). As for training set missingness, F1 decreases with an increase in test missingness most steeply under MNAR (both during training and test). And, once again, clustering is an exception to this trend, instead showing invariance to test missingness rates under MNAR train (irrespective of test missingness) but not under MCAR and MAR train.
5.2. Fairness of the Predictive Model
Training set missingness. Figure 9 shows TPRD of Random Forest on diabetes as a function of training set missingness rate. While we previously found that fairness is generally agnostic to missingness type when it is the same between training and test sets (see Section 4.2), we find that model fairness is highly sensitive to missingness shift — in terms of both different missingness rates and different missingness mechanisms between training and test.
Worryingly, there is no consistent trend across MVI technique, missingness type and training test missingness rate. For example, consider miss-forest, which was the most robust to missingness shift according to F1. Under MCAR training, miss-forest preserves fairness of the model trained on clean data (shown with a dashed grey line) when training and test missingness rates match (at 30% training error rate, fixed at 30% for this experiment), but worsens fairness (higher TPRD) when they are different (at 10% and 50% training error rates). Under MAR training, however, we see the opposite behavior, with miss-forest preserving clean model fairness at 10% and 50% train missingness rates, but worsening fairness when training and test missingness rates are equal (at 30%). Under MNAR training, TPRD increases with an increase in training missingness rate under MCAR and MAR test, and remains constant when there is no shift in missingness mechanism (under MNAR test).
Figure 9. True Positive Rate Difference of Random Forest on diabetes, as a function of training missingness rate. Dashed line shows performance of the model trained on clean data.
Test set missingness.
We measured the impact of test missingness rate on fairness and found that fairness is highly sensitive to shifts in test missingness. Figure 46 in Appendix 45 shows the TPRD of the Random Forest model on diabetes as a function of test missingness rate. As before, fairness is highly sensitive to shifts in test missingness. For boostclean, datawig and auto-ml, TPRD generally increases (fairness worsens) with an increase in test error rate, although this increase is not monotonic. miss-forest is particularly robust, and although it only matches the fairness of the model trained on clean data under MCAR training, it is agnostic to increases in test missingness rate, under all missingness scenarios. For simpler methods such as deletion, clustering, median-mode and median-dummy, there is no consistent trend, even when missingness types remain the same and only missingness rates change between train and test. For example with deletion and clustering, TPRD increases with increase in test missingness rate in scenario S1 (MCAR train, MCAR test) and S3 (MNAR train, MNAR test), but monotonically decreases with test missingness rate in scenario S2 (MAR train, MAR test).
In summary, our results corroborate the findings of Guha et al. (2024), and are a cause for concern as they indicate that the MVI techniques that perform best during development may not preserve fairness post-deployment, where shifts are likely to occur.
5.3. Stability of the Predictive Model
Figure 10. Label Stability of Random Forest on diabetes, as a function of training set missingness rate. Dashed line shows performance of the model trained on clean data.
Training set missingness. Figure 10 shows Label Stability of the Random Forest model on diabetes as a function of training set missingness rate. Under MCAR and MAR missingness, most MVI techniques with the exception of boostclean, deletion, and clustering show good stability (comparable to the model trained on clean data), and are generally insensitive to missingness rates. deletion and clustering are the least stable methods and show a monotonic decrease in stability with increase in missingness rate, with strongest effect under MNAR. auto-ml and miss-forest are the most robust MVI techniques and preserve stability of the clean model under all missingness settings and error rates.
Test set missingness.
Figure 47 in Appendix 46 shows the Label Stability of the Random Forest model on diabetes as a function of test missingness rate. We find that missingness rate has little effect on model stability under all missingness settings, with the exception of clustering, which shows lower label stability at higher missingness rates, most pronounced in scenario S7 (MAR train, MNAR test).
6. Summary of Experimental Findings
Do not drop your nulls! Building on evidence in the scientific community (Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), our findings indicate that, in addition to hurting predictive performance, deletion also worsens stability and fairness, arguing against the use of this method in practice.
Multiple imputation shows mixed results.
There is conflicting evidence in the literature on the performance of multiple imputation (MI) (Joost R. van Ginkel and van der Voort, 2020), and our empirical findings are also mixed and somewhat unexpected.  Feng (2023) and  Le Morvan et al. (2021) argue that MI has superior predictive performance compared to impute-then-classify approaches.  Graham (2009) and McNeish (2017) find that MI performs well even with less data and small error rates. In contrast, we find that MI (specifically, boostclean) is only competitive on small datasets, and is unstable compared to simpler MVI techniques. This can likely be explained by the complexity-stability trade-offs (Domingos, 2000): MI uses a more complex model class and is able to minimize the empirical loss, but due to this very complexity it shows a large variation in predictions under small train set perturbations.
Fairness is highly missingness-specific.
We find that no MVI technique is consistently fairness-preserving, corroborating the findings of  Zhang and Long (2021b) and Guha et al. (2024). Further, we find that fairness is highly sensitive to changes in missingness rates and missingness mechanisms between training and test sets, which are likely to occur in practice, and therefore a cause for ethical concern.
Imputation quality and fairness metrics are often insufficiently predictive of the correctness and fairness of downstream models.
Shadbahr et al. (2023) posit that distributional imputation quality metrics are a more reliable indicator of a model’s predictive performance than discrepancy metrics. We find that neither discrepancy nor distributional metrics are sufficiently indicative of downstream model performance (according to F1), indicating that poor imputation quality can be overcome by a strong learner. Further, we find that imputation fairness metrics are not predictive of model fairness: perfect imputation fairness can result in downstream model unfairness. Conversely, models trained using MVI techniques with imputation unfairness can still have good downstream fairness.
(a) MCAR missingness
(b) MAR missingness
(c) MNAR missingness
(d) Mixed missingness
Figure 11. Spearman correlation ($\rho$) between MVI technique, model type, and performance metrics (F1, fairness and stability) for different train missingness mechanisms (subplots). TPRD and TNRD values close to 0 are ideal (fair), so we compute correlations using $TPRD\_Reversed=1-|TPRD|$ and $TNRD\_Reversed=1-|TNRD|$. Supplementary plots for each combination of training and test missingness are in Appendix E.
Model stability depends more on the dataset size and MVI technique than on the missingness scenario.
We find that for large datasets even simple statistical imputers can preserve stability. In contrast, for small datasets, only a few MVI techniques preserve stability, while deletion, statistical imputation and more complex ML-based approaches all worsen stability.
Sensitivity to train and test missingness rates.
The predictive performance of models (according to F1) is more sensitive to missingness rates during testing than during training. Fairness, on the other hand, is highly sensitive to both training and test missingness rates. Lastly, there is little effect of test missingness rates on model stability, but most MVI techniques are more unstable at higher training missingness rates.
No good methods for MNAR missingness.
MNAR is theoretically the hardest setting to model. Empirically, we find that MVI techniques perform poorly with MNAR during testing. Conversely, even methods that are trained on MNAR but evaluated on “simpler” test missingness (MCAR and MAR) perform poorly. Further, we find that model F1 and stability are more sensitive to missingness rates under MNAR than under MCAR and MAR. Designing MVI techniques that are effective under MNAR is an open research direction.
Deep learning outperforms tree-based models on large datasets.
There has been an explosion of interest in deep learning for tabular data (Hwang and Song, 2023; Gorishniy et al., 2021; Borisov et al., 2022; Ye et al., 2024; Fayaz et al., 2022). Recall that in Figures 3, 4, and 5 we reported results only for the best-performing models (in terms of F1), to streamline exposition. We find that tree-based methods like random forest and gradient boosted trees still outperform new deep table-learning models like GANDALF (Joseph and Raj, 2022) on small datasets (such as diabetes and german), corroborating the findings of Grinsztajn et al. (2022) and Shwartz-Ziv and Armon (2022), on tabular datasets with (simulated) missingness. GANDALF, however, is the best-performing model on larger datasets (heart and folk-employment in our study), underscoring the growing effectiveness of deep learning on a variety of data modalities.
Best performing approaches under single- and multi-mechanism missingness can be different.
A key novelty in our study is the evaluation under mixed or multi-mechanism missingness— which is a more complex but realistic missingness pattern, likely to be encountered in practice. We find that trends are different under single- and multi-mechanism missingness. Figure 11 shows that evaluation metrics and model types such as gandalf__\__clf and lgbm__\__clf are uncorrelated under single-mechanism missingness, but correlated under mixed missingness, for example for F1 ($\rho$=-0.24 for gandalf__\__clf and $\rho$=-0.37 for lgbm__\__clf) and fairness according to TPRD ($\rho$=0.55 for gandalf__\__clf and $\rho$=-0.4 for lgbm__\__clf).
It’s complicated!
Prior experimental work (Hasan et al., 2021; Lin and Tsai, 2020; Joel and Doorsamy, 2022; Shadbahr et al., 2023; García Laencina et al., 2010) concluded that there is no definitive answer to the question of which missing value imputation technique is best according to predictive performance. Our findings support this and demonstrate trade-offs between F1, fairness and stability, based on the architecture of the predictive model. For example, from Figure 11, under multi-mechanism missingness, logistic regression (lr__\__clf) is positively correlated with F1 ($\rho$=0.23) and stability ($\rho$=0.38) but negatively correlated with fairness according to TPRD ($\rho$=-0.23). Random forest (rf__\__clf), on the other hand, is weakly positively correlated with fairness according to TPRD ($\rho$=0.16) and F1 ($\rho$=0.25), but negatively correlated with stability ($\rho$=-0.24). Lastly, deep table-learning (gandalf__\__clf) is strongly positively correlated with fairness according to TPRD ($\rho$=0.55), weakly correlated with stability ($\rho$=0.12), and negatively correlated with F1 ($\rho$=-0.24). This is not bad news, but rather an acknowledgement of the inherent complexity of the task of learning from incomplete data, and a call for more rigorous and holistic evaluation protocols— of the kind demonstrated in this work — to select the most suitable missing value imputation technique and model architecture for a given task.
7. Conclusions, Limitations and Future Work
Conclusions. We presented Shades-of-Null, an evaluation suite for responsible missing value imputation. A key novelty of our work is in evaluating the predictive performance of models trained on imputed data alongside fairness and stability. Another novelty is our modeling of realistic missingness scenarios that go beyond Rubin’s classic settings of MCAR, MAR and MNAR to include multi-mechanism missingness and missingness shift. We showcased results of an extensive study, with 23,940 experimental pipelines, that uses our framework to evaluate a variety of MVI methods in realistic missingness scenarios. We synthesized interesting trends, and identified trade-offs between evaluation metrics.
Limitations.
We started with clean datasets, carefully designed meaningful missingness scenarios, and simulated them using error injectors. A limitation of our work is that we evaluate performance on synthetically generated — rather than on naturally occurring — missingness. This is a common limitation of work in this area: we often do not have a clean ground truth to compare against, and so we make the assumption that injecting synthetic errors will only exacerbate the effect of existing (unknown) errors. With this assumption, any degradation in performance we see is an underestimate of the possible degradation we can expect by comparing with an ideal clean ground truth.
Future work.
It is interesting future work to create research datasets with naturally-occurring missingness shifts. Further, while predictive performance, fairness and stability can be seen as orthogonal dimensions of performance, we have uncovered trade-offs among them that arise due to missingness in the data, the choice of a missing value imputation method, and, in some cases, the architecture of the model trained on the data post-imputation. This, coupled with the finding that there is no universally best-performing imputation method, motivates the development of holistic metrics, evaluation procedures, and imputation techniques that are grounded in the lifecycle of data collection, processing and use. That is, much exciting work remains in responsible data engineering, of which missing value imputation is an important dimension, and to which our work makes a contribution.
Appendix Overview
This supplemental material provides significantly more detail regarding the results presented in this paper. The organization is as follows: We first provide additional details on the implementation of the evaluation suite, which include a broad description of the MVI techniques used in our study (Appendix A.1) and an explanation of the corrections and enhancements made to existing benchmarks and MVI techniques to ensure the development of our truly fair evaluation suite (Appendix A.2). Next, we extend the description of our experimental settings to facilitate reproducibility (Appendix B). Specifically, Appendix B.1 provides detailed information about our experimental setup on the cluster, while Appendix B.2 extends our methodology for simulating missingness in other datasets, including proportions and base rates of protected groups, realistic missingness scenarios, feature correlation coefficients with the target, and feature importance for each dataset. In Appendix C, we offer additional results to those discussed in Section 4. In particular, Appendix LABEL:apdx:single-mechanism-imp-quality presents the imputation quality of all MVI techniques across all datasets from the perspectives of accuracy, distributional alignment, and fairness, while Appendix C.1 showcases the fairness of predictive models measured by metrics beyond the TPRD mentioned in Section 4. Finally, Appendix D presents supplementary experimental findings on the missingness shift discussed in Section 5. Specifically, Appendix D.1 examines the accuracy, stability, and fairness dimensions of model performance under missingness shift with fixed train and test error rates, highlighting its effects on various model types and datasets. Furthermore, Appendix D.2 explores further experimental results on model performance under missingness shift with variable train and test error rates.
Appendix A Additional details on the evaluation suite implementation
A.1. Missing Value Imputation Techniques
In our evaluation suite, we evaluate 8 MVI techniques, classified into 5 main categories based on the taxonomy presented in (Emmanuel et al., 2021). In each category, we selected one or two SOTA methods for our comparison. Each technique is capable of imputing both categorical and numerical features. All configurations and hyperparameters for the MVI techniques described below are specified in the configuration module within the source code.
  1. (1)
    deletion: We simply drop the rows with missingness indicators.
  2. (2)
    Statistical techniques:
    1. (a)
      median-mode: We impute missing values with the median of the complete case for numerical columns and mode (most frequently occurring value) for categorical columns.
    2. (b)
      median-dummy: We impute missing values with the median of the complete case for numerical columns and assign a (new) dummy category for nulls in categorical columns.
  3. (3)
    Machine learning approaches:
    1. (a)
      miss-forest (supervised) (Stekhoven and Bühlmann, 2011): This approach iteratively trains a RF model on a set of clean samples (no missingness) and predicts the missing values.
    2. (b)
      clustering (unsupervised) (Gajawada and Toshniwal, 2012): We assign the missing value to a cluster, based on the distance from the cluster center, and then impute the missing value with the mean or mode of the data points in that cluster. [^8] We used the k-prototypes implementation from https://github.com/nicodv/kmodes We implemented this method ourselves, as no existing solution supports null imputation based on clustering for data with both numerical and categorical features.
  4. (4)
    Deep learning approaches:
    1. (a)
      datawig (Biessmann et al., 2019): This method implements deep learning modules combined with neural architecture search and end-to-end optimization of the imputation pipeline.
    2. (b)
      auto-ml (Jäger et al., 2021): This approach uses the AutoML library autokeras (Jin et al., 2019) to implement the discriminative deep learning imputation method. For categorical columns, autokeras’ StructuredDataClassifier is employed, while StructuredDataRegressor is utilized for numerical columns. These classes manage data encoding and optimize model architecture and hyperparameters autonomously.
  5. (5)
    Multiple imputation (joint) using boostclean (Krishnan et al., 2017): This approach treats the error correction task as a statistical boosting problem where a set of weak learners are composed into a strong learner. To generate the weak learners, BoostClean iteratively selects a single imputation technique, applies it to a training set (with missingness), and fits a new model on the newly imputed training set.
A.2. Corrections and Enhancements in Benchmarks and MVI Techniques
Prior to constructing our evaluation suite, we conducted a thorough analysis of the strengths and weaknesses of existing state-of-the-art benchmarks and MVI techniques by reviewing their codebases. Unfortunately, we identified several methodological flaws and code bugs in some of them. Consequently, alongside creating our software novelty in the evaluation suite, we dedicated considerable effort to correcting these bugs in the null imputers and enhancing existing benchmarking approaches to ensure the development of a truly fair evaluation suite.
Although our evaluation suite pursued different goals, (Abdelaal et al., 2023) served as the primary benchmark against which we compared our work. The authors made significant contributions, establishing a benchmark that set a high standard in the field. However, we identified several critical concerns in their codebase that we aimed to address. Firstly, their comparison did not include hyper-parameter tuning for null imputers, and they did not utilize seeds to control randomization in the null imputers. In contrast, we meticulously tuned hyper-parameters for each null imputer in our study, implemented controlled randomization throughout the entire pipeline using seeds, and rigorously tested the evaluation suite with both unit and integration tests to ensure reproducibility. Secondly, a significant methodological flaw in their benchmark was that null imputers were fitted on the entire dataset without initially splitting it into training and test sets. The dataset split was introduced only during the model training stage, resulting in a clear leakage when null imputers were fitted on samples included in the test set. This issue likely arose because imputers like miss-forest and datawig lack built-in interfaces for fitting only on a training set and then transforming both training and test sets. Presumably, the authors chose to use these imputers without modification. In order to ensure a fair comparison of MVI techniques as they would be used in production systems, in our study, we standardized all 8 MVI techniques to meet the following requirements: 1) fitting a null imputers only on a training set and then applying it to both training and test sets; 2) enabling hyper-parameter tuning of a null imputer (if applicable); 3) controlling randomization of a null imputer using seeds (if applicable). Below is a summary of the enhancements we implemented for MVI techniques used in our evaluation suite:
  • clustering: We implemented this method ourselves to meet our imputation requirements, as existing solutions do not support null imputation based on clustering for datasets containing missingness in both numerical and categorical features at the same time. For this, we used the k-prototypes as a distance function in the imputer. [^9] https://github.com/nicodv/kmodes
  • miss-forest (Stekhoven and Bühlmann, 2011): The main concern with the original implementation was that the null imputer was fitted within a transform method. This meant that, when applying the transform method to a test set, the null imputer would be fitted on testing samples as well. To adhere to our imputation requirements, we moved the predictor fitting to a fit method and reused the fitted predictors in the transform method. Additionally, we introduced hyper-parameter tuning for the base classifier and regressor, which was lacking in the original implementation.
  • datawig (Biessmann et al., 2019): To align this technique with our imputation criteria, we modified the complete method from the
    SimpleImputer, which imputes null values across all categorical and numerical columns in the dataset. We adjusted the methodological approach in the source code and reused built-in fit and transform methods from the SimpleImputer to achieve this.
  • auto-ml: We adopted a conceptual idea from (Jäger et al., 2021) and used their code as a foundation for our implementation. A significant limitation in their source code was its capability to handle only one column with null values at a time. To enhance its functionality, we expanded the method to support imputation across multiple categorical and numerical columns. Additionally, we incorporated ideas inspired by miss-forest (Stekhoven and Bühlmann, 2011), specifically leveraging initial statistical estimates for each column containing null values to train classifiers and regressors used for imputation.
  • boostclean (Krishnan et al., 2017): We enhanced this method by introducing the capability to tune the internal hyper-parameters of boostclean used for statistical boosting. Furthermore, we implemented the ability to fine-tune the hyper-parameters of the underlying models trained on each dataset that was imputed using MVI techniques listed in boostclean.
In summary, our evaluation suite offers users not only a robust evaluation framework but also incorporates fixes to the MVI implementations previously available.
Appendix B Additional Experimental Details for Reproducibility
B.1. Computing Infrastructure
Our large-scale experimental study incorporated 8 null imputation techniques, 6 model types, 7 datasets, 10 evaluation scenarios, and 6 seeds for controlling randomization across the entire pipeline. This configuration led to an overwhelming total of 23,940 experimental pipelines. The complexity was further compounded by the necessity to tune both null imputers and ML models within each pipeline, employing a bootstrap of 50 estimators to assess model stability. Depending on the dataset size, null imputation technique, and model type, the execution time for a single experimental pipeline ranged from 3 minutes to 16.5 hours. Executing such a large number of pipelines would be impossible without a suitable experimental environment comprising a high-performance computing (HPC) cluster for execution and a 3-node MongoDB cluster for storing experimental results.
Our evaluation suite controller provides a useful separation of experimental pipelines, and the MongoDB cluster gives the ability to store results independently from a local file system. To manage the extensive computational demands, we utilized several optimizations (detailed in Section 3.4) and the SLURM software system, which allowed us to execute up to 100 simultaneous jobs on the HPC cluster. To run experimental pipelines, each SLURM job was allocated 12-24 physical cores or one GPU card, along with 16-96 GB of RAM, depending on the dataset size, imputation method, and model type. For null imputation using datawig, we employed an RTX8000 NVIDIA GPU card with 48 GB of GPU memory, while all other tasks were handled by a 2x Intel Xeon Platinum 8268 24C 205W 2.9GHz processor.
To effectively store the results of the experiments in the MongoDB cluster, we designed a data model with unique GUIDs for each experimental pipeline. This model comprehensively encompasses all facets of each experiment: dataset names, all used seeds, imputer and model hyper-parameters, and more. Since we have multiple options of input variables for each stage of the experimental pipeline, we needed to consolidate and unify the data to enable effective analysis of the results of the experiments. Hence, we established GUIDs for each layer of the experimental pipeline, encompassing identifiers for each imputation stage, model tuning stage, model profiling stage, and experimental session. This data model helps us access data on different levels, unify the structure of all experiments, increase the safety of aggregations and result interpretation, and simplify the creation of visualizations.
B.2. Extended Methodology for Simulating Missingness
Similarly to Section 3.1, this subsection outlines realistic missingness scenarios based on intrinsic dataset properties. Tables 4-10 report the demographic composition of all datasets, specifically, the proportions and base rates of each protected group. Figures 12-17 display the Spearman correlation coefficients between covariates and the target variable, as well as feature importance derived using scikit-learn’s built-in functionality for datasets not covered in Section 3.1. For datasets with numerous features, plots include only the most correlated features. These proportions, base rates, correlation coefficients, and feature importance guided the conditions of our missingness scenarios, aiming to reflect realistic instances of non-response caused by disparate access, distrust, or procedural injustice, where disadvantaged groups exhibit more missing values than privileged groups. Tables 11-16 detail the missingness scenarios for the folk-employment, folk-income, bank, heart, law-school, and german datasets, respectively.
Table 4. Table 4: Proportions and Base Rates for diabetes.
overall gender_priv gender_dis
Proportions 1.0 0.621 0.379
Base Rates 0.291 0.272 0.321
Table 5. Table 5: Proportions and Base Rates for german.
overall sex_priv sex_dis age_priv age_dis sex&age_priv sex&age_dis
Proportions 1.0 0.69 0.31 0.81 0.19 0.895 0.105
Base Rates 0.7 0.723 0.648 0.728 0.579 0.717 0.552
Table 6. Table 6: Proportions and Base Rates for folk-income.
overall sex_priv sex_dis race_priv race_dis sex&rac1p_priv sex&race_dis
Proportions 1.0 0.511 0.489 0.678 0.322 0.829 0.171
Base Rates 0.35 0.422 0.274 0.389 0.267 0.374 0.232
Table 7. Table 7: Proportions and Base Rates for law-school.
overall male_priv male_dis race_priv race_dis male&race_priv male&race_dis
Proportions 1.0 0.561 0.439 0.841 0.159 0.917 0.083
Base Rates 0.89 0.899 0.878 0.921 0.723 0.906 0.713
Table 8. Table 8: Proportions and Base Rates for bank.
overall age_priv age_dis
Proportions 1.0 0.955 0.045
Base Rates 0.117 0.106 0.341
Table 9. Table 9: Proportions and Base Rates for heart.
overall gender_priv gender_dis
Proportions 1.0 0.35 0.65
Base Rates 0.5 0.505 0.497
Table 10. Table 10: Proportions and Base Rates for folk-employment.
overall sex_priv sex_dis race_priv race_dis sex&rac1p_priv sex&race_dis
Proportions 1.0 0.490 0.510 0.625 0.375 0.806 0.194
Base Rates 0.57 0.617 0.525 0.563 0.582 0.577 0.541
Table 11. Table 11: Missingness scenarios for an error rate of 30% for folk-employment. AGEP is a numerical column; DIS, MIL, SCHL are categorical columns.
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
MCAR DIS, MIL, AGEP, SCHL N/A 0.3 0.3
MAR MIL, AGEP SEX 0.2 (female) 0.1 (male)
DIS, SCHL RAC1P 0.2 (non-white) 0.1 (white)
MNAR DIS DIS 0.25 (with disability) 0.05 (without disability)
[…]
AGEP AGEP 0.25 (¿ 50) 0.05 ($\leq 50$)
SCHL SCHL 0.25 (¡ 21) 0.05 ($\geq 21$)
(a) Correlation with label
(b) Feature importance
Figure 12. EDA for designing missingness scenarios in folk-employment.
Table 12. Table 12: Missingness scenarios for an error rate of 30% for folk-income. WKHP, AGEP are numerical columns; SCHL, MAR are categorical columns.
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
MCAR WKHP, AGEP, SCHL, MAR N/A 0.3 0.3
MAR WKHP, SCHL SEX 0.2 (female) 0.1 (male)
MAR, AGEP RAC1P 0.2 (non-white) 0.1 (white)
MNAR MAR MAR 0.25 (not married) 0.05 (married)
WKHP WKHP 0.25 (¡ 40) 0.05 ($\geq 40$)
AGEP AGEP 0.25 (¿ 50) 0.05 ($\leq 50$)
SCHL SCHL 0.25 (¡ 21) 0.05 ($\geq 21$)
(a) Correlation with label
(b) Feature importance
Figure 13. EDA for designing missingness scenarios in folk-income.
Table 13. Table 13: Missingness scenarios for an error rate of 30% for bank. balance, campaign are numerical columns; education, job are categorical columns.
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
MCAR balance, campaign, education, job N/A 0.3 0.3
MAR education, job age 0.18 ($\geq 30$) 0.12 (¡ 30)
balance, campaign marital 0.2 (single) 0.1 (married)
MNAR education education 0.2 (tertiary) 0.1 (secondary)
[…]
balance balance 0.2 ($\leq 1000$) 0.1 (¿ 1000)
campaign campaign 0.2 ($\leq 1$) 0.1 (¿ 1)
(a) Correlation with label
(b) Feature importance
Figure 14. EDA for designing missingness scenarios in bank.
Table 14. Table 14: Missingness scenarios for an error rate of 30% for heart. weight, height are numerical columns; cholesterol, gluc are categorical columns.
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
MCAR weight, height, cholesterol, gluc N/A 0.3 0.3
MAR weight, height gender 0.2 (female) 0.1 (male)
cholesterol, gluc age 0.2 ($\geq 50$) 0.1 (¡ 50)
MNAR weight weight 0.25 ($\geq 75$) 0.05 (¡ 75)
height height 0.2 (¡ 160) 0.1 ($\geq 160$)
cholesterol cholesterol 0.16 (not normal) 0.14 (normal)
gluc gluc 0.12 (not normal) 0.18 (normal)
(a) Correlation with label
(b) Feature importance
Figure 15. EDA for designing missingness scenarios in heart.
Table 15. Table 15: Missingness scenarios for an error rate of 30% for law-school. zfygpa, ugpa are numerical columns; fam_inc, tier are categorical columns.
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
MCAR zfygpa, ugpa, fam_inc, tier N/A 0.3 0.3
MAR ugpa, zfygpa male 0.2 (0) 0.1 (1)
am_inc, tier race 0.15 (non-white) 0.15 (white)
MNAR ugpa ugpa 0.2 (¡ 3.0) 0.1 ($\geq 3.0$)
zfygpa zfygpa 0.2 ($\leq 0$) 0.1 (¿ 0)
fam_inc fam_inc 0.2 (¡ 4) 0.1 ($\geq 4$)
tier tier 0.2 (¡ 4) 0.1 ($\geq 4$)
(a) Correlation with label
(b) Feature importance
Figure 16. EDA for designing missingness scenarios in law-school.
Table 16. Table 16: Missingness scenarios for an error rate of 30% for german. duration, credit-amount are numerical columns; checking-account, savings-account, employment-since are categorical columns.
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
MCAR duration, credit-amount, checking-account, savings-account, employment-since N/A 0.3 0.3
MAR savings-account, checking-account, credit-amount age 0.18 ($\leq 25$) 0.12 (¿ 25)
employment-since, duration sex 0.2 (female) 0.1 (male)
MNAR checking-account checking-account 0.25 (no account) 0.05 (not no account)
duration duration 0.25 ($\leq 20$) 0.05 (¿ 20)
savings-account savings-account 0.2 (not no savings account) 0.1 (no savings account)
[…]
credit-amount credit-amount 0.25 (¿ 5000) 0.05 ($\leq 5000$)
(a) Correlation with label
(b) Feature importance
Figure 17. EDA for designing missingness scenarios in german.
Appendix C Additional Experimental Results for Single-mechanism Missingness
C.1. Additional Accuracy and Fairness Metrics of Predictive Model
In this section, we extend the results discussed in Section 4 by presenting plots for additional popular accuracy and fairness metrics of downstream model performance. Figures 18, 19, 21, and 20 illustrate Accuracy, Disparate Impact, True Negative Rate Difference, and Selection Rate Difference of best performing models, respectively, for different imputation techniques, datasets, and missingness mechanisms. Fairness metrics are calculated for an intersectional group when multiple sensitive attributes are present in the dataset, and for a single group when there is only one sensitive attribute (see Table 2 for reference).
Overall, the plots support the assertion made in Section 4.2 that the impact of MVI on fairness (error-disparity) is most strongly influenced by the baseline disparity of the model trained on clean (complete, no missingness) data. This implies that it is hard to determine in advance whether data cleaning will worsen fairness.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 18. Accuracy of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Datasets are ordered in increasing order by size. The blue line shows performance of a model trained on clean data.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 19. Disparate Impact (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 1 are ideal/fair. Datasets are ordered in increasing order by size. The best performing model for each dataset is reported in the figure. The blue line indicates the performance of a model trained on clean data.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 20. Selection Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 0 are ideal/fair. Datasets are ordered in increasing order by size. The best performing model for each dataset is reported in the figure. The blue line indicates the performance of a model trained on clean data.
(a) Missing Completely At Random (MCAR)
(b) Missing At Random (MAR)
(c) Missing Not At Random (MNAR)
Figure 21. True Negative Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots). Values close to 0 are ideal/fair. Datasets are ordered in increasing order by size. The best performing model for each dataset is reported in the figure. The blue line indicates the performance of a model trained on clean data.
Appendix D Additional Experimental Results for Missingness Shift
D.1. Fixed Train and Test Missingness Fractions
F1 of predictive model. F1 scores of predictive models trained with various MVI techniques under different train and test missingness conditions, including missingness shift (scenarios S1-9 from Table 1), are presented in Figure 24. The plot demonstrates that F1 scores of predictive models are sensitive to missingness shifts across different datasets, as also discussed in Section 5.1. The impact of missingness shift is most significant for MCAR train & MNAR test and MAR train & MNAR test scenarios on the diabetes, german, and heart datasets. This observation aligns with the statement in (Shadbahr et al., 2023) that MNAR is more complex to model than MCAR and MAR. Additionally, miss-forest emerges as the most robust technique in handling missingness shifts across all datasets and settings.
Figure 22 further supports these findings by showing the performance of different model types in terms of F1 scores on diabetes after applying various ML and DL-based null imputers.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 22. F1 of different models (colors in legend) on the diabetes dataset for ML and DL-based MVI techniques (subplots) under missingness shift (Scenarios S1-S9).
Stability of predictive model. The stability of various top-performing models using different MVI techniques under missingness shift is illustrated in Figure 25. The plot indicates that missingness shift does not significantly impact model stability for most MVI techniques, compared to the impact observed in F1 scores. However, the plot shows that the choice of MVI technique can lead to notable differences in model stability for small-sized datasets, whereas changes in stability are less significant for datasets with more than 15K rows. This suggests that ML models exhibit better stability with larger datasets. Additionally, the plot supports our observation in Section 4.3 that the state-of-the-art multiple imputation technique boostclean shows lower model stability across different datasets compared to other MVI techniques, even if it holds good model accuracy.
Figure 23 further demonstrates the effect of model type on stability for the diabetes dataset.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 23. Stability of different models (colors in legend) on the diabetes dataset for ML and DL-based MVI techniques (subplots) under missingness shift (Scenarios S1-S9).
(a) Diabetes (905 samples)
(b) German (1000 samples)
(c) Folk Income (15,000 samples)
(d) Law School (20,798 samples)
(e) Bank (40,004 samples)
(f) Heart (70,000 samples)
(g) Folk Employment (302,640 samples)
Figure 24. F1 of best performing models for different imputation strategies (colors in legend) on different datasets (subplots) under missingness shift (Scenarios S1-S9). The blue line indicates the performance of a model trained on clean data.
(a) Diabetes (905 samples)
(b) German (1000 samples)
(c) Folk Income (15,000 samples)
(d) Law School (20,798 samples)
(e) Bank (40,004 samples)
(f) Heart (70,000 samples)
(g) Folk Employment (302,640 samples)
Figure 25. Label stability of best performing models for different imputation strategies (colors in legend) on different datasets (subplots) under missingness shift (Scenarios S1-S9). The blue line indicates the performance of a model trained on clean data.
Fairness of predictive model. The fairness of predictive models according to true positive rate difference is illustrated in Figure 26. The plot reveals that model fairness is highly sensitive to missingness shift. For instance, for the diabetes dataset, a predictive model trained on an MCAR train set can exhibit worse TPRD on an MNAR test set compared to an MCAR test set. Conversely, for the bank dataset, a predictive model trained on an MCAR train set shows better TPRD on an MNAR test set than on an MCAR test set. We further present the impact of missingness shifts on other fairness metrics, decomposed by different model types, in Figures 27-44. These plots illustrate the effects on Disparate Impact, Selection Rate Difference, and True Negative Rate Difference across all ML and DL-based imputers and all our datasets.
(a) Diabetes and Random Forest
(b) Law School and Logistic Regression
(c) Bank and Gradient Boosted Trees
Figure 26. Fairness, measured by True Positive Rate Difference, of best performing models for different imputation strategies (colors in legend) on different datasets (subplots) under missingness shift (Scenarios S1-9)
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 27. Disparate Impact of different models (colors in legend) on the diabetes dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 28. True Negative Rate Difference of different models (colors in legend) on the diabetes dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 29. Selection Rate Difference of different models (colors in legend) on the diabetes dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 30. Disparate Impact of different models (colors in legend) on the german dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 31. True Negative Rate Difference of different models (colors in legend) on the german dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 32. Selection Rate Difference of different models (colors in legend) on the german dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 33. Disparate Impact of different models (colors in legend) on the folk-income dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 34. True Negative Rate Difference of different models (colors in legend) on the folk-income dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 35. Selection Rate Difference of different models (colors in legend) on the folk-income dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 36. Disparate Impact of different models (colors in legend) on the law-school dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 37. True Negative Rate Difference of different models (colors in legend) on the law-school dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 38. Selection Rate Difference of different models (colors in legend) on the law-school dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 39. Disparate Impact of different models (colors in legend) on the bank dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 40. True Negative Rate Difference of different models (colors in legend) on the bank dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 41. Selection Rate Difference of different models (colors in legend) on the bank dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 42. Disparate Impact of different models (colors in legend) on the heart dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 43. True Negative Rate Difference of different models (colors in legend) on the heart dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
(a) AutoML
(b) Datawig
(c) K-Means Clustering
(d) MissForest
Figure 44. Selection Rate Difference of different models (colors in legend) on the heart dataset for ML and DL-based MVI techniques (subplots) under missingness shift.
D.2. Variable Train and Test Missingness Rates
Correctness under variable test set missingness. This subsection supplements the discussion of Section 5.1. Figure 45 shows effect of varying test missingness rates on the F1 of the Random Forest model on diabetes, under missingness shifts.
Figure 45. F1 of the Random Forest model on diabetes, as a function of test missingness rate, under different missingness scenarios. The dashed line indicates performance of the model trained on clean data.
Fairness (TPRD) under variable test set missingness. This subsection supplements the discussion of Section 5.2. Figure 46 shows effect of varying test missingness rates on the True Positive Rate Difference (TPRD) of the Random Forest model on diabetes, under missingness shifts.
Figure 46. True Positive Rate Difference of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
Stability under variable test set missingness. This subsection supplements the discussion of Section 5.3. It includes the label stability plots that were omitted in the main body of the paper due to space constraints. Figure 47 shows effect of varying test missingness rates on the Label Stability of the Random Forest model on diabetes, under missingness shifts.
Figure 47. Label Stability of the Random Forest model on diabetes, as a function of test missingness rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
Other fairness metrics under variable train set missingness. Figures 48, 49, and 50 for the diabetes dataset and the Random Forest model illustrate the impact of train set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively. Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the train set having 10%, 30%, or 50% error rates, while the test set maintained a constant 30% error rate across all settings.
The plots for additional fairness metrics support the assertion in Section 5.2 that null imputation methods are sensitive to the train missingness rate in terms of fairness. However, the impact varies across different fairness metrics. The highest deviation from the baseline value of each fairness metric is observed in the subfigures for MAR train sets at a 30% error rate, matching the test error rate. For instance, the effect on TNRD is similar to that on TPRD in Section 5.2, with most imputers showing an improvement in TNRD, whereas Disparate Impact and Selection Rate Difference worsen as the train error rate increases.
Other fairness metrics under variable test set missingness. Figures 51, 52, and 53 for the diabetes dataset and the Random Forest model show the impact of test set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively. Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the test set having 10%, 20%, 30%, 40%, and 50% error rates, while the train set maintained a constant 30% error rate across all settings.
The plots reinforce the assertion in Section 5.2 that model fairness is highly sensitive to missingness shift. For example, the subfigures for MCAR train & MNAR test and MAR train & MNAR test show the largest deviations with increasing test error rates for all three fairness metrics, while plots for other settings remain relatively flat. Interestingly, when null imputers and models are fitted on the MNAR train set, model performance on the MNAR test set across different test error rates do not change significantly compared to MCAR train & MNAR test and MAR train & MNAR test settings. This indicates that the missingness mechanism in the train set can significantly affect model fairness on the MNAR test set. Additionally, miss-forest and auto-ml proved to be the most robust MVI techniques against test set missingness shifts under all conditions.
Figure 48. True Negative Rate Difference of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
Figure 49. Disparate Impact of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
Figure 50. Selection Rate Difference of the Random Forest model on diabetes, as a function of train set error rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
Figure 51. True Negative Rate Difference of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
Figure 52. Disparate Impact of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
Figure 53. Selection Rate Difference of the Random Forest model on diabetes, as a function of test set error rate, under different missingness scenarios. The dashed line indicates the performance of the model trained on clean data.
Appendix E Spearman Correlation by Train and Test Missingness
In Figure 11 in Section 6, we reported correlations between MVI technique, model type, test missingness and performance metrics (F1, fairness and stability), based on train missingness. In this section, we supplement this analysis by reporting correlations between MVI technique, model type and performance metrics for each combination of train and test missingness.
Figure 54 is for MCAR train and different test missingness, Figure 55 is for MAR train and different test missingness, and Figure 56 is for MNAR train and different test missingness.
(a) MCAR train, MCAR test
(b) MCAR train, MAR test
(c) MCAR train, MNAR test
Figure 54. Spearman correlation ($\rho$) between MVI technique, model type, test missingness and performance metrics (F1, fairness and stability) under MCAR train and different test missingnesses. TPRD and TNRD values close to 0 are ideal (fair), so we compute correlations using $TPRD\_Reversed=1-|TPRD|$ and $TNRD\_Reversed=1-|TNRD|$
(a) MAR train, MCAR test
(b) MAR train, MAR test
(c) MAR train, MNAR test
Figure 55. Spearman correlation ($\rho$) between MVI technique, model type, test missingness and performance metrics (F1, fairness and stability) under MAR train and different test missingnesses. TPRD and TNRD values close to 0 are ideal (fair), so we compute correlations using $TPRD\_Reversed=1-|TPRD|$ and $TNRD\_Reversed=1-|TNRD|$
(a) MNAR train, MCAR test
(b) MNAR train, MAR test
(c) MNAR train, MNAR test
Figure 56. Spearman correlation ($\rho$) between MVI technique, model type, test missingness and performance metrics (F1, fairness and stability) under MNAR train and different test missingnesses. TPRD and TNRD values close to 0 are ideal (fair), so we compute correlations using $TPRD\_Reversed=1-|TPRD|$ and $TNRD\_Reversed=1-|TNRD|$

### v4

@@
 In this paper, we present Shades-of-Null, an evaluation suite for responsible missing value imputation.
-Our work is novel in two ways (i) we model realistic and socially-salient missingness scenarios that go beyond Rubin’s classic Missing Completely at Random (MCAR), Missing At Random (MAR) and Missing Not At Random (MNAR) settings, to include multi-mechanism missingness (when different missingness patterns co-exist in the data) and missingness shift (when the missingness mechanism changes between training and test) (ii) we evaluate imputers holistically, based on imputation quality, as well as on the predictive performance, fairness and stability of the models that are trained and tested on the data post-imputation.
-We use Shades-of-Null to conduct a large-scale empirical study involving 23,940 experimental pipelines, and find that while there is no single best-performing imputation approach for all missingness types, interesting trade-offs arise between predictive performance, fairness and stability, based on the combination of missingness scenario, imputer choice, and the architecture of the predictive model.
+Our work is novel in two ways (i) we model realistic and socially-salient missingness scenarios that go beyond Rubin’s classic Missing Completely at Random (MCAR), Missing At Random (MAR) and Missing Not At Random (MNAR) settings, to include multi-mechanism missingness (when different missingness patterns co-exist in the data) and missingness shift (when the missingness mechanism changes between training and test) (ii) we evaluate imputers holistically, based on imputation quality and imputation fairness, as well as on the predictive performance, fairness and stability of the models that are trained and tested on the data post-imputation.
+We use Shades-of-Null to conduct a large-scale empirical study involving 29,736 experimental pipelines, and find that while there is no single best-performing imputation approach for all missingness types, interesting trade-offs arise between predictive performance, fairness and stability, based on the combination of missingness scenario, imputer choice, and the architecture of the predictive model.
 We make Shades-of-Null publicly available, to enable researchers to rigorously evaluate missing value imputation methods on a wide range of metrics in plausible and socially meaningful scenarios.
@@
 Among the factors affecting data quality is data missingness, a prevailing practical challenge of sustained interest to the data management, statistics and data science communities, and to the scientific community writ large (Little and Rubin, 1987; Rubin, 1976; Pigott, 2010; Mitra et al., 2023; Zhou et al., 2024; Shadbahr et al., 2023; Graham, 2009; Le Morvan et al., 2021; Zhang and Long, 2021a; Howell, 2007; Ilyas and Chu, 2019).
-Specifically in data management, debates have been raging since the very inception of the field regarding whether — and how — to handle missing values, see  Date (1984) for a classic example.
-One difficulty, at the operational level, is that, while missing values are typically denoted with the token null, that may not always be the case.
-There could be hidden missing values in the data, like when ‘AL’ is selected as the state on a job application because it’s the first item in the drop-down list.
-Another difficulty, at the semantic level, is that a null can take on many meanings: it may signify that the value is unknown, inapplicable, intentionally withheld, etc.
-In this paper, we do not concern ourselves with this semantic debate, and we also do not consider hidden missing values (Qahtan et al., 2018).
-Rather, we focus our attention on a very specific case.
-We are given a dataset $X$ (a single relation), in which the values of some features are missing, denoted by the marker null.
-The meaning of null is that the feature does take on a value in the real world (i.e., it is applicable) but that value is unobserved in $X$.
-Further, we intend to use $X$ in a machine learning (ML) setting: to train a model or to make predictions about the tuples in $X$ using a model that has already been trained and deployed.
-This model cannot deal with null values directly, and so these values must be imputed (filled in) as part of data pre-processing.
+Debates on handling missing values in data management date back to the field’s inception, with classic discussions such as Date (1984).
+At the operational level, missing values are typically denoted by null, but hidden missing values can exist (e.g., ‘AL’ being selected by default in a job application).
+At the semantic level, null can have multiple meanings—unknown, inapplicable, or intentionally withheld.
+This paper does not engage in the semantic debate or consider hidden missing values (Qahtan et al., 2018).
+Instead, we focus on a specific case: a dataset $X$ (a single relation) where some features are missing, marked by null, indicating that the feature has a real-world value but is unobserved in $X$.
+Our goal is to use $X$ in a machine learning (ML) setting, either for model training or inference.
+Since ML models cannot handle null directly, missing values must be imputed as part of data preprocessing.
 As our starting point, we will use Rubin’s missingness framework (Rubin, 1976) that, nearly 50 years since it was proposed, still remains the most popular approach to modeling missing data.
@@
 Missing Completely at Random (MCAR)
-Consider a dataset of job applicants that includes their most recent salary and the number of years of relevant work experience.
-The MCAR assumption holds, for example, when a job applicant’s most recent salary is missing from their record for purely administrative reasons that are uncorrelated with either their (potentially missing) salary or with their (observed) experience.
-More formally, the probability that a feature value is missing is independent of $X$: $P(R|X)=P(R)$.
+In a job applicant dataset with salary and years of experience, MCAR holds if salary is missing due to administrative errors, unrelated to the salary itself or work experience. That is: $P(R|X)=P(R)$.
 Missing at Random (MAR)
-Suppose that job applicants with fewer years of relevant work experience are more likely to withhold their most recent salary (in hopes of being offered a higher salary by the prospective employer), and that this can be explained by observed covariates (i.e., by the number of years of work experience).
-The MAR assumption holds in this example, where the probability that some feature value is missing depends only on the observed covariates $X_{\text{obs}}$, and not on the missing feature values: $P(R|X)=P(R|X_{\text{obs}})$.
+If job applicants with fewer years of experience are more likely to withhold their salary, and this can be explained by observed covariates (i.e., years of experience), then MAR holds.
+Here, missingness depends only on observed features, not the missing values themselves: $P(R|X)=P(R|X_{\text{obs}})$.
 Missing Not at Random (MNAR)
-Suppose that a job applicant’s most recent salary does not correlate with their number of years of experience but rather is based on some combination of their geographic location and the results of a skills test, neither of which are captured in the data.
-Further, suppose that individuals whose most recent salary was low may be likely to withhold their salary information (again, in hopes of being offered a higher salary by the prospective employer).
-The MNAR assumption holds in this example, because missingness is correlated with the missing value itself, and it cannot be explained by the observed covariates (i.e., by the number of years of experience): $P(R|X)\not=P(R|X_{\text{obs}})$.
+Consider a job applicant whose salary depends on geographic location and skills test results—neither captured in the data—rather than years of experience.
+Suppose applicants with lower salaries are more likely to withhold this information, hoping for a higher offer.
+In this case, MNAR holds because missingness is correlated with the missing value itself and cannot be explained by observed covariates (i.e., years of experience): $P(R|X)\not=P(R|X_{\text{obs}})$.
 Missing value imputation (MVI)
-Rubin’s framework has formed the basis of a large body of technical work on missing value imputation, which has been covered in several excellent and comprehensive surveys (Abdelaal et al., 2023; Hasan et al., 2021; Le Quy et al., 2022; Mena et al., 2022; Pintz et al., 2022; Emmanuel et al., 2021; Peherstorfer et al., 2018; Lin and Tsai, 2020; Pigott, 2010; Zhou et al., 2024; Alabadla et al., 2022; García Laencina et al., 2010; Joel and Doorsamy, 2022; Ilyas and Chu, 2019).
-Broadly speaking, MVI approaches fall into several categories: (1) statistical, such as imputing missing observations with the median or mode (Schafer and Graham, 2002); (2) learning-based impute-then-classify, which use the observed data to iteratively learn and impute missing values, using k-nearest neighbors (Batista and Monard, 2003), clustering (Gajawada and Toshniwal, 2012), decision trees (Twala, 2009), or ensembles (Stekhoven and Bühlmann, 2012); and (3) joint data cleaning and model training (Karlaš et al., 2020; Krishnan et al., 2017, 2016), based on Rubin’s seminal multiple imputation framework (Rubin, 1987).
+Rubin’s framework has shaped a vast body of work on missing value imputation, extensively reviewed in several comprehensive surveys (Abdelaal et al., 2023; Hasan et al., 2021; Le Quy et al., 2022; Mena et al., 2022; Pintz et al., 2022; Emmanuel et al., 2021; Peherstorfer et al., 2018; Lin and Tsai, 2020; Pigott, 2010; Zhou et al., 2024; Alabadla et al., 2022; García Laencina et al., 2010; Joel and Doorsamy, 2022; Ilyas and Chu, 2019).
+MVI methods fall into three main categories: (1) Statistical methods, such as median or mode imputation (Schafer and Graham, 2002); (2) Learning-based impute-then-classify, which iteratively impute missing values using k-nearest neighbors (Batista and Monard, 2003), clustering (Gajawada and Toshniwal, 2012), decision trees (Twala, 2009), or ensembles (Stekhoven and Bühlmann, 2012); (3) Joint data cleaning and model training, integrating imputation with model learning (Karlaš et al., 2020; Krishnan et al., 2017, 2016), based on Rubin’s multiple imputation framework (Rubin, 1987).
 Beyond Rubin’s framework: Mixing scenarios and dealing with missingness shift.
-Rubin’s framework, while clean and amenable to analysis, does not fully capture real-world missingness.
-First, it has been argued that the assumptions behind MCAR rarely hold in practice, and that most real-world data exists on a continuum between MAR and MNAR, depending on how that data was collected (Graham, 2009).
-Second, while it is useful to distinguish between missingness scenarios based on their statistical properties to draw valid inferences, we are likely to encounter multi-mechanism missingness in practice, where several scenarios may co-exist in the same dataset, affecting subsets of its features, tuples, or both (Zhou et al., 2024).
-For example,  Mitra et al.
-(2023) introduce the notion of a data missingness life cycle and posit that data pre-processing steps such as assimilating data from different sources or from different heterogeneous populations (each with their own missingness patterns) can result in ‘structured missingness’ that cannot be represented using Rubin’s framework.
-Third, it has been observed that, in the context of data-centric AI, the assumptions that hold about missingness in the training data may no longer hold post-deployment, a phenomenon termed missingness shift (by analogy with data distribution shift) (Zhou et al., 2023).
-Evaluating MVI techniques.
-Despite an abundance of MVI techniques using different model architectures being proposed each year, making meaningful progress on learning from incomplete data ultimately comes down to how realistically we model missingness: our proposed methods are only as good as the evaluation scenarios we rigorously test them on.
-If our evaluation scenarios do not realistically capture the kinds of missingness that we encounter in practice, then our data cleaning methods are unlikely to be effective.
-Further, newly proposed MVI techniques are often only evaluated on the correctness of the imputation (imputation quality) and the correctness of the predictive model (accuracy or F1).
-However, in recent years, the scientific community is starting to uncover trade-offs between MVI efficacy and other normatively important dimensions of model performance, including algorithmic fairness (Guha et al., 2024; Feng, 2023), and stability and robustness (Zhou et al., 2024), beyond accuracy.
-Missingness as a form of pre-existing bias.
-Consider again the job applicant screening example, and suppose that the applicants’ gender and age are among the features.
-Female job applicants who suspect that they have been receiving lower salaries than their male colleagues, despite having comparable qualifications and job experience, may withhold salary information more frequently then men, hoping to narrow the gender wage gap in their next position.
-As a result there will be more missing salary values for women, and the probability that the value is missing depends on the observed covariate, gender.
-This situation is consistent with MAR missingness, and it points to pre-existing bias — a situation where data reflects a history of societal discrimination (Friedman and Nissenbaum, 1996).
-For our final example, suppose that disability status is among the features.
-Job applicants who live with a disability are more likely to omit a value for this feature.
-Assuming that this feature does not correlate with the other features, this situation is consistent with MNAR, with missingness itself serving as a proxy for disadvantage.
+Rubin’s framework, while analytically clean, does not fully capture real-world missingness.
+First, MCAR assumptions rarely hold, and real-world data often falls on a continuum between MAR and MNAR, depending on collection methods (Graham, 2009).
+Second, missingness mechanisms frequently co-exist within a dataset (affecting different features or tuples), leading to multi-mechanism missingness (Zhou et al., 2024).
+For instance, Mitra et al.
+(2023) introduce the data missingness life cycle, showing how data integration from diverse sources creates structured missingness beyond Rubin’s model.
+Third, in data-centric AI, missingness assumptions valid during training may shift post-deployment, a phenomenon termed missingness shift, analogous to data distribution shift (Zhou et al., 2023).
+Missingness as a form of bias.
+Consider the job applicant screening example with gender and age as features.
+Female applicants who suspect wage discrimination may withhold salary information more often than men, hoping to narrow the gender pay gap.
+This leads to more missing salary values for women, where missingness depends on the observed covariate (gender), aligning with MAR.
+This reflects pre-existing bias, where data encodes historical societal discrimination (Friedman and Nissenbaum, 1996).
+For another example, suppose disability status is included as a feature.
+Applicants with disabilities may be more likely to omit this information.
+If disability status is uncorrelated with other features, this scenario aligns with MNAR, with missingness itself acting as a proxy for disadvantage.
+When handling missing values, data scientists must also address technical bias (Friedman and Nissenbaum, 1996), where incorrect technical choices create disparities in predictive accuracy, often amplifying pre-existing bias.
+A key example is imputing missing values under incorrect assumptions, which can worsen disparities in classifier performance (Schelter and Stoyanovich, 2020; Stoyanovich et al., 2022; Schelter et al., 2019; Guha et al., 2024).
+For instance, if job applicant salaries are missing under MAR or MNAR (e.g., older women withhold salaries due to perceived discrimination), imputing them under MCAR could further depress salary estimates, reinforcing the gender wage gap and ageism, and leading to discriminatory outcomes.
+Missing value imputation can impact model arbitrariness.
+Missingness is an indication of uncertainty in the data.
+MVI methods “resolve” this uncertainty at the tuple level, but they may induce a change in the data distribution in ways that impacts the stability of predictions of a model trained on this data.
+In some cases, the resulting models produce vastly different — and even arbitrary — predictions under small perturbations in the input (Coston et al., 2023; Creel and Hellman, 2022; Rhea et al., 2022b, a).
+For example, if a job applicant’s salary is imputed in vastly different ways upon two consecutive applications for the same position, and this, in turn, impacts the hiring decision, then the decision-making process violates the principle of process fairness (e.g., (sta, 1983; Tyler, 2006)).
+Importantly, instability and accuracy are orthogonal: models can be accurate in expectation while still being unstable (Long et al., 2023).
+Research gap
+Despite numerous MVI techniques being proposed each year, there has been limited systematic progress in assessing them across key performance aspects, including imputation correctness, predictive accuracy, and fairness—measured as disparities in imputation quality or model performance across groups.
+Moreover, while missingness signals uncertainty, there has been no comprehensive evaluation of the stability of models trained on cleaned data.
+Crucially, realistic modeling of missingness, identifying bias sources, and selecting appropriate stakeholder groups and fairness metrics must be grounded in the specific context of use (Mitchell et al., 2021; Fuster et al., 2022; Obermeyer et al., 2019; Köchling and Wehner, 2020).
+For instance, age-based discrimination is relevant in both hiring and lending, yet older applicants face disadvantages (and legal protections) in hiring, while younger applicants are disadvantaged in lending.
+Thus, MVI techniques must be evaluated in societally meaningful scenarios.
 Summary of contributions
 We implemented an experimental benchmark called Shades-of-Null to rigorously and comprehensively evaluate state-of-the-art MVI techniques on a variety of realistic missingness scenarios (including single- and multiple-mechanism missingness and missingness shift), on a suite of evaluation metrics (including fairness and stability), in the context of data pre-processing in a machine learning pipeline.
-Our work is (1) novel: to the best of our knowledge, the settings of multi-mechanism missingness and missingness shifts have not been empirically studied before; (2) comprehensive: we evaluate a suite of 8 MVI techniques on 7 benchmark datasets using 6 model types, running a total of 23,940 pipelines, and is the first study of such scale in the missing data domain, to the best of our knowledge; (3) normatively grounded: we focus on decision-making contexts such as lending, hiring, and healthcare, where missingness is socially salient.
+Our work is (1) novel: to the best of our knowledge, the settings of multi-mechanism missingness and missingness shifts have not been empirically studied before; (2) comprehensive: we evaluate a suite of 15 MVI techniques on 7 benchmark datasets using 6 model types, running a total of 29,736 pipelines, and is the first study of such scale in the missing data domain, to the best of our knowledge; (3) normatively grounded: we focus on decision-making contexts such as lending, hiring, and healthcare, where missingness is socially salient.
 Mitigating social harm such as algorithmic discrimination is a leading concern in these domains (Barocas and Selbst, 2016), and we evaluate the impact of MVI approaches on downstream model fairness and stability (which have been understudied in the context of missing data), in addition to classically studied imputation quality and model correctness metrics.
@@
 See Appendix A.2 for details.
-We make Shades-of-Null publicly available [^1] https://github.com/FalaahArifKhan/data-cleaning-stability and hope to enable researchers to comprehensively and rigorously evaluate new MVI methods on a wide range of evaluation metrics, under plausible and socially meaningful missingness scenarios.
+We make Shades-of-Null publicly available  [^1] https://github.com/FalaahArifKhan/data-cleaning-stability and hope to enable researchers to comprehensively evaluate new MVI methods on a wide range of evaluation metrics, under plausible and socially meaningful missingness scenarios.
 2. Related Work
@@
 (2023) review deep learning-based approaches (variational auto-encoders and generative adversarial networks) and representation learning (graph neural networks and diffusion-based methods).
-Multiple imputation (Schafer and Graham, 2002; Zhou et al., 2024) and expectation maximization (Pigott, 2010; Wang et al., 2024) are also influential, but too computationally expensive to be popular in practice (Alabadla et al., 2022).
+Multiple imputation (Schafer and Graham, 2002; Zhou et al., 2024) and expectation maximization (Pigott, 2010; Wang et al., 2024a) are also influential, but too computationally expensive to be popular in practice (Alabadla et al., 2022).
+MNAR-specific techniques, like not-MIWAE (Ipsen et al., 2020) and GINA (Ma and Zhang, 2021), tackle the challenge of MNAR data by employing identifiable generative models that effectively account for complex missingness mechanisms.
+Recent methods, including NOMI (Wang et al., 2024b) and TDM (Zhao et al., 2023), introduce advancements like uncertainty-driven networks and transformed distribution matching, which enhance both imputation accuracy and computational efficiency.
 Beyond impute-then-classify, the data management community has proposed holistic methods like CPClean (Karlaš et al., 2020) and ActiveClean (Krishnan et al., 2016), that jointly perform data cleaning and model training, deriving from the multiple imputation framework (Rubin, 1987).
@@
 Evaluating MVI techniques
-Empirical studies have been overwhelmingly focused on medical datasets, and only evaluate missingness under MCAR (Lin and Tsai, 2020; Alabadla et al., 2022; Farhangfar et al., 2008).
-Further, most proposed methods only perform direct evaluation, using metrics such as MAE, MSE, RMSE, and AUC (Hasan et al., 2021; Joel and Doorsamy, 2022).
-Very few studies use both direct and indirect evaluation, and even studies that include indirect evaluation only evaluate predictive accuracy (Lin and Tsai, 2020).
-This leaves a significant research gap, specially for desiderata such as stability and self-consistency, which have recently been shown to be highly sensitive to data-centric interventions (Black and Fredrikson, 2021; Black et al., 2022; Cooper et al., 2023), and unfairness, which can be worsened by improperly chosen data cleaning techniques (Guha et al., 2024).
-Additionally, the performance of MVI techniques under multi-mechanism missingness (Zhou et al., 2024) and missingness shifts (Zhou et al., 2023) remains unexplored, despite these conditions being more likely to occur in practice due to distribution shifts in production deployments (Graham, 2009).
-Notably, and worryingly, overwhelming evidence in the literature indicates that there is no single “best-performing” MVI approach on accuracy  (Hasan et al., 2021; Lin and Tsai, 2020; Shadbahr et al., 2023; García Laencina et al., 2010; Farhangfar et al., 2008), and that model performance (narrowly measured based on ‘correctness’ thus far) depends on dataset characteristics such as size and correlation between variables (Alabadla et al., 2022) and missingness rates in the train and test sets (Lin and Tsai, 2020; Shadbahr et al., 2023).
+We are aware of several surveys of MVI techniques, all conducted with a strong empirical focus (Lin and Tsai, 2020; Alabadla et al., 2022; Farhangfar et al., 2008; Miao et al., 2022; Shahbazian and Greco, 2023).
+Miao et al.
+(2022) compare 19 MVI methods on 15 datasets, and while our results corroborate their findings (see Section 5.1), their evaluation is limited to imputation quality and overall accuracy (but not fairness or stability).
+Other empirical studies have been primarily focused on medical datasets, and only evaluate missingness under MCAR (Lin and Tsai, 2020; Alabadla et al., 2022; Farhangfar et al., 2008; Shahbazian and Greco, 2023).
+Further, most proposed methods only evaluate imputation quality, using metrics such as MAE, MSE, RMSE, and AUC (Hasan et al., 2021; Joel and Doorsamy, 2022), although some also evaluate overall predictor accuracy  (Lin and Tsai, 2020).
+Additionally, the performance of MVI techniques under multi-mechanism missingness (Zhou et al., 2024) and missingness shifts (Zhou et al., 2023) remains unexplored in prior work, despite these conditions being more likely to occur in practice due to distribution shifts in production deployments (Graham, 2009).
+Notably, overwhelming evidence in the literature indicates that there is no single “best-performing” MVI approach on accuracy  (Hasan et al., 2021; Lin and Tsai, 2020; Shadbahr et al., 2023; García Laencina et al., 2010; Farhangfar et al., 2008), and that model performance (narrowly measured based on ‘correctness’ thus far) depends on dataset characteristics such as size and correlation between variables (Alabadla et al., 2022) and missingness rates in the train and test sets (Lin and Tsai, 2020; Shadbahr et al., 2023).
 Fairness and missingness
-There has been some recent interest in studying the social harm that can come from poorly chosen MVI techniques (Caton et al., 2022; Zhang and Long, 2022, 2021a, 2021b; Jeong et al., 2022; Wang and Singh, 2021; Martínez-Plumed et al., 2019; Feng et al., 2024).
-Most empirical studies (Wang and Singh, 2021; Zhang and Long, 2021b, 2022; Caton et al., 2022) have worked with the COMPAS (Larson et al., 2016) and Adult  [^2] https://archive.ics.uci.edu/dataset/2/adult datasets, the latter of which has been “retired” from community use due to issues with provenance (Ding et al., 2021).
+There has been some recent interest in studying the social harm that can come from poorly chosen MVI techniques (Caton et al., 2022; Zhang and Long, 2022, 2021a, 2021b; Jeong et al., 2022a; Wang and Singh, 2021; Martínez-Plumed et al., 2019; Feng et al., 2024).
+Most empirical studies (Wang and Singh, 2021; Zhang and Long, 2021b, 2022; Caton et al., 2022; Jeong et al., 2022b) have worked with the COMPAS (Larson et al., 2016) and Adult (Dua and Graff, 2019) datasets, the latter of which has been “retired” from community use due to issues with provenance (Ding et al., 2021).
 Further, these experimental studies employ randomly-generated missingness: usually by randomly sampling or using a fixed set of columns, and randomly picking rows in which to replace values with null.
@@
 A notable exception is Martínez-Plumed et al.
-(2019), who map social mechanisms such as prejudicial access and self-reporting bias to different categories of missingness such as missing-by-design and item non-response.
-Further, they analyse the correlations between all features to study the effect of different kinds of missingness.
-We use a similar methdology to simulate realistic missingnesses in this work.
-Importantly, however, we find conceptual limitations in their framing of algorithmic fairness.
-Martínez-Plumed et al.
-(2019) write:“The surprising result was to find that, […] the examples with missing values seem to be fairer than the rest.” Asserting that some rows of data are more or less fair does not make sense, in that samples (e.g., job applicants) do not demonstrate the property of being fair.
-It is the model (e.g., used to make hiring decisions) that asserts the property of being fair or not, by the inclusion or exclusion of certain samples amongst those given a positive outcome.
-We re-interpret the findings of  Martínez-Plumed et al.
-(2019) to say that excluding samples with missingness can increase model unfairness, which is further evidence against the use of deletion as a strategy for dealing with missing data.
-Zhang and Long (2021b) evaluate MVI methods on imputation fairness, defined as the imputation accuracy difference on privileged and disadvantaged demographic groups.
-They find that (1) higher missingness disparity leads to higher imputation unfairness; (2) higher rates of missingness lead to higher imputation unfairness; (3) more imbalance in the data (or larger disparity in the sample sizes across groups) leads to higher imputation unfairness; (4) the same imputation method has different prediction fairness under different missingness mechanisms.
-Their evaluation is limited to randomly generated null values in COMPAS.
-We extend their analysis of imputation fairness to additional datasets, missingness scenarios and imputation fairness definitions (such as KL divergence difference).
-Zhang and Long (2022) introduce the notion of imputation fairness risk and provide bounds for “correctly specified” imputation methods.
-While there is very little theoretical work in this space, and this is a commendable contribution, we call into question the broader implications of this result: it is well-known that imputation quality metrics do not fully capture downstream model performance (Zhang and Long, 2021a).
-Put differently, a classifier built on poor imputation quality can still give good downstream performance (Shadbahr et al., 2023).
-With this in mind, we ask the question: Does minimizing imputation unfairness minimize model unfairness?
-Our empirical findings indicate it does not, as is discussed in Section 4.4.
-Missingness and stability. We are not aware of any work investigating the effect of missing value imputation on model stability.
+(2019), who map social mechanisms such as prejudicial access and self-reporting bias to missingness categories like missing-by-design and item non-response.
+They also analyze feature correlations to study the effects of different missingness types.
+We adopt a similar methodology to simulate realistic missingness in this work but identify conceptual limitations in their fairness framing.
+The authors state: “The surprising result was to find that, […] the examples with missing values seem to be fairer than the rest.” However, asserting that some rows of data are more or less fair is misguided, as fairness is not a property of individual samples (e.g., job applicants) but of the model (e.g., in hiring decisions), which determines fairness through inclusion or exclusion in positive outcomes.
+We reinterpret their findings to suggest that excluding samples with missing values can increase model unfairness, reinforcing the case against deletion as a missing data strategy.
+Zhang and Long (2021b) evaluate MVI methods on imputation fairness, defined as the difference in imputation accuracy between privileged and disadvantaged groups.
+They find that imputation unfairness increases with higher missingness disparity, higher overall missingness rates, and greater data imbalance across groups.
+Further, they find that varying missingness mechanisms for the same imputation method impacts prediction fairness.
+Their analysis is limited to randomly generated null values in COMPAS.
+We extend their work to additional datasets, missingness scenarios, and alternative imputation fairness definitions.
+In a follow-up work, Zhang and Long (2022) introduce imputation fairness risk and provide bounds for “correctly specified” imputation methods.
+While this is a commendable theoretical contribution in a largely unexplored area, we question its broader implications: imputation quality metrics do not fully capture downstream model performance (Zhang and Long, 2021a).
+In other words, a classifier can perform well despite poor imputation quality (Shadbahr et al., 2023).
+This raises a key question: Does minimizing imputation unfairness reduce model unfairness?
+Our empirical findings suggest it does not, as discussed in Section 4.4.
+Finally, Jeong et al.
+(2022b) propose a decision tree-based method that integrates fairness into model training while handling missing values.
+Their approach splits only on observed values to mitigate disparities introduced by imputation.
+Their evaluation is limited to MCAR scenarios (with more missingness for disadvantaged groups).
+In contrast, we assess more advanced MVI techniques under diverse missingness scenarios (MCAR, MAR, MNAR, and missingness shift) without applying fairness interventions.
+Nonetheless, we share their broader motivation of assessing and mitigating unfairness holistically throughout the data lifecycle.
+Future work could explore different combinations of MVI and fairness interventions.
+Missingness and stability.
+We are not aware of any work investigating the effect of missing value imputation on model stability.
 3. Benchmark Overview
@@
 Our methodology for simulating missingness is based on evaluation scenarios, defined by the missingness mechanism during training and testing, shown in Table 1: (1) single-mechanism missingness, injected similarly into train and test sets (S1 - S3); (2) single-mechanism missingness, injected differently into train and test sets (missingness shift) (S4 - S9); and (3) missingness is mixed, to include all three missingness mechanisms, and is injected similarly into train and test sets (S10).
+For each dataset in our study (see Section 3.5), we designed socially-salient missingness scenarios corresponding to the three missingness mechanisms (MCAR, MAR, MNAR).
+Following (Joel and Doorsamy, 2022; Martínez-Plumed et al., 2019), we identified features for missing value injection, denoted by $\mathcal{F}^{m}$, based on their Spearman correlation with the target variable and feature importance scores computed using scikit-learn.
+These selected features were chosen to reflect plausible missingness patterns.
+For instance, in the diabetes dataset, while features like blood pressure or cholesterol levels are expected to be consistently observed, others, such as family history or physical activity, might be omitted or withheld due to privacy concerns or reporting biases.
+The remaining features, denoted by $\mathcal{F}^{c}$, were considered complete, with no missing values.
+The three missing mechanisms share the same set of selected features ($\mathcal{F}^{m}$), but differ in their injection strategies.
+For MCAR, the missing values are randomly injected on $\mathcal{F}^{m}$.
+In contrast, the missingness of MAR is based on sensitive attributes within $\mathcal{F}^{c}$ to simulate pre-existing bias, as described in Section 1.
+Specifically, higher rates of missingness were injected to disadvantaged groups wherever possible (in some cases there were too few samples from disadvantaged groups), reflecting realistic disparities caused by unequal access, distrust, or procedural injustice (Akpinar et al., 2024).
+Finally, for MNAR, the missingness is determined by missing values themselves, and the likelihood of missing values depends on the missing features.
+Table 3 presents the selected columns ($\mathcal{F}^{m}$) and injection conditions for the diabetes dataset, based on the correlation coefficients and feature importances in Figure 2.
+Additional information on other datasets is available in Appendix B.2.
+3.2. Missing Value Imputation (MVI) Techniques
+As discussed in Section 2, many competitive MVI techniques have been proposed.
+We selected 15 of them, from 8 broad categories based on taxonomies presented in (Emmanuel et al., 2021; Jäger et al., 2021; Miao et al., 2022; Wang et al., 2024b), namely: (1) deletion; (2) statistical: median-mode and median-dummy; (3) machine learning-based: miss-forest (Stekhoven and Bühlmann, 2011) and clustering (Gajawada and Toshniwal, 2012); (4) discriminative deep learning-based: datawig (Biessmann et al., 2019) and auto-ml (Jäger et al., 2021); (5) generative deep learning-based: gain (Yoon et al., 2018) and hi-vae (Nazabal et al., 2020); (6) MNAR-specific: not-miwae (Ipsen et al., 2020) and mnar-pvae (Ma and Zhang, 2021); (7) multiple imputation: boostclean (Krishnan et al., 2017); and (8) other recent: nomi (Wang et al., 2024b), tdm (Zhao et al., 2023), and edit-gain (Miao et al., 2021).
+See Appendix A.1 for details.
+for details.
 Table 1. Table 1: Evaluation Scenarios
@@
 S10
-For each dataset in our study (see Section 3.5), we designed socially-salient missingness scenarios corresponding to the three missingness mechanisms (MCAR, MAR, MNAR).
-Following (Joel and Doorsamy, 2022; Martínez-Plumed et al., 2019), we selected columns in which to inject null values based on the correlation coefficients between all covariates and the target variable, as well as feature importance computed using scikit-learn’s built-in functionality.
-We used sensitive attributes in the design of MAR missingness to simulate pre-existing bias, as described in Section 1, and injected higher rates of missingness for disadvantaged groups wherever possible (in some cases there were too few samples from disadvantaged groups), aiming to reflect realistic instances of disparate non-response caused by disparate access, distrust, or procedural injustice (Akpinar et al., 2024).
-Furthermore, we selected columns that may plausibly contain missingness.
-For example, blood pressure or cholesterol in diabetes is expected to always be observed, but family history or physical activity might be omitted or withheld.
-Table 3 presents missingness scenarios for the diabetes dataset, based on the correlation coefficients and feature importances in Figure 2. Details for other datasets are available in Appendix B.2.
-3.2. Missing Value Imputation (MVI) Techniques
-As discussed in Section 2, there is a large number of competitive MVI techniques in the literature.
-For our study, we picked 8 techniques, from 5 broad categories in the taxonomy of  Emmanuel et al.
-(2021), namely: (1) deletion; (2) statistical techniques: median-mode and median-dummy; (3) machine learning-based techniques: miss-forest (Stekhoven and Bühlmann, 2011) and clustering (Gajawada and Toshniwal, 2012); (4) deep learning-based techniques: datawig (Biessmann et al., 2019) and auto-ml (Jäger et al., 2021); and (5) multiple imputation: boostclean (Krishnan et al., 2017).
-Additional details about these methods are deferred to Appendix A.1.
 3.3. Evaluation Metrics
@@
 $DI=\frac{TP_{dis}+FP_{dis}}{N_{dis}}\div\frac{TP_{priv}+FP_{priv}}{N_{priv}}$
+Broadly speaking, fairness metrics that are based on error rates are consistent with formal equality of opportunity, while those that are based on selection rates (SRD and DI) are consistent with substantive equality of opportunity (Khan et al., 2022).
+Which specific metric is more relevant depends on the application domain and on the stakeholder (Narayanan, 2018).
+SRD focuses on absolute disparities in selection rates (e.g., fixed quotas used in college admissions), while DI focuses on relative disparities (e.g., enforcing the 4/5th rule for non-discrimination in hiring in the U.S.).
+From the point of view of individuals being selected, TPRD is typically used when assessing fairness in opportunity allocation, such as hiring or loan approvals, where a positive outcome should be equally accessible across groups, while TNRD is most relevant in risk assessment and exclusion decisions, like medical diagnoses, where fairness in avoiding false negatives (e.g., failing to detect a disease) matters.
 3.4. Shades-of-Null Architecture
 The architecture of Shades-of-Null is shown in Figure 1.
-Our goal when developing this evaluation suite was to support the comprehensive evaluation of different MVI techniques (Section 3.2), under different missingness scenarios (Section 3.1), according to different aspects of model performance (Section 3.3).
-Shades-of-Null is novel in several ways.
-First, it is flexible and supports a variety of missingness scenarios, namely, single-mechanism, multi-mechanism and missingness shift.
-Second, it supports holistic evaluation of both imputation quality (including imputation fairness) and model performance — including accuracy, stability and fairness.
-The central component of the architecture in Figure 1 is the benchmark controller that executes user-specified missingness scenarios by applying error injectors to input datasets.
-It then imputes missing values according to user-specified MVI technique(s) and pre-processes the imputed datasets using standard scaling for numerical features and one-hot encoding for categorical features.
-The controller then executes ML model training, including hyperparameter tuning.
-Finally, the evaluation module assesses imputation quality and model performance.
-For comprehensive model profiling, the evaluation module uses Virny (Herasymuk et al., 2024), a Python library that computes accuracy, stability and fairness metrics for a set of initialized models, for multiple sensitive attributes and their intersections.
-We store experimental results in MongoDB.
+The core component is the benchmark controller, which executes user-specified missingness scenarios by applying error injectors to input datasets.
+It then imputes missing values using the selected MVI technique(s) and preprocesses the data with standard scaling for numerical features and one-hot encoding for categorical features.
+Next, the controller trains ML models, including hyperparameter tuning.
+The evaluation module then assesses imputation quality and model performance.
+For comprehensive profiling, it uses Virny (Herasymuk et al., 2024), a Python library that computes accuracy, stability, and fairness metrics across multiple sensitive attributes and their intersections.
 Figure 1. Shades-of-Null architecture
 Table 2. Table 2: Dataset Information
-Name Domain Size Sensitive Attributes
-diabetes healthcare 905 sex
-german finance 1,000 sex, age
-folk-income finance 15,000 sex, race
-law-school education 20,798 sex, race
-bank marketing 40,004 age
-heart healthcare 70,000 sex
-folk-employment hiring 302,640 sex, race
-Shades-of-Null incorporates two optimizations to streamline the execution of multiple experiments.
-Firstly, it decouples missing value imputation and model training.
-This enables the preservation of imputation results, including imputed datasets stored in a file system, facilitating their reuse during subsequent model training and evaluation stages.
-Secondly, the experimental pipeline within Shades-of-Null supports evaluation on multiple test sets simultaneously (e.g., with different rates or types of missingness), to speed up evaluation.
-For example, the running time of executing a pipeline with one training set and one test set is close to that of a pipeline with one train set and multiple test sets.
+name domain # tuples # attrs sensitive attrs
+diabetes healthcare 952 17 sex
+german finance 1,000 21 sex, age
+folk-income finance 15,000 10 sex, race
+law-school education 20,798 11 sex, race
+bank marketing 40,004 13 age
+heart healthcare 70,000 11 sex
+folk-employment hiring 302,640 16 sex, race
+Shades-of-Null incorporates two optimizations to enhance experimental efficiency.
+First, it decouples missing value imputation from model training, allowing imputed datasets to be stored and reused in subsequent training and evaluation stages.
+Second, it supports simultaneous evaluation on multiple test sets (e.g., with varying missingness rates or types), significantly reducing running time, and so executing a pipeline with one training set and multiple test sets takes about the same time as with a single test set.
 3.5. Datasets and Tasks
-As noted in Section 1, in this work we are interested in missingness that is socially salient.
-With this in mind, we selected $7$ datasets from a variety of social decision-making contexts including lending, hiring, marketing, admissions, and healthcare, summarized in Table 2.
-For each dataset, there is a binary classification task, where being assigned a positive label corresponds to the attainment of a desirable social good (such as education, employment, or access to healthcare).
-We selected these datasets to have good coverage of (i) social domains and (ii) dataset sizes and other characteristics.
+As noted in Section 1, we focus on socially salient missingness.
+With this in mind, we selected seven datasets from diverse social decision-making contexts, including lending, hiring, marketing, admissions, and healthcare, summarized in Table 2.
+Each dataset involves a binary classification task, where a positive label represents access to a desirable social good (e.g., education, employment, or healthcare).
+We chose these datasets to ensure broad coverage of (i) social domains, (ii) dataset sizes, and (iii) numerical-to-categorical column ratios.
+Dataset descriptions are deferred to Appendix A.3
 Table 3. Table 3: Missingness scenarios for an error rate of 30% for diabetes. SoundSleep is a numerical column; Family__\__Diabetes, PhysicallyActive and RegularMedicine are categorical columns.
-Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
+[…]
 MCAR SoundSleep, Family__\__Diabetes, PhysicallyActive, RegularMedicine N/A 0.3 0.3
@@
 Figure 2. EDA for designing missingness scenarios in diabetes.
-diabetes [^3] https://www.kaggle.com/datasets/tigganeha4/diabetes-dataset-2019 (Tigga and Garg, 2020) was collected in India through a questionnaire including 18 questions related to health, lifestyle, and family background.
-A total of 952 participants are characterized by 17 attributes (13 categorical, 4 numerical) and a binary target variable that represents whether a person is diabetic.
-Here, sex is the sensitive attribute, with “female” as the disadvantaged group.
-german [^4] https://archive.ics.uci.edu/dataset/144/statlog+german+credit+data (Hofmann, 1994) is a popular fairness dataset that contains records of creditworthiness assessments, classifying individuals as high or low credit risks.
-It contains information on 1,000 individuals characterized by 21 attributes (14 categorical, 7 numerical), including credit history, occupation and housing information.
-Here, sex and age are the sensitive attributes, with “female” and “age25absent25$\leq$ 25≤ 25” as the disadvantaged groups.
-Folktables (Ding et al., 2021) is another popular fairness dataset derived from US Census data from all 50 states between 2014-2018.
-The dataset has several associated tasks, of which we selected two: (i) ACSIncome (folk-income) is a binary classification task to predict whether an individual’s annual income is above \$50,000, from 10 features (8 categorical, 2 numerical) including educational attainment, work hours per week, marital status, and occupation.
-We use data from Georgia from 2018, subsampled to 15k rows.
-(ii) ACSEmployment (folk-employment) is a binary classification task to predict whether an individual is employed, from 16 features (15 categorical, 1 numerical) including educational attainment, employment status of parent, military status, and nativity.
-We use data from California from 2018, with 302,640 rows (we do not subsample).
-In both tasks, sex and race are the sensitive attributes, with “female” and “non-White” as the disadvantaged groups.
-law-school [^5] https://www.kaggle.com/datasets/danofer/law-school-admissions-bar-passage  (Wightman, 1998) was gathered through a survey conducted by the Law School Admission Council (LSAC) across 163 law schools in the US in 1991, and contains admissions records of 20,798 applicants, characterized by 11 attributes (5 categorical, 6 numerical), including LSAT scores and college GPAs.
-The task is to predict whether a candidate would pass the bar exam.
-Here, sex and race are the sensitive attributes, with “female” and “non-White” as the disadvantaged groups.
-bank [^6] https://archive.ics.uci.edu/dataset/222/bank+marketing  (Moro et al., 2014) contains data from direct marketing campaigns by a Portuguese bank, between 2008 and 2013.
-It contains information on 40,004 potential customers, with 16 attributes (10 categorical, 6 numerical), including occupation, marital status, education, and a binary target that indicates whether the individual subscribed for a term deposit.
-Here, age is the sensitive attribute, with $<25$ and $>60$ as the disadvantaged group.
-heart [^7] https://www.kaggle.com/datasets/sulianova/cardiovascular-disease-dataset consists of patient measurements with respect to cardiovascular diseases, with information of 70,000 individuals characterized by 11 attributes (6 categorical, 5 numerical), including age, height, weight, blood pressure, and a binary target indicating whether the patient has heart disease.
-Here, sex is the sensitive attribute, with “female” as the disadvantaged group.
 3.6. Model Types
 We evaluate predictive performance of 6 ML models: (i) decision tree (dt__\__clf) with a tuned maximum tree depth, minimum samples at a leaf node, number of features used to decide the best split, and criteria to measure the quality of a split; (ii) logistic regression (lr__\__clf) with tuned regularization penalty, regularization strength, and optimization algorithm; (iii) gradient boosted trees (lgbm__\__clf) with tuned number of boosted trees, maximum tree depth, maximum tree leaves, and minimum number of samples in a leaf; (iv) random forest (rf__\__clf) with a tuned number of trees, maximum tree depth, minimum samples required to split a node, and minimum samples at a leaf node (v) neural network, historically called the multi-layer perceptron (mlp__\__clf) with two hidden layers, each with 100 neurons, and a tuned activation function, optimization algorithm, and learning rate; (vi) a deep table-learning method called GANDALF (Joseph and Raj, 2022) (gandalf__\__clf) with a tuned learning rate, number of layers in the feature abstraction layer, dropout rate for the feature abstraction layer, and initial percentage of features to be selected in each Gated Feature Learning Unit (GFLU) stage.
+Search grids of hyperparameters for all models are defined in our codebase.
 4. Single and Multi-Mechanism Missingness
@@
 To evaluate model correctness, we report results for F1, see Appendix C.2 for accuracy results.
-For fairness, we use binary group definitions throughout.
+For fairness, we use binary group definitions.
 For datasets with two sensitive attributes, we define the doubly-disadvantaged group as disadvantaged (dis) and everyone else as privileged (priv).
-For example: on the law-school, folk-income and folk-employment datasets, non-White women are the dis group, and White women, non-White men and White men collectively form the priv group.
+For example: on the law-school, folk-income and folk-employment datasets, non-White women are the dis group, and White women, non-White men and White men are the priv group.
 We report results for TPRD, and defer results for other fairness metrics to Appendix C.2.
@@
 We discuss this unexpected performance of MI further in Section 7.
-auto-ml, datawig and miss-forest are generally the best performing MVI techniques, but simpler statistical techniques (e.g., median-mode under MAR) are also competitive.
-This underscores the need to evaluate novel DL-based and ML-based approaches more holistically (e.g., on a variety of missingness scenarios) to ensure that they justify the additional training overhead and complexity they introduce, compared to simple statistical methods.
+auto-ml, datawig and miss-forest are generally the best performing MVI techniques, with nomi a close second, offering an optimal balance between imputation accuracy and training time (see Section 6 and Appendix C.1 for training time analysis).
+However, simpler statistical techniques (e.g., median-mode under MAR) are also competitive.
+This underscores the need to evaluate novel DL-based and ML-based methods holistically (e.g., on a variety of missingness scenarios) to ensure that they justify the additional training overhead and complexity they introduce compared to simple methods.
+Interestingly, not-miwae and mnar-pvae do not demonstrate superior performance compared to other methods in our socially salient MNAR scenario.
+Instead, their performance aligns closely with other leading MVI approaches under MNAR conditions.
+This finding is further discussed in Section 7.
 In line with conventional wisdom (Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), we find that deletion worsens predictive performance.
 This effect is strongest for small datasets like diabetes, with F1 decreasing as much as 0.1 under MNAR, compared to the model trained on clean data.
+This is due to deletion discarding useful information, whereas retaining rows with nulls can still provide valuable signal for model training.
 Note that the F1 score on the bank dataset is extremely low (0.32).
@@
 Impute-then-classify approaches like miss-forest, auto-ml and datawig, which were the best-performing according to F1, also match the stability of the model trained on clean data under all missingness types and dataset sizes.
+nomi, which performs best on accuracy and training time, demonstrates comparable stability to these top techniques on all datasets.
 clustering, which as was poor-performing according to F1 (correctness) is also poor-performing according to stability on small datasets (diabetes and german, with 905 and 1k samples, respectively).
 A notable and surprising exception is under mixed missingness on german, where clustering is competitive with other impute-then-classify approaches, despite not being competitive on MCAR, MAR or MNAR individually.
+The reason for this is that imputation accuracy of MVI techniques impacts data uncertainty, which then directly impacts overall model uncertainty (Gal et al., 2016).
 Deletion worsens stability compared to the model trained on clean data for all missingness types on diabetes, but, notably, only under MNAR on german.
@@
 4.4. Imputation Quality and Fairness
-In Figure 6, we report the imputation quality of different MVI techniques according to F1 score (for categorical columns), RMSE (for numerical columns), and KL divergence (for both numerical and categorical columns, computed over the columns with nulls only) and compare it with the F1 of the downstream model.
+In Figure 6, we report the imputation quality of 10 most accurate MVI techniques per category according to F1 score (for categorical columns), RMSE (for numerical columns), and KL divergence (for both numerical and categorical columns, computed over the columns with nulls only) and compare it with the F1 of the downstream model.
+See Appendix C for an extended comparison of training times and accuracy across all MVI techniques.
 Imputation Quality.
 We find that MVI techniques with vastly different imputation quality can give rise to models with comparable F1 scores, indicating that imputation correctness is not sufficiently predictive of model correctness (Zhang and Long, 2021a).
-For example, in Figure 6(a) on the diabetes dataset, median-dummy has an F1 score (imputation) close to 0; clustering, datawig, median-mode and deletion have an F1 score (imputation) between 0.5-0.6, and auto-ml and miss-forest have an F1 score (imputation) close to 1, but the models trained with all of these techniques have an F1 close to 1.
+For example, in Figure 6(a) on the diabetes dataset, median-dummy has an F1 score (imputation) close to 0; auto-ml, miss-forest and nomi have an F1 score (imputation) close to 1; and all other imputers have an F1 score (imputation) between 0.5-0.6, but the models trained with all of these techniques have an F1 close to 1.
 This trend is consistent across all datasets and missingness types, and for numerical columns as well (see Figure 6(b)).
@@
 (c) KL divergence difference (imputation) vs TPRD (model)
-Figure 7. Imputation fairness vs model fairness: imputation fairness (F1 difference, RMSE difference and KL divergence difference) may not be indicative of model fairness (TPRD).
+Figure 7. Imputation fairness vs. model fairness: imputation fairness (F1 difference, RMSE difference and KL divergence difference) may not be indicative of model fairness (TPRD).
 Imputation Fairness.
@@
 5. Missingness Shift
-Next, we evaluate the correctness, fairness, and stability of MVI techniques under missingness shift.
+Next, we evaluate the correctness, fairness, and stability of 10 most effective MVI techniques from various categories under missingness shift.
 We simulate missingness shift in two ways: (i) by varying the missingness mechanism between training and test (S4-9 in Table 1); and (ii) by varying the missingness rates between training and test.
@@
 All other MVI techniques, including boostclean, are generally robust to higher training missingness rates, and only show a 5% decrease in F1 (compared to 10% decrease with deletion), even at rates as high as 50%.
+This is because even imperfect imputation provides valuable insights for the model, making deletion a less favorable choice.
 A notable exception is clustering, which, surprisingly and somewhat counter-intuitively, has higher F1 at higher training missingness rates, and is actually better under MNAR than under MAR and MCAR, for all missingness rates and scenarios.
@@
 We find that F1 generally decreases with an increase in test missingness, corroborating the findings of  Shadbahr et al.
+(2023) and Miao et al.
 (2022).
-A notable exception is miss-forest, which is robust to both forms of missingness shift (i.e., changing missingness mechanisms and missingness rates).
-As for training set missingness, F1 decreases with an increase in test missingness most steeply under MNAR (both during training and test).
+A notable exception is miss-forest, which is robust to both forms of missingness shift such as changing missingness rates (as observed in (Miao et al., 2022)) and missingness mechanisms.
+As for training set missingness, F1 decreases with an increase in test missingness most steeply under MNAR (both during training and test), further supporting the findings of Miao et al.
+(2022).
 And, once again, clustering is an exception to this trend, instead showing invariance to test missingness rates under MNAR train (irrespective of test missingness) but not under MCAR and MAR train.
@@
 As before, fairness is highly sensitive to shifts in test missingness.
-For boostclean, datawig and auto-ml, TPRD generally increases (fairness worsens) with an increase in test error rate, although this increase is not monotonic.
-miss-forest is particularly robust, and although it only matches the fairness of the model trained on clean data under MCAR training, it is agnostic to increases in test missingness rate, under all missingness scenarios.
+For boostclean, datawig, and mnar-pvae, TPRD generally increases (fairness worsens) with an increase in test error rate, although this increase is not monotonic.
+miss-forest, auto-ml, and nomi are particularly robust to increases in test missingness rate under all missingness scenarios.
 For simpler methods such as deletion, clustering, median-mode and median-dummy, there is no consistent trend, even when missingness types remain the same and only missingness rates change between train and test.
 For example with deletion and clustering, TPRD increases with increase in test missingness rate in scenario S1 (MCAR train, MCAR test) and S3 (MNAR train, MNAR test), but monotonically decreases with test missingness rate in scenario S2 (MAR train, MAR test).
-In summary, our results corroborate the findings of Guha et al.
+In summary,
+our results corroborate the findings of Guha et al.
 (2024), and are a cause for concern as they indicate that the MVI techniques that perform best during development may not preserve fairness post-deployment, where shifts are likely to occur.
@@
 deletion and clustering are the least stable methods and show a monotonic decrease in stability with increase in missingness rate, with strongest effect under MNAR.
-auto-ml and miss-forest are the most robust MVI techniques and preserve stability of the clean model under all missingness settings and error rates.
+miss-forest, auto-ml, and nomi are the most robust MVI techniques and preserve stability of the clean model under all missingness settings and error rates.
+This further highlights how imputation quality directly influences data uncertainty, ultimately impacting the overall uncertainty of the final model, as discussed in Section 4.3.
 Test set missingness.
 Figure 51 in Appendix 50 shows the Label Stability of the Random Forest model on diabetes as a function of test missingness rate.
-We find that missingness rate has little effect on model stability under all missingness settings, with the exception of clustering, which shows lower label stability at higher missingness rates, most pronounced in scenario S7 (MAR train, MNAR test).
+We find that test missingness rate has little effect on model stability under all missingness settings, with the exception of clustering, which shows lower label stability at higher missingness rates, most pronounced in scenario S7 (MAR train, MNAR test).
+6. Running Time
+Table 4.
+Table 4: Training time (in seconds) of MVI techniques averaged across single- and multi-mechanism scenarios (S1-3, S10).
+Imputers are sorted by running time on the folk_emp dataset, and datasets are ordered by the number of rows.
+Dataset shapes reflect training sets with 30% rows with nulls.
+Values represent mean running times across seeds, with standard deviations.
+Imputer diabetes (633, 17) german (700, 21) folk__\__inc (12000, 10) law__\__school (16638, 11) bank (32003, 13) heart (56000, 11) folk__\__emp (242112, 16)
+median-dummy $0.013\pm 0.000$ $0.014\pm 0.001$ $0.021\pm 0.001$ $0.024\pm 0.001$ $0.027\pm 0.000$ $0.053\pm 0.001$ $0.773\pm 0.024$
+median-mode $0.012\pm 0.000$ $0.014\pm 0.001$ $0.023\pm 0.001$ $0.025\pm 0.001$ $0.031\pm 0.001$ $0.067\pm 0.003$ $0.933\pm 0.019$
+deletion $0.013\pm 0.000$ $0.013\pm 0.001$ $0.024\pm 0.001$ $0.025\pm 0.000$ $0.046\pm 0.001$ $0.087\pm 0.004$ $1\pm 0.049$
+mnar__\__pvae $8\pm 0.705$ $14\pm 11$ $14\pm 1$ $22\pm 11$ $55\pm 30$ $42\pm 6$ $206\pm 7$
+edit__\__gain $2\pm 0.119$ $2\pm 0.141$ $13\pm 0.221$ $21\pm 2$ $30\pm 1$ $62\pm 7$ $215\pm 4$
+nomi $11\pm 3$ $14\pm 7$ $22\pm 2$ $22\pm 2$ $29\pm 1$ $38\pm 2$ $356\pm 20$
+tdm $932\pm 74$ $1023\pm 11$ $1297\pm 22$ $1172\pm 92$ $1412\pm 34$ $1310\pm 42$ $1449\pm 31$
+notmiwae $161\pm 101$ $217\pm 82$ $555\pm 2$ $804\pm 8$ $665\pm 284$ $1393\pm 535$ $2944\pm 725$
+gain $261\pm 12$ $298\pm 5$ $1115\pm 38$ $1484\pm 30$ $1964\pm 36$ $2892\pm 48$ $6148\pm 178$
+hivae $68\pm 0.784$ $95\pm 1$ $745\pm 15$ $1073\pm 11$ $2450\pm 130$ $3794\pm 129$ $7163\pm 317$
+k__\__means__\__clustering $10\pm 0.402$ $13\pm 0.442$ $27\pm 0.640$ $323\pm 7$ $998\pm 49$ $1037\pm 64$ $7427\pm 778$
+miss__\__forest $111\pm 17$ $244\pm 86$ $1758\pm 526$ $2530\pm 851$ $4307\pm 1310$ $6337\pm 1601$ $20358\pm 4934$
+datawig $596\pm 185$ $277\pm 59$ $604\pm 46$ $2361\pm 492$ $5089\pm 651$ $7592\pm 854$ $31060\pm 3743$
+automl $1953\pm 195$ $1805\pm 212$ $5559\pm 581$ $6803\pm 565$ $13893\pm 1687$ $19055\pm 2710$ $104476\pm 14743$
+Table 4 presents the training time of MVI techniques for each dataset, averaged across all unique training scenarios (single- and multi-mechanism S1-S3, S10).
+Our time efficiency analysis approach aligns with prior work (Wang et al., 2024b; Miao et al., 2021), who also focused on training time as inference times are comparably fast across all techniques.
+Our results reveal that statistical imputers deletion, median-mode, and median-dummy are the fastest, while still delivering competitive accuracy for larger datasets like heart and folk_emp, as shown in Figures 18 and 19.
+In contrast, miss-forest, datawig, and auto-ml exhibit the longest training times, with at least one of these methods achieving the highest imputation accuracy in most cases.
+Interestingly, auto-ml requires three times more training time than datawig, the second most computationally intensive technique.
+This difference is due to the auto-ML nature of auto-ml, which involves extensive hyperparameter and network architecture tuning.
+Among non-statistical techniques, mnar-pvae, edit-gain, and nomi are the most efficient in terms of runtime.
+Notably, nomi provides competitive accuracy comparable to miss-forest, datawig, and auto-ml, making it the most optimal technique in terms of the trade-off between imputation accuracy and training time.
+A particularly noteworthy comparison is between gain and edit-gain.
+As explained in Appendix A.1, EDIT accelerates the training of parametric imputation models.
+Applying EDIT to gain results in a 28x speedup on the folk_emp dataset, with even greater improvements for smaller datasets, as shown in Table 4.
+Importantly, with this running time acceleration, edit-gain achieves accuracy comparable to gain on most datasets, as shown in Figures 18 and 19.
 7. Summary of Experimental Findings
 Do not drop your nulls!
-Building on evidence in the scientific community (Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), our findings indicate that, in addition to hurting predictive performance, deletion also worsens stability and fairness, arguing against the use of this method in practice.
+Building on evidence in the scientific community (Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), our experiments show that deletion is the least effective strategy for model accuracy, fairness, and stability, especially in setups where each row contributes valuable information.
+Logically, deletion results in data loss, but its appropriateness depends on data quality rather than quantity.
+If rows are complete duplicates or contain erroneous values, deletion may be justified.
 Multiple imputation shows mixed results.
-There is conflicting evidence in the literature on the performance of multiple imputation (MI) (Joost R.
-van Ginkel and van der Voort, 2020), and our empirical findings are also mixed and somewhat unexpected.
+There is conflicting evidence on the performance of multiple imputation (MI) (Joost R.
+van Ginkel and van der Voort, 2020), and our empirical findings are similarly mixed and somewhat unexpected.
 Feng (2023) and Le Morvan et al.
-(2021) argue that MI has superior predictive performance compared to impute-then-classify approaches.
-Graham (2009) and McNeish (2017) find that MI performs well even with less data and small error rates.
-In contrast, we find that MI (specifically, boostclean) is only competitive on small datasets, and is unstable compared to simpler MVI techniques.
-This can likely be explained by the complexity-stability trade-offs (Domingos, 2000): MI uses a more complex model class and is able to minimize the empirical loss, but due to this very complexity it shows a large variation in predictions under small train set perturbations.
+(2021) argue that MI outperforms impute-then-classify approaches in predictive performance, while Graham (2009) and McNeish (2017) find MI effective even with limited data and small error rates.
+In contrast, we find that MI (specifically, boostclean) is only competitive on small datasets and is less stable than simpler MVI techniques.
+This likely stems from the complexity-stability trade-off (Domingos, 2000; Breiman, 2001): MI employs a more complex model class that minimizes empirical loss but exhibits greater prediction variance under small training set perturbations.
 Fairness is highly missingness-specific.
@@
 Shadbahr et al.
-(2023) posit that distributional imputation quality metrics are a more reliable indicator of a model’s predictive performance than discrepancy metrics.
-We find that neither discrepancy nor distributional metrics are sufficiently indicative of downstream model performance (according to F1), indicating that poor imputation quality can be overcome by a strong learner.
-Further, we find that imputation fairness metrics are not predictive of model fairness: perfect imputation fairness can result in downstream model unfairness.
-Conversely, models trained using MVI techniques with imputation unfairness can still have good downstream fairness.
+(2023) argue that distributional imputation quality metrics better predict model performance than discrepancy metrics.
+However, we find that neither reliably indicates downstream performance (F1), as strong learners can compensate for poor imputation quality.
+Moreover, imputation fairness metrics do not predict model fairness: perfect imputation fairness can still lead to model unfairness, while models trained with imputation-unfair MVI techniques can achieve good downstream fairness.
 (a) MCAR missingness
@@
 TPRD and TNRD values close to 0 are ideal (fair), so we compute correlations using $TPRD\_Reversed=1-|TPRD|$ and $TNRD\_Reversed=1-|TNRD|$.
-Supplementary plots for each combination of training and test missingness are in Appendix E.
+Supplementary plots are in Appendix E.
 Model stability depends more on the dataset size and MVI technique than on the missingness scenario.
@@
 Sensitivity to train and test missingness rates.
-The predictive performance of models (according to F1) is more sensitive to missingness rates during testing than during training.
-Fairness, on the other hand, is highly sensitive to both training and test missingness rates.
-Lastly, there is little effect of test missingness rates on model stability, but most MVI techniques are more unstable at higher training missingness rates.
-No good methods for MNAR missingness.
+Model performance (F1) is more affected by test missingness than training missingness.
+Fairness, however, is highly sensitive to both.
+Model stability is largely unaffected by test missingness, but most MVI techniques become more unstable with higher training missingness.
+Existing MNAR-specific methods are insufficient.
 MNAR is theoretically the hardest setting to model.
-Empirically, we find that MVI techniques perform poorly with MNAR during testing.
-Conversely, even methods that are trained on MNAR but evaluated on “simpler” test missingness (MCAR and MAR) perform poorly.
-Further, we find that model F1 and stability are more sensitive to missingness rates under MNAR than under MCAR and MAR.
-Designing MVI techniques that are effective under MNAR is an open research direction.
+Empirically, we find that MVI techniques perform poorly with MNAR during testing, even the two existing MNAR-specific methods (not-miwae and mnar-pvae).
+For instance, experimental results show that model F1 and stability are more sensitive to missingness rates under MNAR than under MCAR and MAR.
+This performance gap arises because these methods rely on assumptions that often fail in real-world settings and are evaluated on limited, mostly large datasets with a single scenario, lacking the diversity of our benchmark.
+Developing improved MNAR-specific MVI methods and establishing more comprehensive benchmarks for their evaluation are critical directions for future research.
 Deep learning outperforms tree-based models on large datasets.
-There has been an explosion of interest in deep learning for tabular data (Hwang and Song, 2023; Gorishniy et al., 2021; Borisov et al., 2022; Ye et al., 2024; Fayaz et al., 2022).
-Recall that in Figures 3, 4, and 5 we reported results only for the best-performing models (in terms of F1), to streamline exposition.
-We find that tree-based methods like random forest and gradient boosted trees still outperform new deep table-learning models like GANDALF (Joseph and Raj, 2022) on small datasets (such as diabetes and german), corroborating the findings of Grinsztajn et al.
-(2022) and Shwartz-Ziv and Armon (2022), on tabular datasets with (simulated) missingness.
-GANDALF, however, is the best-performing model on larger datasets (heart and folk-employment in our study), underscoring the growing effectiveness of deep learning on a variety of data modalities.
+Interest in deep learning for tabular data has surged (Hwang and Song, 2023; Gorishniy et al., 2021; Borisov et al., 2022; Ye et al., 2024; Fayaz et al., 2022).
+Recall that Figures 3, 4, and 5 show only the best-performing models (by F1) for clarity.
+Tree-based methods like random forests and gradient-boosted trees outperform deep table-learning models like GANDALF (Joseph and Raj, 2022) on small datasets (diabetes, german), aligning with prior findings (Grinsztajn et al., 2022; Shwartz-Ziv and Armon, 2022).
+However, GANDALF excels on larger datasets (heart, folk-employment), highlighting the increasing effectiveness of deep learning across diverse data modalities.
 Best performing approaches under single- and multi-mechanism missingness can be different.
-A key novelty in our study is the evaluation under mixed or multi-mechanism missingness— which is a more complex but realistic missingness pattern, likely to be encountered in practice.
-We find that trends are different under single- and multi-mechanism missingness.
-Figure 11 shows that evaluation metrics and model types such as gandalf__\__clf and lgbm__\__clf are uncorrelated under single-mechanism missingness, but correlated under mixed missingness, for example for F1 ($\rho$=-0.24 for gandalf__\__clf and $\rho$=-0.37 for lgbm__\__clf) and fairness according to TPRD ($\rho$=0.55 for gandalf__\__clf and $\rho$=-0.4 for lgbm__\__clf).
+A key novelty of our study is the evaluation under mixed or multi-mechanism missingness—a more complex but realistic pattern likely encountered in practice.
+We find that performance trends vary between single- and multi-mechanism missingness.
+Figure 11 shows that model types like gandalf__\__clf and lgbm__\__clf are uncorrelated with performance metrics under single-mechanism missingness but become correlated under mixed missingness: F1 correlations are $\rho$=-0.24 for gandalf__\__clf and $\rho$=-0.37 for lgbm__\__clf, while fairness (TPRD) correlations are $\rho$=0.55 for gandalf__\__clf and $\rho$=-0.4 for lgbm__\__clf.
 It’s complicated!
-Prior experimental work (Hasan et al., 2021; Lin and Tsai, 2020; Joel and Doorsamy, 2022; Shadbahr et al., 2023; García Laencina et al., 2010) concluded that there is no definitive answer to the question of which missing value imputation technique is best according to predictive performance.
-Our findings support this and demonstrate trade-offs between F1, fairness and stability, based on the architecture of the predictive model.
-For example, from Figure 11, under multi-mechanism missingness, logistic regression (lr__\__clf) is positively correlated with F1 ($\rho$=0.23) and stability ($\rho$=0.38) but negatively correlated with fairness according to TPRD ($\rho$=-0.23).
-Random forest (rf__\__clf), on the other hand, is weakly positively correlated with fairness according to TPRD ($\rho$=0.16) and F1 ($\rho$=0.25), but negatively correlated with stability ($\rho$=-0.24).
-Lastly, deep table-learning (gandalf__\__clf) is strongly positively correlated with fairness according to TPRD ($\rho$=0.55), weakly correlated with stability ($\rho$=0.12), and negatively correlated with F1 ($\rho$=-0.24).
-This is not bad news, but rather an acknowledgement of the inherent complexity of the task of learning from incomplete data, and a call for more rigorous and holistic evaluation protocols— of the kind demonstrated in this work — to select the most suitable missing value imputation technique and model architecture for a given task.
+Prior studies (Hasan et al., 2021; Lin and Tsai, 2020; Joel and Doorsamy, 2022; Shadbahr et al., 2023; García Laencina et al., 2010) have found no universally best missing value imputation technique for predictive performance.
+Our findings reinforce this, revealing trade-offs between F1, fairness, and stability that depend on the predictive model’s architecture.
+For instance, under multi-mechanism missingness (Figure 11), logistic regression (lr__\__clf) correlates positively with F1 ($\rho$=0.23) and stability ($\rho$=0.38) but negatively with fairness (TPRD, $\rho$=-0.23).
+Random forest (rf__\__clf) shows weak positive correlations with fairness (TPRD, $\rho$=0.16) and F1 ($\rho$=0.25) but a negative correlation with stability ($\rho$=-0.24).
+Deep table-learning (gandalf__\__clf) is strongly correlated with fairness (TPRD, $\rho$=0.55), weakly with stability ($\rho$=0.12), and negatively with F1 ($\rho$=-0.24).
+This is not bad news, but rather an acknowledgement of the complexity of learning from incomplete data and the need for rigorous, holistic evaluation protocols, such as those demonstrated in this work, to select the most suitable imputation method and model architecture for a given task.
 8. Conclusions, Limitations and Future Work
 Conclusions.
-We presented Shades-of-Null, an evaluation suite for responsible missing value imputation.
-A key novelty of our work is in evaluating the predictive performance of models trained on imputed data alongside fairness and stability.
-Another novelty is our modeling of realistic missingness scenarios that go beyond Rubin’s classic settings of MCAR, MAR and MNAR to include multi-mechanism missingness and missingness shift.
-We showcased results of an extensive study, with 23,940 experimental pipelines, that uses our framework to evaluate a variety of MVI methods in realistic missingness scenarios.
-We synthesized interesting trends, and identified trade-offs between evaluation metrics.
+We introduced Shades-of-Null, an evaluation suite for responsible missing value imputation.
+A key contribution is evaluating fairness and stability alongside predictive performance.
+Additionally, we model realistic missingness scenarios beyond Rubin’s MCAR, MAR, and MNAR, incorporating multi-mechanism missingness and missingness shift.
+Through 29,736 experimental pipelines, we assessed various MVI methods under realistic missingness conditions, revealing key trends and trade-offs across evaluation metrics.
 Limitations.
@@
 Future work.
-It is interesting future work to create research datasets with naturally-occurring missingness shifts.
-Further, while predictive performance, fairness and stability can be seen as orthogonal dimensions of performance, we have uncovered trade-offs among them that arise due to missingness in the data, the choice of a missing value imputation method, and, in some cases, the architecture of the model trained on the data post-imputation.
-This, coupled with the finding that there is no universally best-performing imputation method, motivates the development of holistic metrics, evaluation procedures, and imputation techniques that are grounded in the lifecycle of data collection, processing and use.
-That is, much exciting work remains in responsible data engineering, of which missing value imputation is an important dimension, and to which our work makes a contribution.
+Creating research datasets with naturally occurring missingness shifts is a promising direction.
+While predictive performance, fairness, and stability are often treated as orthogonal, we uncover trade-offs driven by missingness, imputation choice, and, in some cases, model architecture.
+The absence of a universally best imputation method highlights the need for holistic metrics, evaluation procedures, and techniques grounded in the data lifecycle.
+Missing value imputation remains a critical aspect of responsible data engineering, to which our work makes a contribution.
 Appendix Overview
 This supplemental material provides significantly more detail regarding the results presented in this paper.
-The organization is as follows: We first provide additional details on the implementation of the evaluation suite, which include a broad description of the MVI techniques used in our study (Appendix A.1) and an explanation of the corrections and enhancements made to existing benchmarks and MVI techniques to ensure the development of our truly fair evaluation suite (Appendix A.2).
+The organization is as follows: We first provide additional details on the implementation of the evaluation suite, which include a broad description and configuration of the MVI techniques used in our study (Appendix A.1) and an explanation of the corrections and enhancements made to existing benchmarks and MVI techniques to ensure the development of our truly fair evaluation suite (Appendix A.2).
 Next, we extend the description of our experimental settings to facilitate reproducibility (Appendix B).
@@
 In Appendix C, we offer additional results to those discussed in Section 4.
-In particular, Appendix LABEL:apdx:single-mechanism-imp-quality presents the imputation quality of all MVI techniques across all datasets from the perspectives of accuracy, distributional alignment, and fairness, while Appendix C.1 showcases the fairness of predictive models measured by metrics beyond the TPRD mentioned in Section 4.
-Finally, Appendix D presents supplementary experimental findings on the missingness shift discussed in Section 5.
+In particular, Appendix C.1 presents the imputation quality of MVI techniques across all datasets from the perspectives of accuracy and training time, while Appendix C.2 showcases the fairness of predictive models measured by metrics beyond the TPRD mentioned in Section 4 and includes scatter plots for all imputers that extend scatter plots from Section 4.4.
+Then, Appendix D presents supplementary experimental findings on the missingness shift discussed in Section 5.
 Specifically, Appendix D.1 examines the accuracy, stability, and fairness dimensions of model performance under missingness shift with fixed train and test error rates, highlighting its effects on various model types and datasets.
 Furthermore, Appendix D.2 explores further experimental results on model performance under missingness shift with variable train and test error rates.
+Finally, Appendix E presents correlation plots for other missingness settings than in Section 7.
 Appendix A Additional details on the evaluation suite implementation
 A.1. Missing Value Imputation Techniques
-In our evaluation suite, we evaluate 8 MVI techniques, classified into 5 main categories based on the taxonomy presented in (Emmanuel et al., 2021).
-In each category, we selected one or two SOTA methods for our comparison.
+In our evaluation suite, we evaluate 15 MVI techniques, classified into 8 broad categories, as outlined in Section 3.2.
+These categories comprehensively cover the spectrum of MVI methods discussed in recent surveys (Emmanuel et al., 2021; Jäger et al., 2021; Miao et al., 2022; Abdelaal et al., 2023) and peer-reviewed papers (Wang et al., 2024b; Wu et al., 2023; Miao et al., 2021), which propose new MVI techniques and compare against them.
+In each category, we selected several SOTA methods for our comparison.
 Each technique is capable of imputing both categorical and numerical features.
-All configurations and hyperparameters for the MVI techniques described below are specified in the configuration module within the source code.
+For all MVI techniques, we adopt the default parameter settings from original papers or their source code.
   1. (1)
@@
   2. (2)
-    Statistical techniques:
-    1. (a)
-      median-mode: We impute missing values with the median of the complete case for numerical columns and mode (most frequently occurring value) for categorical columns.
-    2. (b)
-      median-dummy: We impute missing values with the median of the complete case for numerical columns and assign a (new) dummy category for nulls in categorical columns.
+    median-mode (statistical): We impute missing values with the median of the complete case for numerical columns and mode (most frequently occurring value) for categorical columns.
   3. (3)
-    Machine learning approaches:
-    1. (a)
-      miss-forest (supervised) (Stekhoven and Bühlmann, 2011): This approach iteratively trains a RF model on a set of clean samples (no missingness) and predicts the missing values.
-    2. (b)
-      clustering (unsupervised) (Gajawada and Toshniwal, 2012): We assign the missing value to a cluster, based on the distance from the cluster center, and then impute the missing value with the mean or mode of the data points in that cluster.
+    median-dummy (statistical): We impute missing values with the median of the complete case for numerical columns and assign a (new) dummy category for nulls in categorical columns.
+  4. (4)
+    miss-forest (Stekhoven and Bühlmann, 2011) (supervised ML): This approach iteratively trains an RF model on a set of clean samples (no missingness) and predicts the missing values.
+We tune RandomForestClassifier for categorical columns with nulls and RandomForestRegressor for numerical columns with nulls.
+The maximum number of iterations for miss-forest is set to 10.
+  5. (5)
+    clustering (Gajawada and Toshniwal, 2012) (unsupervised ML): We assign the missing value to a cluster, based on the distance from the cluster center, and then impute the missing value with the mean or mode of the data points in that cluster.
 [^2] We used the k-prototypes implementation from https://github.com/nicodv/kmodes We implemented this method ourselves, as no existing solution supports null imputation based on clustering for data with both numerical and categorical features.
   6. (6)
-    Deep learning approaches:
-    1. (a)
-      datawig (Biessmann et al., 2019): This method implements deep learning modules combined with neural architecture search and end-to-end optimization of the imputation pipeline.
-    2. (b)
-      auto-ml (Jäger et al., 2021): This approach uses the AutoML library autokeras (Jin et al., 2019) to implement the discriminative deep learning imputation method.
+    datawig (Biessmann et al., 2019) (discriminative DL): This method implements deep learning modules combined with neural architecture search and end-to-end optimization of the imputation pipeline.
+We train datawig for 100 epochs and a single iteration using a batch size of 64.
+The final_fc_hidden_units parameter is tuned for each dataset by selecting the optimal value from the set {1, 10, 50, 100}.
+  7. (7)
+    auto-ml (Jäger et al., 2021) (discriminative DL): This approach uses the AutoML library autokeras (Jin et al., 2019) to implement the discriminative deep learning imputation method.
 For categorical columns, autokeras’ StructuredDataClassifier is employed, while StructuredDataRegressor is utilized for numerical columns.
 These classes manage data encoding and optimize model architecture and hyperparameters autonomously.
+We train auto-ml for 100 epochs with a validation split of 0.2, using up to 50 trials to tune hyperparameters.
+  8. (8)
+    gain (Yoon et al., 2018) (generative DL): This method adapts the Generative Adversarial Networks (GAN) framework to impute missing data.
+The generator observes partially observed data and imputes the missing components conditioned on the observed data, outputting a completed data vector.
+The discriminator distinguishes between observed and imputed components, using a hint vector to focus on specific imputation tasks.
+The adversarial training ensures that the generator learns to produce imputations that closely match the true data distribution.
+We train gain by tuning its hyperparameters based on the following grid: $\alpha\in\{1,10\}$, hint rate $\in\{0.7,0.9\}$, batch size $\in\{64,128\}$, generator learning rate $\in\{1\times 10^{-5},1\times 10^{-4},0.0005\}$, and discriminator learning rate $\in\{1\times 10^{-6},1\times 10^{-5},0.00005\}$.
+  9. (9)
+    hi-vae (Nazabal et al., 2020) (generative DL): This approach extends the Variational Autoencoder (VAE) framework to handle incomplete and heterogeneous datasets.
+HI-VAE models mixed numerical (e.g., real-valued, positive real-valued, and count) and categorical (e.g., ordinal and nominal) data types.
+It employs a hierarchical architecture with a shared latent space to capture correlations among attributes and a likelihood model tailored to each data type.
+By training on observed data and utilizing an Evidence Lower Bound (ELBO) computed only for observed entries, HI-VAE achieves robust imputation without overfitting.
+We train hi-vae for 2000 epochs using a batch size of 128, with latent dimensions set as $z=10$, $y=5$, and $s=10$, and a learning rate of $1\times 10^{-3}$.
+  10. (10)
+    not-miwae (Ipsen et al., 2020) (MNAR-specific): This method extends deep generative modeling to handle Missing Not At Random (MNAR) data by explicitly modeling the missing data mechanism alongside the data distribution.
+It leverages importance-weighted variational inference to jointly optimize the parameters of the data and missing mechanisms.
+The method uses stochastic gradients derived via reparameterization in both latent and data spaces, allowing efficient training.
+By incorporating prior knowledge about the missingness process into a flexible missing model, not-miwae achieves robust imputation under MNAR scenarios.
+We train not-miwae for a maximum of 100,000 iterations using a batch size of 16, with 128 hidden units, and a latent dimensionality of $L=10,000$, employing the self-masking process selfmasking_known.
+  11. (11)
+    mnar-pvae (Ma and Zhang, 2021) (MNAR-specific): This approach introduces an identifiable deep generative model for handling Missing Not At Random (MNAR) data.
+By leveraging identifiability principles and extending variational autoencoders (VAEs), GINA ensures that the underlying data-generating process can be uniquely recovered.
+It jointly models the data distribution and the missingness mechanism, integrating auxiliary variables to achieve identifiability under mild assumptions.
+Through a combination of importance-weighted variational inference and a flexible neural architecture, GINA enables robust and unbiased imputation of MNAR data.
+We train mnar-pvae (GINA) for 400 epochs and a single iteration using a batch size of 100 and a learning rate of $1\times 10^{-3}$.
+The network is configured with an embedding dimension of 20, a latent dimension of 20, encoder and decoder layers each containing 10 units, and uses Tanh as the non-linearity.
   12. (12)
@@
 To generate the weak learners, BoostClean iteratively selects a single imputation technique, applies it to a training set (with missingness), and fits a new model on the newly imputed training set.
+We train boostclean for 5 iterations and tune its underlying prediction model for each iteration.
+  13. (13)
+    nomi (Wang et al., 2024b) (recent): This method introduces an uncertainty-driven network for missing data imputation.
+NOMI integrates three key components: a retrieval module, a Neural Network Gaussian Process Imputator (NNGPI), and an uncertainty-based calibration module.
+The retrieval module identifies local neighbors of incomplete samples, while the NNGPI combines the probabilistic modeling of Gaussian Processes with the representation power of neural networks to impute missing values and quantify uncertainty.
+The uncertainty-based calibration module dynamically refines the imputations by balancing predictions across iterations, leveraging uncertainty to improve reliability.
+We train nomi with a maximum of 3 iterations, using 10 neighbors for imputation, the $l_{2}$ similarity metric, a temperature parameter $\tau=1.0$, and a weighting coefficient $\beta=0.8$.
+  14. (14)
+    tdm (Zhao et al., 2023) (recent): This approach proposes Transformed Distribution Matching (TDM), a novel approach to missing data imputation that leverages optimal transport in a transformed latent space.
+TDM employs deep invertible neural networks to map data into a latent space where the geometry better reflects the underlying data structure.
+It minimizes the Wasserstein distance between the transformed distributions of two batches of data to achieve imputation, while avoiding overfitting through mutual information constraints.
+We train tdm for 10,000 iterations with a batch size of 512 and a learning rate of $1\times 10^{-2}$.
+  15. (15)
+    edit-gain (Miao et al., 2021) (recent): This technique introduces an efficient and effective data imputation framework that leverages influence functions to accelerate the training of parametric imputation models.
+EDIT consists of two key modules: the Imputation Influence Evaluation (IIE) module, which estimates the influence power of samples on the imputation model’s predictions, and the Representative Sample Selection (RSS) module, which constructs a minimal representative sample set to satisfy user-specified accuracy guarantees.
+Additionally, EDIT employs a weighted loss function that emphasizes high-influence samples, boosting imputation accuracy while reducing training cost.
+We employ gain (Yoon et al., 2018) as the underlying imputer for EDIT, as it is the only method shared with us directly by the authors of the original paper.
+We train edit-gain using a batch size of 128 for 30 epochs, with an $\alpha$ parameter value of 1.
+The initial sample size is set to 6000 for all our datasets, matching the initial sample size of the smallest dataset used in the original paper, since all our datasets are of the same size or smaller.
 A.2. Corrections and Enhancements in Benchmarks and MVI Techniques
@@
 Presumably, the authors chose to use these imputers without modification.
-In order to ensure a fair comparison of MVI techniques as they would be used in production systems, in our study, we standardized all 8 MVI techniques to meet the following requirements: 1) fitting a null imputers only on a training set and then applying it to both training and test sets; 2) enabling hyper-parameter tuning of a null imputer (if applicable); 3) controlling randomization of a null imputer using seeds (if applicable).
+In order to ensure a fair comparison of MVI techniques as they would be used in production systems, in our study, we standardized our MVI techniques to meet the following requirements: 1) fitting a null imputers only on a training set and then applying it to both training and test sets; 2) enabling hyper-parameter tuning of a null imputer (if applicable); 3) controlling randomization of a null imputer using seeds (if applicable).
 Below is a summary of the enhancements we implemented for MVI techniques used in our evaluation suite:
@@
 In summary, our evaluation suite offers users not only a robust evaluation framework but also incorporates fixes to the MVI implementations previously available.
+A.3. Datasets and Tasks
+diabetes [^4] https://www.kaggle.com/datasets/tigganeha4/diabetes-dataset-2019 (Tigga and Garg, 2020) was collected in India through a questionnaire including 18 questions related to health, lifestyle, and family background.
+A total of 952 participants are characterized by 17 attributes (13 categorical, 4 numerical) and a binary target variable that represents whether a person is diabetic.
+Here, sex is the sensitive attribute, with “female” as the disadvantaged group.
+german [^5] https://archive.ics.uci.edu/dataset/144/statlog+german+credit+data (Hofmann, 1994) is a popular fairness dataset that contains records of creditworthiness assessments, classifying individuals as high or low credit risks.
+It contains information on 1,000 individuals characterized by 21 attributes (14 categorical, 7 numerical), including credit history, occupation and housing information.
+Here, sex and age are the sensitive attributes, with “female” and “age25absent25$\leq$ 25≤ 25” as the disadvantaged groups.
+Folktables (Ding et al., 2021) is another popular fairness dataset derived from US Census data from all 50 states between 2014-2018.
+The dataset has several associated tasks, of which we selected two: (i) ACSIncome (folk-income) is a binary classification task to predict whether an individual’s annual income is above \$50,000, from 10 features (8 categorical, 2 numerical) including educational attainment, work hours per week, marital status, and occupation.
+We use data from Georgia from 2018, subsampled to 15k rows.
+(ii) ACSEmployment (folk-employment) is a binary classification task to predict whether an individual is employed, from 16 features (15 categorical, 1 numerical) including educational attainment, employment status of parent, military status, and nativity.
+We use data from California from 2018, with 302,640 rows (we do not subsample).
+In both tasks, sex and race are the sensitive attributes, with “female” and “non-White” as the disadvantaged groups.
+law-school [^6] https://www.kaggle.com/datasets/danofer/law-school-admissions-bar-passage  (Wightman, 1998) was gathered through a survey conducted by the Law School Admission Council (LSAC) across 163 law schools in the US in 1991, and contains admissions records of 20,798 applicants, characterized by 11 attributes (5 categorical, 6 numerical), including LSAT scores and college GPAs.
+The task is to predict whether a candidate would pass the bar exam.
+Here, sex and race are the sensitive attributes, with “female” and “non-White” as the disadvantaged groups.
+bank [^7] https://archive.ics.uci.edu/dataset/222/bank+marketing  (Moro et al., 2014) contains data from direct marketing campaigns by a Portuguese bank, between 2008 and 2013.
+It contains information on 40,004 potential customers, with 13 attributes (7 categorical, 6 numerical), including occupation, marital status, education, and a binary target that indicates whether the individual subscribed for a term deposit.
+Here, age is the sensitive attribute, with $<25$ and $>60$ as the disadvantaged group.
+heart [^8] https://www.kaggle.com/datasets/sulianova/cardiovascular-disease-dataset consists of patient measurements with respect to cardiovascular diseases, with information of 70,000 individuals characterized by 11 attributes (6 categorical, 5 numerical), including age, height, weight, blood pressure, and a binary target indicating whether the patient has heart disease.
+Here, sex is the sensitive attribute, with “female” as the disadvantaged group.
 Appendix B Additional Experimental Details for Reproducibility
 B.1. Computing Infrastructure
-Our large-scale experimental study incorporated 8 null imputation techniques, 6 model types, 7 datasets, 10 evaluation scenarios, and 6 seeds for controlling randomization across the entire pipeline.
-This configuration led to an overwhelming total of 23,940 experimental pipelines.
+Our large-scale experimental study incorporated 15 null imputation techniques, 6 model types, 7 datasets, 10 evaluation scenarios, and 6 seeds for controlling randomization across the entire pipeline.
+This configuration led to an overwhelming total of 29,736 experimental pipelines.
 The complexity was further compounded by the necessity to tune both null imputers and ML models within each pipeline, employing a bootstrap of 50 estimators to assess model stability.
@@
 For null imputation using datawig, we employed an RTX8000 NVIDIA GPU card with 48 GB of GPU memory, while all other tasks were handled by a 2x Intel Xeon Platinum 8268 24C 205W 2.9GHz processor.
+All evaluated imputation algorithms were implemented using Python 3.9, utilizing their original source code.
+Some of this code was generously shared directly by the respective authors, for which we express our sincere gratitude.
+All dependencies are specified in our repository, along with detailed instructions in the README for installing and running the evaluation suite.
 To effectively store the results of the experiments in the MongoDB cluster, we designed a data model with unique GUIDs for each experimental pipeline.
@@
 Similarly to Section 3.1, this subsection outlines realistic missingness scenarios based on intrinsic dataset properties.
-Tables 4-10 report the demographic composition of all datasets, specifically, the proportions and base rates of each protected group.
+Tables 5-11 report the demographic composition of all datasets, specifically, the proportions and base rates of each protected group.
 Figures 12-17 display the Spearman correlation coefficients between covariates and the target variable, as well as feature importance derived using scikit-learn’s built-in functionality for datasets not covered in Section 3.1.
@@
 These proportions, base rates, correlation coefficients, and feature importance guided the conditions of our missingness scenarios, aiming to reflect realistic instances of non-response caused by disparate access, distrust, or procedural injustice, where disadvantaged groups exhibit more missing values than privileged groups.
-Tables 11-16 detail the missingness scenarios for the folk-employment, folk-income, bank, heart, law-school, and german datasets, respectively.
+Tables 12-17 detail the missingness scenarios for the folk-employment, folk-income, bank, heart, law-school, and german datasets, respectively.
 Table 5. Table 5: Proportions and Base Rates for diabetes.
@@
 Table 12. Table 12: Missingness scenarios for an error rate of 30% for folk-employment. AGEP is a numerical column; DIS, MIL, SCHL are categorical columns.
-Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
+[…]
 MCAR DIS, MIL, AGEP, SCHL N/A 0.3 0.3
@@
 Table 13. Table 13: Missingness scenarios for an error rate of 30% for folk-income. WKHP, AGEP are numerical columns; SCHL, MAR are categorical columns.
-Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
+[…]
 MCAR WKHP, AGEP, SCHL, MAR N/A 0.3 0.3
@@
 Table 14. Table 14: Missingness scenarios for an error rate of 30% for bank. balance, campaign are numerical columns; education, job are categorical columns.
-Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
+[…]
 MCAR balance, campaign, education, job N/A 0.3 0.3
@@
 Table 15. Table 15: Missingness scenarios for an error rate of 30% for heart. weight, height are numerical columns; cholesterol, gluc are categorical columns.
-Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
+[…]
 MCAR weight, height, cholesterol, gluc N/A 0.3 0.3
@@
 Table 16. Table 16: Missingness scenarios for an error rate of 30% for law-school. zfygpa, ugpa are numerical columns; fam_inc, tier are categorical columns.
-Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
+[…]
 MCAR zfygpa, ugpa, fam_inc, tier N/A 0.3 0.3
@@
 duration, credit-amount are numerical columns; checking-account, savings-account, employment-since are categorical columns.
-Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
+[…]
 MCAR duration, credit-amount, checking-account, savings-account, employment-since N/A 0.3 0.3
@@
 Figure 17. EDA for designing missingness scenarios in german.
-Appendix C Additional Experimental Results for Single-mechanism Missingness
-C.1. Additional Accuracy and Fairness Metrics of Predictive Model
+Appendix C Additional Experimental Results for Single- and Multi-mechanism Missingness
+C.1. Accuracy and Efficiency of MVI Techniques
+In this section, we extend Section 4.4 by presenting additional results on the accuracy and efficiency of MVI techniques across single- and multi-mechanism scenarios (S1-S3, S10).
+Figures 18 and 19 illustrate imputation accuracy, measured by F1 scores for categorical columns and RMSE for numerical columns, respectively.
+Tables 18–22 summarize the training time of various MVI techniques (in seconds), averaged across these scenarios.
+In the tables, imputers are sorted by runtime on the folk_emp dataset, while datasets are ordered by their number of rows.
+Dataset shapes indicate training sets consisting of 70% complete rows and 30% rows with missing values.
+In the tables, the reported values represent mean runtimes across different seeds, with standard deviations included.
+Higher F1 scores and lower RMSE and runtimes are desirable for better performance.
+Notably, boostclean is excluded as it is a joint cleaning-and-training technique that does not generate imputed rows.
+Table 18 (same as Table 4 from Section 6, duplicated for convenience) reveals that statistical imputers such as deletion,
+median-mode, and median-dummy are the fastest, while still delivering competitive accuracy for larger datasets like heart and folk_emp, as shown in Figures 18 and 19.
+In contrast, miss-forest, datawig, and auto-ml exhibit the longest training times, with at least one of these methods achieving the highest imputation accuracy in most cases.
+Interestingly, auto-ml requires three times more training time than datawig, the second most computationally intensive technique.
+This difference is due to the auto-ML nature of auto-ml, which involves extensive hyperparameter and network architecture tuning.
+Among non-statistical techniques, mnar-pvae, edit-gain, and nomi are the most efficient in terms of runtime.
+Notably, nomi provides competitive accuracy comparable to miss-forest, datawig, and auto-ml, making it the most optimal technique in terms of the trade-off between imputation accuracy and training time.
+A particularly noteworthy comparison is between gain and edit-gain.
+As explained in Appendix A.1, EDIT accelerates the training of parametric imputation models.
+Applying EDIT to gain results in a 28x speedup on the folk_emp dataset, with even greater improvements for smaller datasets, as shown in Table 18.
+Importantly, with this runtime acceleration edit-gain achieves accuracy comparable to gain on most datasets, as shown in Figures 18 and 19.
+Tables 19–22 provide training times for individual scenarios (S1, S2, S3, and S10).
+These values are generally consistent with those in Table 18, with minor exceptions: nomi outperforms mnar-pvae and edit-gain in terms of speed for MCAR and MAR scenarios, and clustering is faster than hi-vae in the MAR scenario.
+It is worth noting that the training times in Tables 18–22 exhibit substantial standard deviations for many MVI techniques.
+These deviations arise because the statistics are computed across pipelines executed with different random seeds, resulting in varying training-test splits and model initializations.
+Consequently, standard deviations can be significant for all MVI techniques.
+Furthermore, mnar-pvae, not-miwae, miss-forest, datawig, and auto-ml show large standard deviations.
+For mnar-pvae and not-miwae, their MNAR-specific nature makes them highly sensitive to the data characteristics, leading to variability across splits.
+In the case of miss-forest, datawig, and auto-ml, the presence of parameters such as maximum iterations or trials introduces early stopping mechanisms that can be influenced by the data from different splits, further contributing to runtime variability.
+Table 18: Training time (in seconds) of MVI techniques averaged across single- and multi-mechanism scenarios (S1-3, S10).
+Imputers are sorted by runtime on the folk_emp dataset, and datasets are ordered by the number of rows.
+Dataset shapes reflect training sets with 30% rows with nulls.
+Values represent mean runtime across seeds, with standard deviations included.
+Imputer diabetes (633, 17) german (700, 21) folk__\__inc (12000, 10) law__\__school (16638, 11) bank (32003, 13) heart (56000, 11) folk__\__emp (242112, 16)
+median-dummy $0.013\pm 0.000$ $0.014\pm 0.001$ $0.021\pm 0.001$ $0.024\pm 0.001$ $0.027\pm 0.000$ $0.053\pm 0.001$ $0.773\pm 0.024$
+median-mode $0.012\pm 0.000$ $0.014\pm 0.001$ $0.023\pm 0.001$ $0.025\pm 0.001$ $0.031\pm 0.001$ $0.067\pm 0.003$ $0.933\pm 0.019$
+deletion $0.013\pm 0.000$ $0.013\pm 0.001$ $0.024\pm 0.001$ $0.025\pm 0.000$ $0.046\pm 0.001$ $0.087\pm 0.004$ $1\pm 0.049$
+mnar__\__pvae $8\pm 0.705$ $14\pm 11$ $14\pm 1$ $22\pm 11$ $55\pm 30$ $42\pm 6$ $206\pm 7$
+edit__\__gain $2\pm 0.119$ $2\pm 0.141$ $13\pm 0.221$ $21\pm 2$ $30\pm 1$ $62\pm 7$ $215\pm 4$
+nomi $11\pm 3$ $14\pm 7$ $22\pm 2$ $22\pm 2$ $29\pm 1$ $38\pm 2$ $356\pm 20$
+tdm $932\pm 74$ $1023\pm 11$ $1297\pm 22$ $1172\pm 92$ $1412\pm 34$ $1310\pm 42$ $1449\pm 31$
+notmiwae $161\pm 101$ $217\pm 82$ $555\pm 2$ $804\pm 8$ $665\pm 284$ $1393\pm 535$ $2944\pm 725$
+gain $261\pm 12$ $298\pm 5$ $1115\pm 38$ $1484\pm 30$ $1964\pm 36$ $2892\pm 48$ $6148\pm 178$
+hivae $68\pm 0.784$ $95\pm 1$ $745\pm 15$ $1073\pm 11$ $2450\pm 130$ $3794\pm 129$ $7163\pm 317$
+k__\__means__\__clustering $10\pm 0.402$ $13\pm 0.442$ $27\pm 0.640$ $323\pm 7$ $998\pm 49$ $1037\pm 64$ $7427\pm 778$
+miss__\__forest $111\pm 17$ $244\pm 86$ $1758\pm 526$ $2530\pm 851$ $4307\pm 1310$ $6337\pm 1601$ $20358\pm 4934$
+datawig $596\pm 185$ $277\pm 59$ $604\pm 46$ $2361\pm 492$ $5089\pm 651$ $7592\pm 854$ $31060\pm 3743$
+automl $1953\pm 195$ $1805\pm 212$ $5559\pm 581$ $6803\pm 565$ $13893\pm 1687$ $19055\pm 2710$ $104476\pm 14743$
+(a) Missing Completely At Random (MCAR)
+(b) Missing At Random (MAR)
+(c) Missing Not At Random (MNAR)
+(d) Mixed missingness (MCAR & MAR & MNAR)
+Figure 18.
+F1 of different imputation strategies (colors in the legend) averaged across categorical columns with nulls per dataset (x-axis) and missingness mechanisms (subplots).
+Datasets are ordered in increasing order by size.
+(a) Missing Completely At Random (MCAR)
+(b) Missing At Random (MAR)
+(c) Missing Not At Random (MNAR)
+(d) Mixed missingness (MCAR & MAR & MNAR)
+Figure 19.
+RMSE of different imputation strategies (colors in the legend) averaged across numerical columns with nulls per dataset (x-axis) and missingness mechanisms (subplots).
+Datasets are ordered in increasing order by size.
+Table 19: Training time (in seconds) of MVI techniques for the MCAR scenario (S1).
+Imputers are sorted by runtime on the folk_emp dataset, and datasets are ordered by the number of rows.
+Dataset shapes reflect training sets with 30% rows with nulls.
+Values represent mean runtime across seeds, with standard deviations included.
+Imputer diabetes (633, 17) german (700, 21) folk__\__inc (12000, 10) law__\__school (16638, 11) bank (32003, 13) heart (56000, 11) folk__\__emp (242112, 16)
+median-dummy $0.015\pm 0.001$ $0.016\pm 0.001$ $0.021\pm 0.000$ $0.025\pm 0.001$ $0.029\pm 0.000$ $0.057\pm 0.002$ $0.846\pm 0.042$
+median-mode $0.013\pm 0.001$ $0.016\pm 0.001$ $0.023\pm 0.000$ $0.029\pm 0.001$ $0.034\pm 0.000$ $0.072\pm 0.002$ $1\pm 0.037$
+deletion $0.014\pm 0.001$ $0.015\pm 0.000$ $0.027\pm 0.001$ $0.028\pm 0.001$ $0.049\pm 0.001$ $0.100\pm 0.006$ $1\pm 0.036$
+nomi $11\pm 4$ $16\pm 9$ $25\pm 2$ $23\pm 2$ $31\pm 1$ $36\pm 2$ $206\pm 5$
+edit__\__gain $2\pm 0.133$ $2\pm 0.155$ $14\pm 0.188$ $22\pm 2$ $31\pm 0.792$ $62\pm 7$ $216\pm 4$
+mnar__\__pvae $8\pm 1$ $10\pm 1$ $13\pm 1$ $44\pm 43$ $129\pm 121$ $43\pm 0.986$ $228\pm 6$
+tdm $897\pm 11$ $1034\pm 6$ $1289\pm 9$ $1095\pm 19$ $1720\pm 68$ $1195\pm 22$ $1448\pm 25$
+notmiwae $157\pm 84$ $249\pm 149$ $600\pm 8$ $751\pm 3$ $749\pm 349$ $1472\pm 637$ $2978\pm 578$
+gain $257\pm 26$ $261\pm 5$ $1112\pm 72$ $1421\pm 39$ $1902\pm 33$ $2748\pm 48$ $5668\pm 146$
+hivae $70\pm 1$ $90\pm 0.228$ $701\pm 8$ $1017\pm 6$ $2531\pm 320$ $3874\pm 86$ $7161\pm 415$
+k__\__means__\__clustering $0.921\pm 0.318$ $1\pm 0.326$ $1\pm 0.117$ $18\pm 1$ $26\pm 1$ $114\pm 24$ $7499\pm 720$
+miss__\__forest $94\pm 14$ $240\pm 92$ $1837\pm 577$ $2150\pm 989$ $4488\pm 1404$ $6269\pm 1610$ $15154\pm 5008$
+datawig $238\pm 69$ $279\pm 66$ $601\pm 70$ $2104\pm 336$ $4654\pm 755$ $7791\pm 921$ $33102\pm 1695$
+automl $2219\pm 367$ $1969\pm 131$ $5751\pm 604$ $6936\pm 252$ $14203\pm 1105$ $19315\pm 3505$ $104290\pm 19094$
+Table 20: Training time (in seconds) of MVI techniques for the MAR scenario (S2).
+Imputers are sorted by runtime on the folk_emp dataset, and datasets are ordered by the number of rows.
+Dataset shapes reflect training sets with 30% rows with nulls.
+Values represent mean runtime across seeds, with standard deviations included.
+Imputer diabetes (633, 17) german (700, 21) folk__\__inc (12000, 10) law__\__school (16638, 11) bank (32003, 13) heart (56000, 11) folk__\__emp (242112, 16)
+median-dummy $0.015\pm 0.001$ $0.016\pm 0.000$ $0.021\pm 0.000$ $0.025\pm 0.001$ $0.029\pm 0.000$ $0.057\pm 0.001$ $0.883\pm 0.044$
+median-mode $0.013\pm 0.001$ $0.016\pm 0.001$ $0.023\pm 0.000$ $0.027\pm 0.000$ $0.033\pm 0.000$ $0.070\pm 0.003$ $1\pm 0.018$
+deletion $0.014\pm 0.001$ $0.015\pm 0.001$ $0.026\pm 0.000$ $0.028\pm 0.000$ $0.048\pm 0.000$ $0.095\pm 0.004$ $1\pm 0.031$
+nomi $10\pm 3$ $14\pm 8$ $23\pm 1$ $24\pm 3$ $29\pm 1$ $38\pm 3$ $191\pm 3$
+edit__\__gain $2\pm 0.088$ $2\pm 0.120$ $13\pm 0.193$ $20\pm 2$ $30\pm 1$ $63\pm 7$ $209\pm 8$
+mnar__\__pvae $8\pm 0.480$ $10\pm 1$ $11\pm 0.646$ $16\pm 0.633$ $35\pm 1$ $41\pm 1$ $213\pm 10$
+tdm $907\pm 19$ $984\pm 11$ $1303\pm 8$ $1182\pm 43$ $1335\pm 12$ $1556\pm 57$ $1477\pm 8$
+notmiwae $235\pm 140$ $273\pm 0.540$ $586\pm 1$ $797\pm 10$ $803\pm 395$ $1498\pm 553$ $2894\pm 487$
+gain $299\pm 11$ $288\pm 4$ $1140\pm 25$ $1590\pm 34$ $1971\pm 34$ $2880\pm 44$ $6659\pm 171$
+k__\__means__\__clustering $0.886\pm 0.341$ $1\pm 0.319$ $1\pm 0.122$ $19\pm 2$ $26\pm 1$ $107\pm 22$ $7395\pm 715$
+hivae $71\pm 0.788$ $99\pm 1$ $734\pm 13$ $1134\pm 11$ $2337\pm 28$ $3585\pm 84$ $7557\pm 468$
+miss__\__forest $112\pm 14$ $251\pm 87$ $1874\pm 606$ $2976\pm 914$ $4493\pm 1218$ $7443\pm 2031$ $19323\pm 5650$
+datawig $271\pm 83$ $325\pm 85$ $642\pm 18$ $2398\pm 735$ $4858\pm 695$ $7820\pm 908$ $29815\pm 5383$
+automl $2172\pm 187$ $1822\pm 236$ $5495\pm 797$ $7036\pm 507$ $15469\pm 2313$ $19357\pm 3697$ $96490\pm 8667$
+Table 21: Training time (in seconds) of MVI techniques for the MNAR scenario (S3).
+Imputers are sorted by runtime on the folk_emp dataset, and datasets are ordered by the number of rows.
+Dataset shapes reflect training sets with 30% rows with nulls.
+Values represent mean runtime across seeds, with standard deviations included.
+Imputer diabetes (633, 17) german (700, 21) folk__\__inc (12000, 10) law__\__school (16638, 11) bank (32003, 13) heart (56000, 11) folk__\__emp (242112, 16)
+median-dummy $0.015\pm 0.001$ $0.015\pm 0.000$ $0.021\pm 0.001$ $0.025\pm 0.001$ $0.029\pm 0.001$ $0.057\pm 0.002$ $0.773\pm 0.005$
+median-mode $0.015\pm 0.000$ $0.016\pm 0.001$ $0.023\pm 0.000$ $0.027\pm 0.000$ $0.034\pm 0.001$ $0.067\pm 0.003$ $0.867\pm 0.006$
+deletion $0.014\pm 0.000$ $0.013\pm 0.001$ $0.025\pm 0.000$ $0.027\pm 0.000$ $0.046\pm 0.000$ $0.086\pm 0.004$ $1\pm 0.118$
+mnar__\__pvae $8\pm 0.560$ $26\pm 44$ $14\pm 0.825$ $16\pm 0.755$ $34\pm 0.941$ $51\pm 23$ $215\pm 8$
+edit__\__gain $2\pm 0.129$ $2\pm 0.144$ $13\pm 0.250$ $21\pm 2$ $30\pm 0.386$ $62\pm 7$ $219\pm 2$
+nomi $12\pm 5$ $14\pm 5$ $25\pm 2$ $24\pm 3$ $38\pm 1$ $51\pm 1$ $722\pm 32$
+tdm $877\pm 4$ $1050\pm 12$ $1361\pm 51$ $1295\pm 287$ $1312\pm 17$ $1276\pm 73$ $1409\pm 66$
+notmiwae $132\pm 100$ $226\pm 96$ $527\pm 1$ $1049\pm 20$ $608\pm 267$ $1518\pm 543$ $3414\pm 814$
+gain $248\pm 4$ $349\pm 4$ $1084\pm 29$ $1531\pm 29$ $1875\pm 41$ $2980\pm 56$ $5869\pm 153$
+hivae $64\pm 0.228$ $100\pm 1$ $761\pm 7$ $1023\pm 11$ $2529\pm 61$ $3887\pm 82$ $7374\pm 267$
+k__\__means__\__clustering $0.887\pm 0.328$ $1\pm 0.323$ $1\pm 0.120$ $18\pm 2$ $26\pm 1$ $106\pm 23$ $7382\pm 805$
+miss__\__forest $123\pm 28$ $275\pm 90$ $1442\pm 365$ $2558\pm 818$ $4281\pm 1601$ $4384\pm 972$ $19227\pm 4033$
+datawig $247\pm 46$ $232\pm 30$ $581\pm 48$ $2142\pm 380$ $4952\pm 674$ $7048\pm 1119$ $30238\pm 3335$
+automl $1704\pm 78$ $1677\pm 309$ $5147\pm 463$ $6707\pm 992$ $13011\pm 2231$ $19299\pm 2315$ $106635\pm 17043$
+Table 22: Training time (in seconds) of MVI techniques for the multi-mechanism scenario (S10).
+Imputers are sorted by runtime on the folk_emp dataset, and datasets are ordered by the number of rows.
+Dataset shapes reflect training sets with 30% rows with nulls.
+Values represent mean runtime across seeds, with standard deviations included.
+Imputer diabetes (633, 17) german (700, 21) folk__\__inc (12000, 10) law__\__school (16638, 11) bank (32003, 13) heart (56000, 11) folk__\__emp (242112, 16)
+median-dummy $0.009\pm 0.000$ $0.010\pm 0.001$ $0.019\pm 0.002$ $0.021\pm 0.000$ $0.021\pm 0.001$ $0.040\pm 0.001$ $0.590\pm 0.007$
+median-mode $0.009\pm 0.000$ $0.010\pm 0.000$ $0.023\pm 0.002$ $0.019\pm 0.000$ $0.024\pm 0.000$ $0.059\pm 0.003$ $0.748\pm 0.015$
+deletion $0.008\pm 0.000$ $0.009\pm 0.001$ $0.017\pm 0.001$ $0.018\pm 0.000$ $0.039\pm 0.001$ $0.067\pm 0.002$ $0.927\pm 0.013$
+mnar__\__pvae $8\pm 0.717$ $9\pm 0.997$ $18\pm 4$ $12\pm 0.505$ $22\pm 0.776$ $33\pm 0.939$ $169\pm 3$
+edit__\__gain $2\pm 0.126$ $2\pm 0.145$ $14\pm 0.252$ $21\pm 2$ $29\pm 1$ $62\pm 9$ $215\pm 2$
+nomi $10\pm 2$ $13\pm 5$ $17\pm 1$ $16\pm 0.530$ $19\pm 0.962$ $26\pm 1$ $306\pm 38$
+tdm $1048\pm 260$ $1024\pm 16$ $1236\pm 21$ $1114\pm 20$ $1281\pm 40$ $1212\pm 16$ $1463\pm 24$
+notmiwae $122\pm 81$ $123\pm 83$ $507\pm 0.743$ $618\pm 1$ $502\pm 128$ $1082\pm 406$ $2492\pm 1023$
+gain $241\pm 7$ $294\pm 7$ $1126\pm 26$ $1392\pm 19$ $2106\pm 38$ $2960\pm 43$ $6396\pm 241$
+hivae $68\pm 0.624$ $93\pm 1$ $785\pm 31$ $1119\pm 16$ $2405\pm 111$ $3828\pm 264$ $6559\pm 119$
+k__\__means__\__clustering $39\pm 0.618$ $50\pm 0.800$ $104\pm 2$ $1239\pm 23$ $3914\pm 194$ $3820\pm 185$ $7432\pm 873$
+miss__\__forest $114\pm 13$ $207\pm 76$ $1878\pm 553$ $2434\pm 685$ $3966\pm 1017$ $7253\pm 1789$ $27730\pm 5046$
+datawig $1627\pm 541$ $272\pm 55$ $591\pm 46$ $2802\pm 518$ $5892\pm 480$ $7708\pm 468$ $31086\pm 4558$
+automl $1715\pm 149$ $1750\pm 172$ $5843\pm 461$ $6531\pm 509$ $12888\pm 1099$ $18248\pm 1324$ $110488\pm 14166$
+C.2. Additional Accuracy and Fairness Metrics of Predictive Model
 In this section, we extend the results discussed in Section 4 by presenting plots for additional popular accuracy and fairness metrics of downstream model performance.
-Figures 18, 19, 21, and 20 illustrate Accuracy, Disparate Impact, True Negative Rate Difference, and Selection Rate Difference of best performing models, respectively, for different imputation techniques, datasets, and missingness mechanisms.
+Figures 20, 21, 22, and 23 illustrate Accuracy, Disparate Impact, Selection Rate Difference, and True Negative Rate Difference of best performing models, respectively, for different imputation techniques, datasets, and missingness mechanisms.
+Figures 24 and 25 present scatter plots for all imputers that extend scatter plots from Section 4.4.
 Fairness metrics are calculated for an intersectional group when multiple sensitive attributes are present in the dataset, and for a single group when there is only one sensitive attribute (see Table 2 for reference).
@@
 (c) Missing Not At Random (MNAR)
+(d) Mixed missingness (MCAR & MAR & MNAR)
 Figure 20.
@@
 (c) Missing Not At Random (MNAR)
+(d) Mixed missingness (MCAR & MAR & MNAR)
 Figure 21.
@@
 Datasets are ordered in increasing order by size.
-The best performing model for each dataset is reported in the figure.
 The blue line indicates the performance of a model trained on clean data.
@@
 (c) Missing Not At Random (MNAR)
+(d) Mixed missingness (MCAR & MAR & MNAR)
 Figure 22.
@@
 Datasets are ordered in increasing order by size.
-The best performing model for each dataset is reported in the figure.
 The blue line indicates the performance of a model trained on clean data.
@@
 (c) Missing Not At Random (MNAR)
-Figure 21.
-True Negative Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots).
-Values close to 0 are ideal/fair.
-Datasets are ordered in increasing order by size.
-The best performing model for each dataset is reported in the figure.
-The blue line indicates the performance of a model trained on clean data.
+(d) Mixed missingness (MCAR & MAR & MNAR)
+Figure 23. True Negative Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots).
+(a) F1 (imputation) vs F1 (model)
+(b) RMSE (imputation) vs F1 (model)
+(c) KL divergence (imputation) vs F1 (model)
+Figure 24. Imputation quality vs. model performance: imputation correctness (F1, RMSE and KL divergence) may not be indicative of model correctness (F1).
+(a) F1 score difference (imputation) vs TPRD (model)
+(b) RMSE difference (imputation) vs TPRD (model)
+(c) KL divergence difference (imputation) vs TPRD (model)
+Figure 25. Imputation fairness vs model fairness: imputation fairness (F1 difference, RMSE difference and KL divergence difference) may not be indicative of model fairness (TPRD).
 Appendix D Additional Experimental Results for Missingness Shift
@@
 Conversely, for the bank dataset, a predictive model trained on an MCAR train set shows better TPRD on an MNAR test set than on an MCAR test set.
-We further present the impact of missingness shifts on other fairness metrics, decomposed by different model types, in Figures 27-44.
-These plots illustrate the effects on Disparate Impact, Selection Rate Difference, and True Negative Rate Difference across all ML and DL-based imputers and all our datasets.
+We further present the impact of missingness shifts on other fairness metrics, decomposed by different model types, in Figures 31-48.
+These plots illustrate the effects on Disparate Impact, Selection Rate Difference, and True Negative Rate Difference across ML- and DL-based imputers and all our datasets.
 (a) Diabetes and Random Forest
@@
 Other fairness metrics under variable train set missingness.
-Figures 48, 49, and 50 for the diabetes dataset and the Random Forest model illustrate the impact of train set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively.
+Figures 52, 53, and 54 for the diabetes dataset and the Random Forest model illustrate the impact of train set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively.
 Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the train set having 10%, 30%, or 50% error rates, while the test set maintained a constant 30% error rate across all settings.
@@
 Other fairness metrics under variable test set missingness.
-Figures 51, 52, and 53 for the diabetes dataset and the Random Forest model show the impact of test set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively.
+Figures 55, 56, and 57 for the diabetes dataset and the Random Forest model show the impact of test set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively.
 Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the test set having 10%, 20%, 30%, 40%, and 50% error rates, while the train set maintained a constant 30% error rate across all settings.
@@
 This indicates that the missingness mechanism in the train set can significantly affect model fairness on the MNAR test set.
-Additionally, miss-forest and auto-ml proved to be the most robust MVI techniques against test set missingness shifts under all conditions.
+Additionally, miss-forest, auto-ml, and nomi proved to be the most robust MVI techniques against test set missingness shifts under all conditions.
 Figure 52.

### v5

@@
 Different models are the best-performing on different datasets.
-In Figures 34 and 5, we report on the best-performing models (according to F1) for each dataset, and compare performance against a model trained on the clean training set using the same model architecture (performance shown as a blue line in the plots).
+In Figures 2223 and 24, we report on the best-performing models (according to F1) for five most representative datasets per experiment, and compare performance against a model trained on clean data.
+Complete results are available in Appendix C.2.
 (a) Missing Completely At Random (MCAR)
@@
 Datasets are in increasing order by size.
-Blue line shows performance of the model trained on clean data.
+Blue line shows median performance of the model trained on clean data.
 4.1. Correctness of the Predictive Model
@@
 Further, from Figure 25 in Appendix C.2, we can see that the accuracy of the lgbm__\__clf is 0.89, which is marginally better than the accuracy reported using Logistic Regression (0.885) in (Le Quy et al., 2022) in Table 15.
+Also note that a model trained on imputed data can sometimes outperform one trained on clean data, as for german and heart on MCAR.
+We hypothesize that this happens when the model trained on cleaned data learns spurious correlations (e.g., due to noisy or erroneous values, especially in continuous features).
+However, when we simulate nulls, MVI techniques that leverage information from other rows could impute more accurate values, breaking some of these spurious correlations and resulting in slightly better performance.
 (a) Missing Completely At Random (MCAR)
@@
 Datasets are in increasing order by size.
-Blue line shows TPRD of the model trained on clean data.
+Blue line shows median TPRD of the model trained on clean data.
 4.2. Fairness of the Predictive Model
@@
 Figure 5.
-Label Stability of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots).
+Label Stability of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots).
 Values close to 1 are desirable.
-Datasets are ordered in increasing order by size.
-The blue line indicates performance of the model trained on clean data.
+Blue line shows median performance of the model trained on clean data.
 4.3. Stability of the Predictive Model
@@
 clustering, which as was poor-performing according to F1 (correctness) is also poor-performing according to stability on small datasets (diabetes and german, with 905 and 1k samples, respectively).
-A notable and surprising exception is under mixed missingness on german, where clustering is competitive with other impute-then-classify approaches, despite not being competitive on MCAR, MAR or MNAR individually.
-The reason for this is that imputation accuracy of MVI techniques impacts data uncertainty, which then directly impacts overall model uncertainty (Gal et al., 2016).
+A surprising exception is under mixed missingness on german, where clustering is competitive with other impute-then-classify approaches, despite not being competitive on MCAR, MAR or MNAR individually.
+The reason is that imputation accuracy of MVI techniques impacts data uncertainty, which then directly impacts overall model uncertainty (Gal et al., 2016).
 Deletion worsens stability compared to the model trained on clean data for all missingness types on diabetes, but, notably, only under MNAR on german.
-boostclean, which constitutes its own model type, shows a deviation from the stability of the model trained on clean data on all datasets except bank and all missingness types: worsening stability compared to the clean model on folk-income, law-school and folk-employment (except under MCAR), but, surprisingly, improving it on heart, even under mixed missingness.
+boostclean, which constitutes its own model type, shows a deviation from the stability of the model trained on clean data on all datasets except bank: worsening stability compared to the clean model on folk-income, law-school and folk-employment (except under MCAR), but, surprisingly, improving it on heart, even under mixed missingness.
 (a) F1 (imputation) vs F1 (model)
@@
 For example with deletion and clustering, TPRD increases with increase in test missingness rate in scenario S1 (MCAR train, MCAR test) and S3 (MNAR train, MNAR test), but monotonically decreases with test missingness rate in scenario S2 (MAR train, MAR test).
-In summary,
-our results corroborate the findings of Guha et al.
+In summary, our results corroborate the findings of Guha et al.
 (2024), and are a cause for concern as they indicate that the MVI techniques that perform best during development may not preserve fairness post-deployment, where shifts are likely to occur.
@@
 Figure 54 in Appendix 53 shows the Label Stability of the Random Forest model on diabetes as a function of test missingness rate.
-We find that test missingness rate has little effect on model stability under all missingness settings, with the exception of clustering, which shows lower label stability at higher missingness rates, most pronounced in scenario S7 (MAR train, MNAR test).
+We find that test missingness rate has little effect on model stability under all settings except clustering, which shows lower label stability at higher missingness rates, most pronounced in scenario S7 (MAR train, MNAR test).
 6. Running Time
@@
 Interest in deep learning for tabular data has surged (Hwang and Song, 2023; Gorishniy et al., 2021; Borisov et al., 2022; Ye et al., 2024; Fayaz et al., 2022).
-Recall that Figures 3, 4, and 5 show only the best-performing models (by F1) for clarity.
+Recall that Figures 22, 23, and 24 show only the best-performing models (by F1) for clarity.
 Tree-based methods like random forests and gradient-boosted trees outperform deep table-learning models like GANDALF (Joseph and Raj, 2022) on small datasets (diabetes, german), aligning with prior findings (Grinsztajn et al., 2022; Shwartz-Ziv and Armon, 2022).
@@
 In Appendix C, we offer additional results to those discussed in Section 4.
-In particular, Appendix C.1 presents the imputation quality of MVI techniques across all datasets from the perspectives of accuracy and training time, while Appendix C.2 showcases the fairness of predictive models measured by metrics beyond the TPRD mentioned in Section 4 and includes scatter plots for all imputers that extend scatter plots from Section 4.4.
+In particular, Appendix C.1 presents the imputation quality of MVI techniques across all datasets from the perspectives of accuracy and training time, while Appendix C.2 showcases additional metrics for predictive model performance beyond those mentioned in Section 4 and includes scatter plots for all imputers that extend scatter plots from Section 4.4.
 Then, Appendix D presents supplementary experimental findings on the missingness shift discussed in Section 5.
@@
 automl $1715\pm 149$ $1750\pm 172$ $5843\pm 461$ $6531\pm 509$ $12888\pm 1099$ $18248\pm 1324$ $110488\pm 14166$
-C.2. Additional Accuracy and Fairness Metrics of Predictive Model
-In this section, we extend the results discussed in Section 4 by presenting plots for additional popular accuracy and fairness metrics of downstream model performance.
-Figures 20, 21, 22, and 23 illustrate Accuracy, Disparate Impact, Selection Rate Difference, and True Negative Rate Difference of best performing models, respectively, for different imputation techniques, datasets, and missingness mechanisms.
-Figures 24 and 25 present scatter plots for all imputers that extend scatter plots from Section 4.4.
+C.2. Additional Experimental Results for Predictive Model Performance
+In this section, we extend the results discussed in Section 4 by presenting plots for additional metrics of downstream model performance.
+Figures 22, 23, 24, 25, 26, 27, and 28 illustrate F1, True Positive Rate Difference, Label Stability, Accuracy, Disparate Impact, Selection Rate Difference, and True Negative Rate Difference of best performing models, respectively, for different imputation techniques, datasets, and missingness mechanisms.
+Figures 20 and 21 present scatter plots for all imputers that extend scatter plots from Section 4.4.
 Fairness metrics are calculated for an intersectional group when multiple sensitive attributes are present in the dataset, and for a single group when there is only one sensitive attribute (see Table 2 for reference).
@@
 This implies that it is hard to determine in advance whether data cleaning will worsen fairness.
+(a) F1 (imputation) vs F1 (model)
+(b) RMSE (imputation) vs F1 (model)
+(c) KL divergence (imputation) vs F1 (model)
+Figure 20. Imputation quality vs. model performance: imputation correctness (F1, RMSE and KL divergence) may not be indicative of model correctness (F1).
+(a) F1 score difference (imputation) vs TPRD (model)
+(b) RMSE difference (imputation) vs TPRD (model)
+(c) KL divergence difference (imputation) vs TPRD (model)
+Figure 21. Imputation fairness vs model fairness: imputation fairness (F1 difference, RMSE difference and KL divergence difference) may not be indicative of model fairness (TPRD).
+(a) Missing Completely At Random (MCAR)
+(b) Missing At Random (MAR)
+(c) Missing Not At Random (MNAR)
+(d) Mixed missingness (MCAR & MAR & MNAR)
+Figure 22.
+F1 of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots).
+Datasets are in increasing order by size.
+Blue line shows median performance of the model trained on clean data.
+(a) Missing Completely At Random (MCAR)
+(b) Missing At Random (MAR)
+(c) Missing Not At Random (MNAR)
+(d) Mixed missingness (MCAR & MAR & MNAR)
+Figure 23.
+True Positive Rate Difference (unfairness) of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets ($x$-axis), and missingness mechanisms (subplots).
+Values close to 0 are desirable.
+Datasets are in increasing order by size.
+Blue line shows median TPRD of the model trained on clean data.
+(a) Missing Completely At Random (MCAR)
+(b) Missing At Random (MAR)
+(c) Missing Not At Random (MNAR)
+(d) Mixed missingness (MCAR & MAR & MNAR)
+Figure 24.
+Label Stability of best performing models (shown in figure) for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots).
+Values close to 1 are desirable.
+Datasets are ordered in increasing order by size.
+The blue line indicates median performance of the model trained on clean data.
 (a) Missing Completely At Random (MCAR)
@@
 Datasets are ordered in increasing order by size.
-The blue line shows performance of a model trained on clean data.
+The blue line shows median performance of a model trained on clean data.
 (a) Missing Completely At Random (MCAR)
@@
 Datasets are ordered in increasing order by size.
-The blue line indicates the performance of a model trained on clean data.
+The blue line indicates the median performance of a model trained on clean data.
 (a) Missing Completely At Random (MCAR)
@@
 Datasets are ordered in increasing order by size.
-The blue line indicates the performance of a model trained on clean data.
+The blue line indicates the median performance of a model trained on clean data.
 (a) Missing Completely At Random (MCAR)
@@
 Figure 28. True Negative Rate Difference (unfairness) of best performing models for different imputation strategies (colors in the legend), datasets (x-axis), and missingness mechanisms (subplots).
-(a) F1 (imputation) vs F1 (model)
-(b) RMSE (imputation) vs F1 (model)
-(c) KL divergence (imputation) vs F1 (model)
-Figure 24. Imputation quality vs. model performance: imputation correctness (F1, RMSE and KL divergence) may not be indicative of model correctness (F1).
-(a) F1 score difference (imputation) vs TPRD (model)
-(b) RMSE difference (imputation) vs TPRD (model)
-(c) KL divergence difference (imputation) vs TPRD (model)
-Figure 25. Imputation fairness vs model fairness: imputation fairness (F1 difference, RMSE difference and KL divergence difference) may not be indicative of model fairness (TPRD).
 Appendix D Additional Experimental Results for Missingness Shift
@@
 Conversely, for the bank dataset, a predictive model trained on an MCAR train set shows better TPRD on an MNAR test set than on an MCAR test set.
-We further present the impact of missingness shifts on other fairness metrics, decomposed by different model types, in Figures 31-48.
+We further present the impact of missingness shifts on other fairness metrics, decomposed by different model types, in Figures 34-51.
 These plots illustrate the effects on Disparate Impact, Selection Rate Difference, and True Negative Rate Difference across ML- and DL-based imputers and all our datasets.
@@
 Other fairness metrics under variable train set missingness.
-Figures 52, 53, and 54 for the diabetes dataset and the Random Forest model illustrate the impact of train set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively.
+Figures 55, 56, and 57 for the diabetes dataset and the Random Forest model illustrate the impact of train set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively.
 Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the train set having 10%, 30%, or 50% error rates, while the test set maintained a constant 30% error rate across all settings.
@@
 Other fairness metrics under variable test set missingness.
-Figures 55, 56, and 57 for the diabetes dataset and the Random Forest model show the impact of test set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively.
+Figures 58, 59, and 60 for the diabetes dataset and the Random Forest model show the impact of test set missingness on True Negative Rate Difference, Disparate Impact, and Selection Rate Difference, respectively.
 Missingness shifts were modeled for all combinations of MCAR, MAR, and MNAR missingness patterns in the train and test sets, with the test set having 10%, 20%, 30%, 40%, and 50% error rates, while the train set maintained a constant 30% error rate across all settings.

### v6

@@
 Finally, for MNAR, the missingness is determined by missing values themselves, and the likelihood of missing values depends on the missing features.
-Table 3 presents the selected columns ($\mathcal{F}^{m}$) and injection conditions for the diabetes dataset, based on the correlation coefficients and feature importances in Figure 2.
+Table 3 presents the selected columns ($\mathcal{F}^{m}$) and injection conditions for the diabetes dataset, based on the correlation coefficients and feature importance values in Figure 2.
 Additional information on other datasets is available in Appendix B.2.
@@
 $DI=\frac{TP_{dis}+FP_{dis}}{N_{dis}}\div\frac{TP_{priv}+FP_{priv}}{N_{priv}}$
-Broadly speaking, fairness metrics that are based on error rates are consistent with formal equality of opportunity, while those that are based on selection rates (SRD and DI) are consistent with substantive equality of opportunity (Khan et al., 2022).
-Which specific metric is more relevant depends on the application domain and on the stakeholder (Narayanan, 2018).
-SRD focuses on absolute disparities in selection rates (e.g., fixed quotas used in college admissions), while DI focuses on relative disparities (e.g., enforcing the 4/5th rule for non-discrimination in hiring in the U.S.).
-From the point of view of individuals being selected, TPRD is typically used when assessing fairness in opportunity allocation, such as hiring or loan approvals, where a positive outcome should be equally accessible across groups, while TNRD is most relevant in risk assessment and exclusion decisions, like medical diagnoses, where fairness in avoiding false negatives (e.g., failing to detect a disease) matters.
+Fairness metrics based on error rates align with formal equality of opportunity, while those based on selection rates (SRD, DI) reflect substantive equality (Khan et al., 2022).
+The choice of metric depends on the domain and stakeholder (Narayanan, 2018).
+SRD captures absolute disparities (e.g., fixed quotas in college admissions), while DI measures relative disparities (e.g., the 4/5th rule in U.S.
+hiring).
+From the individual’s perspective, TPRD suits opportunity allocation (e.g., hiring, loans), ensuring access to positive outcomes, whereas TNRD applies to exclusion decisions (e.g., medical diagnoses), emphasizing fairness in avoiding false negatives.
 3.4. Shades-of-Null Architecture
 The architecture of Shades-of-Null is shown in Figure 1.
-The core component is the benchmark controller, which executes user-specified missingness scenarios by applying error injectors to input datasets.
-It then imputes missing values using the selected MVI technique(s) and preprocesses the data with standard scaling for numerical features and one-hot encoding for categorical features.
-Next, the controller trains ML models, including hyperparameter tuning.
-The evaluation module then assesses imputation quality and model performance.
+Its core component, the benchmark controller, executes user-specified missingness scenarios by applying error injectors to input datasets.
+It then imputes missing values using selected MVI technique(s) and preprocesses data via standard scaling (numerical) and one-hot encoding (categorical), and then trains ML models with hyperparameter tuning.
+The evaluation module assesses imputation quality and model performance.
 For comprehensive profiling, it uses Virny (Herasymuk et al., 2024), a Python library that computes accuracy, stability, and fairness metrics across multiple sensitive attributes and their intersections.
@@
 This is due to deletion discarding useful information, whereas retaining rows with nulls can still provide valuable signal for model training.
-Note that the F1 score on the bank dataset is extremely low (0.32).
-This is due to severe class imbalance in the dataset (the base rate is 0.117, as reported in Table 9 in Appendix B).
-Further, from Figure 25 in Appendix C.2, we can see that the accuracy of the lgbm__\__clf is 0.89, which is marginally better than the accuracy reported using Logistic Regression (0.885) in (Le Quy et al., 2022) in Table 15.
-Also note that a model trained on imputed data can sometimes outperform one trained on clean data, as for german and heart on MCAR.
-We hypothesize that this happens when the model trained on cleaned data learns spurious correlations (e.g., due to noisy or erroneous values, especially in continuous features).
-However, when we simulate nulls, MVI techniques that leverage information from other rows could impute more accurate values, breaking some of these spurious correlations and resulting in slightly better performance.
+The F1 score on the bank dataset is low (0.32), due to severe class imbalance (base rate 0.117, see Table 9 in Appendix B).
+Figure 25 in Appendix C.2 shows that lgbm__\__clf achieves 0.89 accuracy, slightly outperforming Logistic Regression (0.885) reported in (Le Quy et al., 2022) in Table 15.
+Interestingly, models trained on imputed data can sometimes outperform those trained on clean data, as seen for german and heart under MCAR.
+We hypothesize this occurs when models trained on cleaned data learns spurious correlations (e.g., from noisy or erroneous values), while MVI methods may impute more accurate values, mitigating such artifacts and improving performance.
 (a) Missing Completely At Random (MCAR)
@@
 4.3. Stability of the Predictive Model
-Figure 24 shows label stability of models trained using different MVI techniques.
-In line with conventional wisdom (Domingos, 2000), stability depends most strongly on dataset characteristics (mainly size) and model type.
-Impute-then-classify approaches like miss-forest, auto-ml and datawig, which were the best-performing according to F1, also match the stability of the model trained on clean data under all missingness types and dataset sizes.
-nomi, which performs best on accuracy and training time, demonstrates comparable stability to these top techniques on all datasets.
-clustering, which as was poor-performing according to F1 (correctness) is also poor-performing according to stability on small datasets (diabetes and german, with 905 and 1k samples, respectively).
-A surprising exception is under mixed missingness on german, where clustering is competitive with other impute-then-classify approaches, despite not being competitive on MCAR, MAR or MNAR individually.
-The reason is that imputation accuracy of MVI techniques impacts data uncertainty, which then directly impacts overall model uncertainty (Gal et al., 2016).
+Figure 24 shows label stability of models trained with different MVI techniques.
+In line with conventional wisdom (Domingos, 2000), stability depends primarily on dataset characteristics (especially size) and model type.
+Impute-then-classify methods like miss-forest, auto-ml, and datawig, which perform best on F1, also match the stability of models trained on clean data across missingness types and dataset sizes.
+nomi, which performs best on accuracy and training time, shows comparable stability to these top methods across datasets.
+clustering, which performed poorly on F1, is likewise unstable on small datasets (diabetes and german, with 905 and 1k samples, respectively).
+A notable exception is german under mixed missingness, where clustering is competitive despite underperforming on MCAR, MAR, or MNAR individually.
+This may be because imputation accuracy affects data uncertainty, which in turn drives model uncertainty (Gal et al., 2016).
 Deletion worsens stability compared to the model trained on clean data for all missingness types on diabetes, but, notably, only under MNAR on german.
@@
 Imputation Quality.
-We find that MVI techniques with vastly different imputation quality can give rise to models with comparable F1 scores, indicating that imputation correctness is not sufficiently predictive of model correctness (Zhang and Long, 2021a).
-For example, in Figure 6(a) on the diabetes dataset, median-dummy has an F1 score (imputation) close to 0; auto-ml, miss-forest and nomi have an F1 score (imputation) close to 1; and all other imputers have an F1 score (imputation) between 0.5-0.6, but the models trained with all of these techniques have an F1 close to 1.
-This trend is consistent across all datasets and missingness types, and for numerical columns as well (see Figure 6(b)).
-We observe similar trends for KL divergence (see Figure 6(c)), indicating that neither discrepancy nor distributional imputation quality metrics are sufficiently indicative of downstream model performance, refuting Shadbahr et al.
+MVI techniques with widely varying imputation quality can yield models with similar F1, suggesting that imputation correctness is not a reliable predictor of downstream performance (Zhang and Long, 2021a).
+For instance, in Figure 6(a) on diabetes, median-dummy has imputation F1 near 0; auto-ml, miss-forest, and nomi are near 1; others fall between 0.5–0.6, yet all produce models with F1 close to 1.
+This pattern holds across datasets, missingness types, and for numerical columns (Figure 6(b)).
+Similar trends are observed for KL divergence (Figure 6(c)), reinforcing that neither discrepancy-based nor distributional metrics reliably predict downstream model performance, contradicting Shadbahr et al.
 (2023)’s claim.
@@
 Further, imputation fairness is insufficiently predictive of model fairness.
-For example, on german under mixed missingness, median-dummy has KL difference close to 0, datawig and miss-forest have KL difference between 1 and 1.5, and median-mode and clustering have KL difference between -1 and -1.5, but the models trained using all five of these techniques have a TPRD close to -0.02.
-Conversely, on diabetes under MAR, auto-ml, miss-forest, median-dummy and clustering all have near perfect imputation fairness (KL difference close to 0), but different model fairness (TPRD between 0.04 and 0.12).
+For example, on german under mixed missingness, median-dummy has KL difference close to 0, datawig and miss-forest have KL difference between 1 and 1.5, clustering and median-mode have KL difference between -1 and -1.5, but the models trained using all five of these techniques have a TPRD close to -0.02.
+Conversely, on diabetes under MAR, auto-ml, clustering, miss-forest, and median-dummy all have near perfect imputation fairness (KL difference close to 0), but different model fairness (TPRD between 0.04 and 0.12).
 We see similar trends for other imputation fairness metrics such as F1 score difference (Figure 7(a)) and RMSE difference (Figure 7(b)).
@@
 Test set missingness.
-We measured the impact of test missingness rate on fairness and found that fairness is highly sensitive to shifts in test missingness.
-Figure 53 in Appendix 52 shows the TPRD of the Random Forest model on diabetes as a function of test missingness rate.
+We measured the impact of test missingness rate on fairness and found that fairness is highly sensitive to such shifts.
+Figure 53 in Appendix 52 shows the TPRD of the Random Forest model on diabetes as test missingness increases.
 As before, fairness is highly sensitive to shifts in test missingness.
-For boostclean, datawig, and mnar-pvae, TPRD generally increases (fairness worsens) with an increase in test error rate, although this increase is not monotonic.
-miss-forest, auto-ml, and nomi are particularly robust to increases in test missingness rate under all missingness scenarios.
-For simpler methods such as deletion, clustering, median-mode and median-dummy, there is no consistent trend, even when missingness types remain the same and only missingness rates change between train and test.
-For example with deletion and clustering, TPRD increases with increase in test missingness rate in scenario S1 (MCAR train, MCAR test) and S3 (MNAR train, MNAR test), but monotonically decreases with test missingness rate in scenario S2 (MAR train, MAR test).
+For boostclean, datawig, and mnar-pvae, TPRD generally increases (fairness worsens) as test error rate rises, though not always monotonically.
+miss-forest, auto-ml, and nomi are robust to increases in test missingness rate under all scenarios.
+Simpler methods such as deletion, clustering, median-mode and median-dummy show no consistent trend, even when missingness types remain the same and only missingness rates change between train and test.
+For example, with deletion and clustering, TPRD increases with test missingness in scenarios S1 (MCAR train, MCAR test) and S3 (MNAR train, MNAR test), but decreases in S2 (MAR train, MAR test).
 In summary, our results corroborate the findings of Guha et al.
@@
 This difference is due to the auto-ML nature of auto-ml, which involves extensive hyperparameter and network architecture tuning.
-Among non-statistical techniques, mnar-pvae, edit-gain, and nomi are the most efficient in terms of runtime.
-Notably, nomi provides competitive accuracy comparable to miss-forest, datawig, and auto-ml, making it the most optimal technique in terms of the trade-off between imputation accuracy and training time.
-A particularly noteworthy comparison is between gain and edit-gain.
-As explained in Appendix A.1, EDIT accelerates the training of parametric imputation models.
-Applying EDIT to gain results in a 28x speedup on the folk_emp dataset, with even greater improvements for smaller datasets, as shown in Table 4.
-Importantly, with this running time acceleration, edit-gain achieves accuracy comparable to gain on most datasets, as shown in Figures 18 and 19.
+Among non-statistical techniques, mnar-pvae, edit-gain, and nomi are the most efficient.
+Notably, nomi delivers accuracy on par with miss-forest, datawig, and auto-ml, successfully balancing imputation accuracy and training time.
+A key comparison is between gain and edit-gain.
+As explained in Appendix A.1, edit-gain achieves a 28x speedup on folk_emp, with even greater improvements for smaller datasets, as shown in Table 4, while maintaining comparable accuracy to gain , as shown in Figures 18 and 19.
 7. Summary of Experimental Findings
 Do not drop your nulls!
-Building on evidence in the scientific community (Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), our experiments show that deletion is the least effective strategy for model accuracy, fairness, and stability, especially in setups where each row contributes valuable information.
-Logically, deletion results in data loss, but its appropriateness depends on data quality rather than quantity.
-If rows are complete duplicates or contain erroneous values, deletion may be justified.
+Building on prior evidence (Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), we confirm that deletion is the least effective strategy for model accuracy, fairness, and stability—especially when each row holds valuable information.
+While deletion leads to data loss by design, its suitability depends more on data quality than quantity.
+If rows are duplicates or contain errors, deletion may be warranted.
 Multiple imputation shows mixed results.
@@
 Further, we find that fairness is highly sensitive to changes in missingness rates and missingness mechanisms between training and test sets, which are likely to occur in practice, and therefore a cause for ethical concern.
-Imputation quality and fairness metrics are often insufficiently predictive of the correctness and fairness of downstream models.
+Imputation quality and fairness metrics often fail to predict the correctness or fairness of downstream models.
 Shadbahr et al.
 (2023) argue that distributional imputation quality metrics better predict model performance than discrepancy metrics.
-However, we find that neither reliably indicates downstream performance (F1), as strong learners can compensate for poor imputation quality.
-Moreover, imputation fairness metrics do not predict model fairness: perfect imputation fairness can still lead to model unfairness, while models trained with imputation-unfair MVI techniques can achieve good downstream fairness.
+However, we find that neither reliably predicts downstream performance, as strong learners can compensate for poor imputations.
+Moreover, imputation fairness does not predict model fairness: fair imputers can still yield unfair models, while models trained with fairness-poor MVI techniques can achieve good downstream fairness.
 (a) MCAR missingness
@@
 We find that for large datasets even simple statistical imputers can preserve stability.
-In contrast, for small datasets, only a few MVI techniques preserve stability, while deletion, statistical imputation and more complex ML-based approaches all worsen stability.
+In contrast, for small datasets, only a few MVI techniques do so, while deletion, statistical imputation and complex ML-based MVI all worsen stability.
 Sensitivity to train and test missingness rates.
@@
 MNAR is theoretically the hardest setting to model.
-Empirically, we find that MVI techniques perform poorly with MNAR during testing, even the two existing MNAR-specific methods (not-miwae and mnar-pvae).
-For instance, experimental results show that model F1 and stability are more sensitive to missingness rates under MNAR than under MCAR and MAR.
-This performance gap arises because these methods rely on assumptions that often fail in real-world settings and are evaluated on limited, mostly large datasets with a single scenario, lacking the diversity of our benchmark.
-Developing improved MNAR-specific MVI methods and establishing more comprehensive benchmarks for their evaluation are critical directions for future research.
+MVI techniques, including MNAR-specific not-miwae and mnar-pvae, perform poorly under MNAR.
+For example, F1 and stability are more sensitive to missingness rates under MNAR than under MCAR or MAR.
+This performance gap stems from unrealistic assumptions and limited evaluations, typically on large datasets and a single scenario, lacking the diversity of our benchmark.
+Advancing MNAR-specific MVI methods and building comprehensive benchmarks are key future directions.
 Deep learning outperforms tree-based models on large datasets.
@@
 However, GANDALF excels on larger datasets (heart, folk-employment), highlighting the increasing effectiveness of deep learning across diverse data modalities.
-Best performing approaches under single- and multi-mechanism missingness can be different.
-A key novelty of our study is the evaluation under mixed or multi-mechanism missingness—a more complex but realistic pattern likely encountered in practice.
-We find that performance trends vary between single- and multi-mechanism missingness.
-Figure 11 shows that model types like gandalf__\__clf and lgbm__\__clf are uncorrelated with performance metrics under single-mechanism missingness but become correlated under mixed missingness: F1 correlations are $\rho$=-0.24 for gandalf__\__clf and $\rho$=-0.37 for lgbm__\__clf, while fairness (TPRD) correlations are $\rho$=0.55 for gandalf__\__clf and $\rho$=-0.4 for lgbm__\__clf.
+Best-performing approaches under single- and multi-mechanism missingness can differ.
+A key contribution of our study is the evaluation under mixed (multi-mechanism) missingness — a complex pattern likely to arise in practice.
+We find that performance trends differ between single- and multi-mechanism settings.
+Figure 11 shows that model types like gandalf_clf and lgbm_clf are uncorrelated with performance metrics under single-mechanism missingness, but become correlated under mixed missingness: F1 correlations are $\rho=-0.24$ for gandalf_clf and $\rho=-0.37$ for lgbm_clf, while fairness (TPRD) correlations are $\rho=0.55$ and $\rho=-0.4$, respectively.
 It’s complicated!
@@
 Deep table-learning (gandalf__\__clf) is strongly correlated with fairness (TPRD, $\rho$=0.55), weakly with stability ($\rho$=0.12), and negatively with F1 ($\rho$=-0.24).
-This is not bad news, but rather an acknowledgement of the complexity of learning from incomplete data and the need for rigorous, holistic evaluation protocols, such as those demonstrated in this work, to select the most suitable imputation method and model architecture for a given task.
+This is not bad news, but a recognition of the complexity of learning from incomplete data, and of the need for rigorous, holistic evaluation protocols, like those used here, to identify the best imputation method and model architecture for a given task.
 8. Conclusions, Limitations and Future Work
@@
 Limitations.
-We started with clean datasets, carefully designed meaningful missingness scenarios, and simulated them using error injectors.
-A limitation of our work is that we evaluate performance on synthetically generated — rather than on naturally occurring — missingness.
-This is a common limitation of work in this area: we often do not have a clean ground truth to compare against, and so we make the assumption that injecting synthetic errors will only exacerbate the effect of existing (unknown) errors.
-With this assumption, any degradation in performance we see is an underestimate of the possible degradation we can expect by comparing with an ideal clean ground truth.
+We began with clean datasets, designed meaningful missingness scenarios, and simulated them using error injectors.
+A key limitation is that we evaluate performance on synthetically generated—rather than naturally occurring—missingness.
+This is common in the field: clean ground truth is rarely available, so we assume that injecting synthetic errors only worsens the effects of existing (unknown) ones.
+Under this assumption, any observed performance degradation likely underestimates the true degradation relative to an ideal ground truth.
 Future work.